### v1

# The separating Noether number of small groups

Abstract.
The separating Noether number of a finite group is the minimal positive integer $d$ such that for any finite dimensional complex linear representation of the group, any two dictinct orbits can be distinguished by a polynomial invariant of degree at most $d$. In the present paper its exact value is determined for all groups of order less than $32$, for all finite groups with a cyclic subgroup of index at most $2$, and for the direct product of a dihedral group with the $2$-element group. The results show in particular that unlike the ordinary Noether number, the separating Noether number of a non-abelian finite group may well be equal to the separating Noether number of a proper direct factor of the group. Most of the results are proved for the case of a general (possibly finite) base field containing an element whose multiplicative order equals the size of the group.
Key words and phrases:
polynomial invariants, separating sets, degree bounds, finite groups
2020 Mathematics Subject Classification:
Primary 13A50; Secondary 13P15, 20C15
Partially supported by the Hungarian National Research, Development and Innovation Office, NKFIH K 138828.
1. Introduction
For a finite dimensional vector space $V$ over a field $K$ we write $K[V]$ for the coordinate ring of $V$ identified with the $n$-variable polynomial algebra $K[x_{1},\dots,x_{n}]$, where the variables $x_{1},\dots,x_{n}$ represent the coordinate functions on $V$ with respect to a chosen basis. Let $G$ be a finite group, and $\rho:G\to\mathrm{GL}(V)$ a representation of $G$ on $V$. We shall say that $(V,\rho)$ is a $G$-module in this case; mostly we shall suppress $\rho$ from the notation and we will say that $V$ is a $G$-module. Sometimes we write $V_{K}$ instead of $V$ when we want to make explicit the actual base field. The action of $G$ on $V$ induces an action on $K[V]$ via $K$-algebra automorphisms, such that the variables $x_{1},\dots,x_{n}$ span a $G$-submodule isomorphic to $V^{*}$ (the $G$-module dual to $V$), and $K[V]$ is the symmetric tensor algebra of $V^{*}$. Consider
$K[V]^{G}=\{f\in K[V]\mid\forall g\in G:\quad g\cdot f=f\},$
the subalgebra of $G$-invariants in $K[V]$. For $v\in V$ we shall denote by $G\cdot v$ the $G$-orbit of $V$, and by $\mathrm{Stab}_{G}(v):=\{g\in G\mid gv=v\}$ the stabilizer subgroup in $G$ of $v$.
It is well known that for $v,w\in V$ with $G\cdot v\neq G\cdot w$, there exists an invariant $f\in K[V]^{G}$ with $f(v)\neq f(w)$ (see for example [16, Theorem 3.12.1], [33, Theorem 16] or [23, Lemma 3.1]). Therefore the general notion of a ”separating set of invariants” (see [16, Definition 2.4.1]) reduces to the following in the case of finite groups:
Definition 1.1.
A subset $S$ of $K[V]^{G}$ is called a separating set if for any $v,w\in V$, $f(v)=f(w)$ for all $f\in S$ implies that $v$ and $w$ have the same $G$-orbit.
The result mentioned before Definition 1.1 implies that every generating set of the algebra $K[V]^{G}$ is a separating set. In the structural study of algebras of polynomial invariants of finite groups the focus on separating sets (instead of the more special generating sets) turned out to be useful in finding strengthened versions of known theorems in the invariant theory of finite groups, see for example [24], [25]. Considering explicit computations of invariants, the class of modular representations for which generators of the corresponding algebra of invariants is known is very limited. On the other hand, in several such cases explicit finite separating sets are known, see for example [48], [37]. In the non-modular case, an explicit separating set of multisymmetric polynomials is given in [39]. The systematic study of separating invariants over finite fields began recently in [34], [38].
1.1. Degree bounds
Endow $K[V]$ with the standard grading, so the variables $x_{1},\dots,x_{n}$ span the degree $1$ homogeneous component of $K[V]$. We shall denote by $\deg(f)$ the degree of a non-zero homogeneous element of $K[V]$. Obviously, $K[V]^{G}$ is spanned by homogeneous elements, so it inherits the grading from $K[V]=\bigoplus_{d=0}^{\infty}K[V]_{d}$. We shall write $K[V]^{G}_{\leq d}$ for the sum of the homogeneous components of degree at most $d$. Degree bounds for generators of algebras of invariants have always been in the focus of invariant theory (see e.g. [49]). We set
$\beta(G,V):=\min\{d\in\mathbb{N}_{0}\mid K[V]^{G}_{\leq d}\text{ generates the}K\text{-algebra }K[V]^{G}\},$
$\beta^{K}(G):=\sup\{\beta(G,V)\mid V\text{ is a finite dimensional }G\text{-module}\text{ over }K\}.$
The quantity $\beta^{K}(G)$ is called the Noether number of $G$ (over $K$) because of the classical inequality $\beta^{\mathbb{C}}(G)\leq|G|$ due to E. Noether [40] (the inequality was extended to the non-modular case in [28], [29]). Its systematic study was initiated in [46]. The results of [46] were improved first in [19], [47] and later in [11], [12]. In particular, [15, Table 1] gives the exact value of the Noether number for all non-abelian groups of order less than $32$, and [12] gives the Noether number for all groups containing a cyclic subgroup of index two.
Following [36, Definition 2] (see also [33]) we set
$\beta_{\mathrm{sep}}(G,V):=\min\{d\in\mathbb{N}_{0}\mid K[V]^{G}_{\leq d}\text{ is a separating set}\},$
and
$\beta_{\mathrm{sep}}^{K}(G):=\sup\{\beta_{\mathrm{sep}}(G,V)\mid V\text{ is a finite dimensional }G\text{-module}\text{ over }K\}.$
We shall refer to the above quantities as the separating Noether number. As different orbits of a finite group can be separated by polynomial invariants, we have the obvious inequality
$\beta_{\mathrm{sep}}(G,V)\leq\beta(G,V)\qquad\text{ for any }V,$
and hence
(1) $\beta_{\mathrm{sep}}^{K}(G)\leq\beta^{K}(G).$
A major difference between the Noether number and its separating version is that if the characteristic of the base field $K$ divides the order of $G$, then $\beta^{K}(G)=\infty$ by [42], whereas $\beta_{\mathrm{sep}}^{K}(G)\leq|G|$ also in the modular case (see for example the proof of [16, Theorem 3.12.1], [33, Theorem 16] or [23, Lemma 3.1]) However, in the present paper our focus is on the non-modular setup, when the characteristic of the base field does not divide the order of $G$. In this context an interesting and generally open question is the following:
Question 1.2.
To what extent is the change from ”generating sets” to ”separating sets” reflected in the corresponding degree bounds?
Moreover, interest in these degree bounds arose also in recent applications for data science and machine learning, see [2], [5], [6], [7]. We mention also that problems in signal processing motivated the study of a variant of the Noether number for rational invariants in [4].
The exact value of $\beta_{\mathrm{sep}}^{K}(G)$ for certain finite abelian groups $G$ and an algebraically closed base field of non-modular characteristic was determined recently in [44], [45]. As far as we know, the only non-abelian finite groups whose separating Noether number in the non-modular case is determined in the literature are the non-abelian groups of order $3p$ for some odd prime $p$ (see [9]). For some results on the modular case see [36, Theorem C], [26].
The aim of the present work is to produce a counterpart for the separating Noether number of the results on the Noether number from [12] and [15].
Running assumption. Throughout this paper $G$ will stand for a finite group, and $K$ for a field. Unless explicitly stated otherwise, it is assumed that the characteristic of $K$ does not divide $|G|$.
2. Main results
Our first result builds on [12] (see the equality (12)) and shows that for a group with a cyclic subgroup of index $2$ (their classification will be recalled in Section 4) the inequality (1) holds with equality, namely:
Theorem 2.1.
Let $G$ be a non-cyclic finite group with a cyclic subgroup of index $2$, and assume that $K$ contains an element of multiplicative order $|G|$. Then we have the equality
$\beta_{\mathrm{sep}}^{K}(G)=\frac{1}{2}|G|+\begin{cases}2&\text{ if }G=\mathrm{Dic}_{4m},\text{ $m>1$};\\ 1&\text{ otherwise. }\end{cases}$
Using [12, Corollary 5.5] we are able to determine the separating Noether number for another infinite sequence of non-abelian finite groups:
Theorem 2.2.
For $n\geq 2$ even consider the direct product $G:=\mathrm{D}_{2n}\times\mathrm{C}_{2}$ of the dihedral group of order $2n$ and the cyclic group $\mathrm{C}_{2}$ of order $2$. Assume that $K$ has an element with multiplicative order $n$. Then we have
$\beta_{\mathrm{sep}}^{K}(\mathrm{D}_{2n}\times\mathrm{C}_{2})=n+2.$
There are $16$ non-abelian groups of order less than $32$ not covered by Theorem 2.1 or Theorem 2.2. We discuss them in the following theorem:
Theorem 2.3.
Let $G$ be a non-abelian group of order less than $32$ that does not contain a cyclic subgroup of index at most $2$, and $G$ is not isomorphic to $\mathrm{D}_{2n}\times\mathrm{C}_{2}$. Assume that $K$ contains an element of multiplicative order $|G|$. The value of $\beta_{\mathrm{sep}}^{K}(G)$ is given in the following table, in some cases under the additional assumption that $\mathrm{char}(K)=0$ (this is indicated in the corresponding line of the table):
[…]
The first column in the above table contains the identifier of $G$ in the Small Groups Library of GAP (see [30]); that is, a pair of the form [order, i], where the GAP command $\mathsf{SmallGroup(id)}$ returns the i-th group of order order in the catalogue. The second column gives $G$ in a standard notation or indicates its structure as a direct product or semidirect product (the symbol $\rtimes$ always stands for a semidirect product that is not a direct product, and $\psi$ and $\phi$ in the rows $(16,3)$ and $(16,13)$ stand for two different involutive automorphisms of $\mathrm{C}_{4}\times\mathrm{C}_{2}$). The presentation of $G$ in terms of generators and relations can be found in the beginning of the section that contains the theorem referred to in the last column for the value of $\beta_{\mathrm{sep}}^{K}(G)$. The third column of the table contains $\beta^{K}(G)$; these values are taken from [15]. The fourth column contains $\beta_{\mathrm{sep}}^{K}(G)$.
Note that in some cases the value for $\beta_{\mathrm{sep}}^{K}(G)$ given in the table holds under weaker restrictions on $K$ than the restrictions stated in Theorem 2.3. The exact conditions on $K$ needed for the validity of our arguments are contained in the statement quoted in the last column of the table. In a few instances we used computer calculations to determine generators of concrete invariant algebras. This reduces the validity of some of our results to the case of a base field of characteristic zero. The use of computer could have been avoided at the price of a significant increase of the length of the paper.
For sake of completeness we include a similar table containing $\beta_{\mathrm{sep}}^{K}(G)$ for those abelian groups $G$ of order less than $32$ that do not have a cyclic subgroup of index at most $2$, under the assumption that $K$ contains an element whose multiplicative order equals the exponent of $G$. These results follow from [18], [44], [45], and the explanations in Section 3.1.
[…]
2.1. Failure of strict monotonicity
For a normal subgroup $N$ of $G$, all representations of $G/N$ can be viewed as a representation of $G$ whose kernel contains $N$. Therefore we have the obvious inequalities
$\beta^{K}(G)\geq\beta^{K}(G/N)\text{ and }\beta_{\mathrm{sep}}^{K}(G)\geq\beta_{\mathrm{sep}}^{K}(G/N).$
Moreover, both the Noether number and the separating Noether number are monotone for subgroups as well; that is, for any subgroup $H$ of a finite group $G$ we have
(2) $\beta^{K}(G)\geq\beta^{K}(H)$
(see [46]) and
(3) $\beta_{\mathrm{sep}}^{K}(G)\geq\beta_{\mathrm{sep}}^{K}(H)$
(see [36, Theorem B]). The inequality (2) was sharpened in [13, Theorem 1.2], where the following strict monotonicity of the Noether number was proved:
(4) $\beta^{K}(G)>\beta^{K}(H)\text{ for any subgroup }H\neq G;$
$\beta^{K}(G)>\beta^{K}(G/N)\text{ for any normal subgroup }N\neq\{1_{G}\}\text{ of }G.$
The results of Theorem 2.1 and Theorem 2.3 show that the analogues of (4) (i.e. strict monotonicity) does not hold for the separating Noether number. Indeed, for $K$ as in the cited theorems, we have
$\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{8})=6=\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{8}\times\mathrm{C}_{2}),$
and $\mathrm{Dic}_{8}\times\mathrm{C}_{2}$ has a proper direct factor isomorphic to $\mathrm{Dic}_{8}$. Another similar example is
$\beta_{\mathrm{sep}}^{K}(\mathrm{A}_{4})=6=\beta_{\mathrm{sep}}^{K}(\mathrm{A}_{4}\times\mathrm{C}_{2}).$
Furthermore,
$\beta_{\mathrm{sep}}^{K}(\mathrm{S}_{3}\times\mathrm{C}_{3})=6=\beta_{\mathrm{sep}}^{K}(\mathrm{C}_{6}),$
and the group $\mathrm{S}_{3}\times\mathrm{C}_{3}$ has a subgroup isomorphic to $\mathrm{C}_{6}$. Similar examples exist among the abelian groups. For example, we have
$\beta_{\mathrm{sep}}^{K}(\mathrm{C}_{6}\oplus\mathrm{C}_{6}\oplus\mathrm{C}_{6})=12=\beta_{\mathrm{sep}}^{K}(\mathrm{C}_{6}\oplus\mathrm{C}_{6}\oplus\mathrm{C}_{3})$
by [44, Theorem 6.1].
2.2. Comments on the methods used
We make an essential use of the known results on exact values of the Noether numbers (as summarized in [15]) of the groups considered here. In particular, in each case when the Noether number and the separating Noether number happen to coincide, we just need to come up with a representation and two points with distinct orbits that can not be separated by invariants of strictly smaller degree than the Noether number. In all such cases it turns out that rather small dimensional representations suffice to provide such examples. On the other hand, to settle the cases when the separating Noether number is strictly smaller than the Noether number requires elaborate work: we need a thorough analysis of concrete representations, discussions of stabilizers, group automorphisms, construction of invariants and relative invariants, as well as several ad hoc ideas that help to undertand the set of solutions of systems of polynomial equations. What makes these computations feasible is that by general principles (see the notion of Helly dimension and Lemma 3.7 below) it is sufficient to deal with representations having a small number of irreducible summands.
2.3. Organization of the paper
We begin in Sections 3.1 and 3.2 with a discussion of the dependence of $\beta_{\mathrm{sep}}^{K}(G)$ on the base field $K$, which seems to be stronger than the dependence of $\beta^{K}(G)$ on $K$. In the rest of Section 3 we introduce notation and collect general facts used throughout the paper. We turn in Section 4 to the proof of Theorem 2.1 and Theorem 2.2. In Section 5 we deal with the groups of order less than $32$ for which the separating Noether number equals the Noether number. In Section 6 we cover three groups of the form $H\times\mathrm{C}_{2}$ satisfying $\beta_{\mathrm{sep}}^{K}(H\times\mathrm{C}_{2})=\beta_{\mathrm{sep}}^{K}(H)$. Section 7 is devoted to the study of the separating Noether number of $\mathrm{S}_{3}\times\mathrm{C}_{3}$. In the final Section 8 the separating Noether number is computed for four groups of the form $G\cong\mathrm{C}_{n}\rtimes\mathrm{C}_{k}$ with $\beta_{\mathrm{sep}}^{K}(G)<\beta^{K}(G)$.
3. Preliminaries
3.1. Dependence on the base field of $\beta_{\mathrm{sep}}^{K}(G)$
Lemma 3.1.
Let $K$ be a subfield of a field $L$, and let $V_{K}$ be a $G$-module. Write $V_{L}$ for the $G$-module $L\otimes_{K}V$ over $L$.
  • (i)
    For any non-negative integer $d$, the $L$-vector space $L[V_{L}]^{G}_{d}$ is spanned by its subset $K[V_{K}]^{G}_{d}$.
  • (ii)
    We have the inequality $\beta_{\mathrm{sep}}(G,V_{K})\leq\beta_{\mathrm{sep}}(G,V_{L})$.
  • (iii)
    We have the inequality $\beta_{\mathrm{sep}}^{K}(G)\leq\beta_{\mathrm{sep}}^{L}(G)$.
Proof.
(i) It is well known, and follows from the fact that the space of solutions over $L$ of a system of homogeneous linear equations with coefficients in $K$ is spanned by the subset of solutions over $K$.
(ii) Take any $v,v^{\prime}\in V_{K}\subseteq V_{L}$ with $G\cdot v\neq G\cdot v^{\prime}$. Then there exists an $f\in L[V_{L}]^{G}_{d}$ with $d\leq\beta_{\mathrm{sep}}(G,V_{L})$ such that $f(v)\neq f(v^{\prime})$. It follows by (i) that there exists an $h\in K[V]^{G}_{d}$ with $h(v)\neq h(v^{\prime})$. This clearly shows the desired inequality.
(iii) is an immediate consequence of (ii). ∎
Remark 3.2.
The inequalities in Lemma 3.1 (ii), (iii) may be strict:
(i) First we mention an example in the modular case. Consider the $\mathrm{S}_{n}$-module $\mathbb{F}_{q}^{n}$ endowed with the natural permutation representation, where $q$ is a prime power and $\mathbb{F}_{q}$ is the finite field with $q$ elements. Separating $\mathrm{S}_{n}$-invariants on $\mathbb{F}_{q}^{n}$ were studied (much before the notion of ’separating invariants’ was formally introduced) in [27], [1], and recently in [34], [20]. In particular, we have $\beta_{\mathrm{sep}}(\mathrm{S}_{3},\mathbb{F}_{2}^{3})=2$ by [34, Lemma 4.3] and $\beta_{\mathrm{sep}}(\mathrm{S}_{3},\mathbb{F}_{4}^{3})=3$ by [20, Corollary 4.9].
(ii) An example where the inequality of Lemma 3.1 (ii) is strict and $\mathrm{char}(K)$ does not divide $|G|$ will be provided in the present paper in Proposition 5.6. In the subsequent paper [21] we shall give an example where the inequality of Lemma 3.1 (iii) is strict and $\mathrm{char}(K)$ does not divide $|G|$.
Lemma 3.1 and Remark 3.2 raise the following question (not answered in the present paper):
Question 3.3.
Can the inequality in Lemma 3.1 (ii) be strict for $K$ infinite?
Proposition 3.4.
Let $G$ be a finite abelian group. Then we have
(5) $\beta_{\mathrm{sep}}^{K}(G)\leq\beta_{\mathrm{sep}}^{\mathbb{C}}(G),\text{ with equality when }K\text{ contains an element }$
whose multiplicative order equals the exponent of $\text{whose multiplicative order equals the exponent of }G.$
Proof.
In [18, Corollary 2.6], a characterization purely in terms of $G$ (and not referring to $K$) is given for $\beta_{\mathrm{sep}}^{K}(G)$, valid when $K$ contains an element whose multiplicative order equals the exponent of $G$ (in fact [18] assumes that the base field is algebraically closed, but the proof of the quoted result works verbatim under the present weaker assumption on $K$). So $\beta_{\mathrm{sep}}^{K}(G)=\beta_{\mathrm{sep}}^{\mathbb{C}}(G)$ when $K$ contains an element of multiplicative order $|G|$. In particular, for any $K$ we have $\beta_{\mathrm{sep}}^{\overline{K}}(G)=\beta_{\mathrm{sep}}^{\mathbb{C}}(G)$, where $\overline{K}$ is the algebraic closure of $K$ (recall the running assumption that $|G|$ is invertible in $K$). Note finally that by Lemma 3.1 (iii) we have $\beta_{\mathrm{sep}}^{K}(G)\leq\beta_{\mathrm{sep}}^{\overline{K}}(G)$. ∎
3.2. Comparison with the dependence on the base field of $\beta^{K}(G)$.
In contrast with Remark 3.2, it is shown in [35, Corollary 4.2] that $\beta^{K}(G)$ depends only on $\mathrm{char}(K)$, so
$\beta^{K}(G)=\begin{cases}\beta^{\mathbb{F}_{p}}(G)\text{ if }\mathrm{char}(K)=p>0;\\ \beta^{\mathbb{C}}(G)\text{ if }\mathrm{char}(K)=0.\end{cases}$
Moreover, by [35, Theorem 4.7],
(6) $\beta^{\mathbb{F}_{p}}(G)\geq\beta^{\mathbb{C}}(G)\text{ for all }p\nmid|G|,\text{ with equality for all but finitely many }p.$
It is interesting to compare this with (5), where the inequality for the corresponding separating Noether numbers is proved in the reverse direction. Of course, it maybe that both (5) and (6) always hold with equality. At least no example is known for $p$ and $G$ (where $p\nmid|G|$) for which any of (5) or (6) is a strict inequality. In fact (6) holds with equality for all $p\nmid|G|$ when $G$ is abelian, or when $G$ is one of the groups for which the exact value of their Noether number is given in [15]. For sake of completeness of the picture, we repeat that $\beta^{\mathbb{F}_{p}}(G)=\infty$ if $p$ divides $|G|$ by [42], whereas $\beta_{\mathrm{sep}}^{K}(G)\leq|G|$ for any $K$ (including the case when $\mathrm{char}(K)$ divides $|G|$).
3.3. Direct sum decompositions and multigradings
Assume that we have a $G$-module direct sum decomposition
(7) $V=V_{1}\oplus\cdots\oplus V_{k}.$
View the dual space $V_{i}^{*}$ as the subspace of $V^{*}$ consisting of the linear forms vanishing on $V_{j}$ for $j\neq i$. Then $V^{*}=\bigoplus_{i=1}^{k}V_{i}^{*}$. Choose a dual basis $x_{i}^{(j)}$ ($i=1,\dots,\dim(V_{j})$) in $V_{i}^{*}$, then our set of variables in $K[V]$ is $\{x_{i}^{(j)}\mid\ j=1,\dots,k;\ i=1,\dots,\dim(V_{j})\}$. Associated to the decomposition (7) there is an $\mathbb{N}_{0}^{k}$-grading on $K[V]$: the multihomogeneous component of multidegree $\alpha=(\alpha_{1},\dots,\alpha_{k})$ in $K[V]$ is spanned by the monomials having degree $\alpha_{j}$ in the variables belonging to $V_{j}^{*}$ for each $j=1,\dots,k$. Clearly, the $G$-action on $K[V]$ preserves this multigrading, and thus $K[V]^{G}$ is spanned by multihomogeneous elements. In the sequel whenever we mention a ‘multigrading’ not explicitly defined, this should refer to the multigrading that comes from a direct sum decomposition as in (7), which should be clear from the context. A typical way by which we shall exploit the multigrading is captured by the following obvious statement:
Lemma 3.5.
We have $\beta_{\mathrm{sep}}^{K}(G)\geq d$ if and only if there exists a $G$-module $V=V_{1}\oplus\cdots\oplus V_{k}$, $v,v^{\prime}\in V$, $h\in K[V]^{G}$ with $h(v)\neq h(v^{\prime})$, such that for any multihomogeneous invariant $f\in K[V]^{G}$ with $\deg(f)<d$ we have $f(v)=f(v^{\prime})$.
Note that the above multigrading on $K[V]$ is induced by the following action of the torus $T:=K^{\times}\times\cdots\times K^{\times}$ ($k$ direct factors) on $V$: for $\lambda=(\lambda_{1},\dots,\lambda_{k})\in T$ and $v=(v_{1},\dots,v_{k})\in V$ set $\lambda\cdot v:=(\lambda_{1}v_{1},\dots,\lambda_{k}v_{k})$. This action commutes with the $G$-action on $V$. Consequently, for $v,v^{\prime}\in V$ we have that
(8) $G\cdot v=G\cdot v^{\prime}\text{ if and only if }G\cdot(\lambda v)=G\cdot(\lambda v^{\prime}).$
Later in the text when we say that ’after rescaling, we may assume that $v$ has some special form’, then we mean that we replace $v$ by an appropriate element in its $T$-orbit, and we refer to (8).
3.4. Reduction to multiplicity-free representations
We record a consequence of the results of [23] and Lemma 3.1:
Lemma 3.6.
Let $V_{1},\dots,V_{q}$ be a complete irredundant list of representatives of the isomorphism classes of simple $G$-modules. Assume that
$|K|>(\max\{\dim(V_{j})\mid j=1,\dots,q\}-1)|G|.$
Then we have the equality
$\beta_{\mathrm{sep}}^{K}(G)=\beta_{\mathrm{sep}}(G,V_{1}\oplus\cdots\oplus V_{q}).$
Proof.
An arbitrary $G$-module $V$ is isomorphic to $V_{1}^{m_{1}}\oplus\cdots\oplus V_{q}^{m_{q}}$, where $W^{m}$ stands for the direct sum $W\oplus\cdots\oplus W$ of $m$ copies of $W$. By [23, Corollary 2.6] we have $\beta_{\mathrm{sep}}(G,V)\leq\beta_{\mathrm{sep}}(G,V_{1}^{\dim(V_{1})}\oplus\cdots\oplus V_{q}^{\dim(V_{q})})$, and by [23, Theorem 3.4 (ii)] and [23, Proposition 3.3 (ii)] we have $\beta_{\mathrm{sep}}(G,V_{1}^{\dim(V_{1})}\oplus\cdots\oplus V_{q}^{\dim(V_{q})})=\beta_{\mathrm{sep}}(G,V_{1}\oplus\cdots\oplus V_{q})$. ∎
Therefore to compute $\beta_{\mathrm{sep}}^{K}(G)$ it is sufficient to deal with multiplicity-free representations of $G$ (unless possibly if $K$ is a finite field with too few elements).
3.5. Helly dimension
Lemma 3.6 can be improved for groups whose Helly dimension is ’small’. The Helly dimension $\kappa(G)$ is the minimal positive integer $k$ having the following property: any set of cosets in $G$ with empty intersection contains a subset of at most $k$ cosets with empty intersection (see [17, Section 4]).
Lemma 3.7.
Under the assumptions and notation of Lemma 3.6, there exists a $k\leq\kappa(G)$ and $1\leq j_{1}<\dots<j_{k}\leq q$ with $\beta_{\mathrm{sep}}^{K}(G)=\beta_{\mathrm{sep}}(G,V_{j_{1}}\oplus\cdots\oplus V_{j_{k}})$.
Proof.
This follows from Lemma 3.6 by the same argument as the proof of [22, Lemma 4.1] (formulated in a bit more general setup). ∎
In view of Lemma 3.7, it is helpful to find upper bounds on the Helly dimension. By [17, Lemma 4.2] we have the inequality
(9) $\kappa(G)\leq\mu(G)+1,$
where $\mu(G)$ stands for the maximal size of an intersection independent set of subgroups of $G$. Recall that a set of subgroups of $G$ is intersection independent if none contains the intersection of the others.
Lemma 3.8.
If a set $S$ of intersection independent subgroups in a group contains a cyclic subgroup of prime power order, then $|S|\leq 2$.
Proof.
Let $S$ be an intersection independent set of subgroups of a group $G$ with $|S|\geq 2$ and $H\in S$ with $|H|$ a prime power. Take $J\in S$ with $J\cap H$ of minimal possible order. Since the lattice of subgroups of $H$ is a chain, for any $L\in S$ we have $L\cap H\supseteq H\cap J$, hence in particular $L\supseteq H\cap J$. Since $S$ is intersection independent, it follows that $L\in\{H,J\}$. As $L\in S$ was arbitrary, it means that $S=\{H,J\}$. ∎
3.6. Notational convention.
Denote by $\widehat{G}$ the group of characters of $G$ (i.e. the group of homomorphisms $G\to K^{\times}$, written multiplicatively). For $\chi\in\widehat{G}$ denote by $U_{\chi}$ the $1$-dimensional vector space $K$ endowed with the representation $\chi$; that is, $g\cdot\lambda=\chi(g)\lambda$ for $g\in G$ and $\lambda\in K$. In several cases we shall deal with a $G$-module
(10) $V=W\oplus U,\qquad W=W_{1}\oplus\cdots\oplus W_{l},\qquad U=U_{\chi_{1}}\oplus\cdots\oplus U_{\chi_{m}},$
where the $W_{i}$ are pairwise non-isomorphic irreducible $G$-modules of dimension at least $2$, the $\chi,\dots,\chi_{m}$ are distinct characters of $G$. We shall denote by $x_{1},x_{2},\dots$ the coordinate functions on $W_{1}$, by $y_{1},y_{2},\dots$ the coordinate functions on $W_{2}$, $z_{1},z_{2},\dots$ the coordinate functions on $W_{3}$, etc., and denote by $t_{\chi}$ the coordinate function on $U_{\chi}$. According to the above conventions, we have $g\cdot t_{\chi}=\chi(g^{-1})t_{\chi}$. Associated to the above direct sum decomposition of $V$ is an $\mathbb{N}_{0}^{l+m}$-grading on $K[V]$ explained at the beginning of Section 3. The phrase ‘multihomogeneous’ will refer to this multigrading.
Usually we shall specify a concrete representation of $G$ (in other words, a $G$-module) by giving a group homomorphism $\psi:G\to\mathrm{GL}_{n}(K)$. The underlying vector space is then the space of column vectors $V=K^{n}$, on which $g\in G$ operates via multiplication by the matrix $\psi(g)$. The action of $G$ on the variables $x_{1},\dots,x_{n}$ (which form a basis in the dual space $V^{*}$ of $V$) is given explicitly by the following formula:
$g\cdot x_{j}=\sum_{i=1}^{n}\psi(g^{-1})_{ji}x_{i}.$
We shall write $A^{T}$ for the transpose of the matrix (e.g. row vector) $A$.
3.7. Relative invariants and the Hilbert ideal
Given a $G$-module $V$, denote by $K[V]^{G}_{+}$ the sum of the positive degree homogeneous components of $K[V]^{G}$. A homogeneous invariant $f\in K[V]^{G}$ is said to be indecomposable if $f\notin(K[V]^{G}_{+})^{2}$. Clearly, $K[V]^{G}$ is generated by the positive degree homogeneous indecomposable invariants. The so-called Hilbert ideal $\mathcal{H}(G,V)$ is the ideal in the polynomial algebra $K[V]$ generated by the positive degree homogeneous $G$-invariants. The Hilbert ideal plays an essential role in studying generators of the algebra of invariants, because a positive degree homogeneous invariant $f\in K[V]^{G}$ is indecomposable if and only if $f$ is not contained in the ideal $\mathcal{H}(G,V)K[V]_{+}$ of $K[V]$. More precisely, a set of homogeneous invariants form a minimal generating set of $K[V]^{G}$ if and only if it gives a vector space basis in a vector space direct complement of $(K[V]^{G}_{+})^{2}$ in $K[V]^{G}_{+}$ (the sum of the homogeneous components of $K[V]^{G}$ of positive degree).
An element $f\in K[V]$ is a relative invariant of weight $\chi\in\widehat{G}$ if $g\cdot f=\chi(g^{-1})f$ for all $g\in G$ (so $f(gv)=\chi(g)f(v)$ for all $g\in G$ and $v\in V$). We set
$K[V]^{G,\chi}:=\{f\in K[V]\mid f\text{ is a relative invariant of weight }\chi\}.$
For example, if $V$ is as in (10), then the variable $t_{\chi}$ is a relative invariant of weight $\chi$. Moreover, for a sequence $\chi^{(1)},\dots,\chi^{(k)}$ of elements of $\widehat{G}$, $t_{\chi^{(1)}}\cdots t_{\chi^{(k)}}$ is an indecomposable $G$-invariant if and only if $\chi^{(1)}\cdots\chi^{(k)}=1\in\widehat{G}$ and there is no $\ell<k$, $1\leq i_{1}<\cdots<i_{\ell}\leq k$ with $\chi^{(i_{1})}\cdots\chi^{(i_{\ell})}=1\in\widehat{G}$ (we say in this case that $\chi^{(1)},\dots,\chi^{(k)}$ is an irreducible product-one sequence over $\widehat{G}$, and we call $k$ its length). We say that $\chi^{(1)},\dots,\chi^{(k)}$ is a product-one free sequence over $\widehat{G}$ if there is no $1\leq i_{1}<\dots<i_{\ell}\leq k$ with $\chi^{(i_{1})}\cdots\chi^{(i_{\ell})}=1\in\widehat{G}$. The Davenport constant $\mathsf{D}(\widehat{G})$ of the finite abelian group $G$ is defined as the maximal length of an irreducible product-one sequence over $G$. Clearly, the maximal possible length of a product-one free sequence is $\mathsf{D}(\widehat{G})-1$.
Lemma 3.9.
Let $V=W\oplus U$ be a $G$-module as in (10). Let $A$ be a homogeneous generating system of $K[W]^{G}$, and for $\chi\in\{\chi_{1},\dots,\chi_{m}\}$ let $A_{\chi}\subset K[W]^{G,\chi}$ be a set of homogeneous relative invariants of weight $\chi$ that span a vector space direct complement in $K[W]^{G,\chi}$ of $\mathcal{H}(G,W)\cap K[W]^{G,\chi}=K[W]^{G}_{+}K[W]^{G,\chi}$. Set
$B:=\{t_{\chi^{(1)}}\cdots t_{\chi^{(k)}}\mid\chi^{(1)},\dots\chi^{(k)}\text{ is an irreducible product-one sequence over }\widehat{G}\},$
and
$C:=\{ht_{\chi^{(1)}}\cdots t_{\chi^{(k)}}\mid\chi^{(1)},\dots\chi^{(k)}\text{ is a product-one free sequence over }\widehat{G},$
$\quad h\in A_{\chi}\text{ where }\chi^{-1}=\chi^{(1)}\cdots\chi^{(k)}\}.$
Then $A\cup B\cup C$ is a homogeneous generating system of $K[V]^{G}$.
Proof.
As we pointed out above, the algebra $K[V]^{G}$ is generated by indecomposable multihomogeneous invariants. A multihomogeneous invariant $f$ belongs to $K[W]^{G}$ or $K[U]^{G}$ or is of the form $ht$, where $t$ is a monomial in $K[U]$ of positive degree, and $h$ is a multihomogeneous element of positive degree in $K[W]$. In the first two cases $f$ belongs to the $K$-subalgebra of $K[V]$ generated by $A\cup B$. In the third case $t=t_{\chi^{(1)}}\cdots t_{\chi^{(k)}}$, where $\chi^{(1)},\dots\chi^{(k)}$ is a product-one free sequence over $\widehat{G}$. Then $t$ is a relative $G$-invariant of some weight $\chi$, and hence necessarily, $h\in K[V]^{G,\chi^{-1}}$. There exists a linear combination $a$ of elements from $A_{\chi^{-1}}$ with $h-a\in K[W]^{G}_{+}K[W]^{G,\chi^{-1}}$. Then $at$ belongs to the $K$-subspace of $K[V]$ generated by $C$, and $f-at=(h-a)t\in(K[V]^{G}_{+})^{2}$. So we showed that any indecomposable multihomogenous element of $K[V]^{G}$ can be reduced modulo $(K[V]^{G}_{+})^{2}$ to an element in the $K$-subalgebra of $K[V]$ generated by $A\cup B\cup C$. This clearly implies our statement. ∎
Given a set $S\subset K[V]$ of polynomials write
$\mathcal{V}(S):=\{v\in V\mid\forall f\in S\colon f(v)=0\}$
for the common zero locus in $V$ of the elements of $S$. Next we restate [9, Lemma 4.3] in a sharp form:
Lemma 3.10.
We have the equality
$\mathcal{V}(K[V]^{G,\chi})=\{v\in V\mid\mathrm{Stab}_{G}(v)\nsubseteq\ker(\chi)\}.$
Proof.
Suppose that $v\in V$, $g\in G$ with $g\cdot v=v$. For any $f\in K[V]^{G,\chi}$ we have
$f(v)=f(g\cdot v)=\chi(g)f(v),$
therefore $\chi(g)\neq 1$ implies $f(v)=0$. This shows the inclusion $\{v\in V\mid\mathrm{Stab}_{G}(v)\nsubseteq\ker(\chi)\}\subseteq\mathcal{V}(K[V]^{G,\chi}$. The reverse inclusion is stated and proved in [9, Lemma 4.3]. ∎
3.8. The use of automorphisms
The natural action of the automorphism group of $G$ on the isomorphism classes of representations of $G$ helps to save some computations later, thanks to the following statements:
Lemma 3.11.
Let $(V,\rho)$ be a $G$-module and $\alpha$ an automorphism of the group $G$.
  • (i)
    For any $\chi\in\widehat{G}$ we have
    $K[(V,\rho)]^{G,\chi}=K[(V,\rho\circ\alpha)]^{G,\chi\circ\alpha}.$
    In particular,
    $K[(V,\rho)]^{G}=K[(V,\rho\circ\alpha)]^{G}.$
  • (ii)
    We have
    $\beta(G,(V,\rho))=\beta(G,(V,\rho\circ\alpha))\text{ and }\beta_{\mathrm{sep}}(G,(V,\rho))=\beta_{\mathrm{sep}}(G,(V,\rho\circ\alpha)).$
  • (iii)
    Assume that $(V,\rho)\cong(V,\rho\circ\alpha)$ as $G$-modules, so there exists a $K$-vector space isomorphism $T:V\to V$ with $T\circ\rho(g)=(\rho\circ\alpha)(g)\circ T$ for all $g\in G$. Then the map $f\mapsto f\circ T$ induces a $K$-vector space isomorphism
    $K[(V,\rho)]^{G,\chi}\cong K[[(V,\rho)]^{G,\chi\circ\alpha}.$
Proof.
(i) and (ii): The subset $\rho(G)$ of $\mathrm{GL}(V)$ coincides with the subset $(\rho\circ\alpha)(G)$ of $\mathrm{GL}(V)$, therefore the representations $\rho$ and $\rho\circ\alpha$ yield the same partition of $V$ into the union of $G$-orbits, and the same polynomials in $K[V]$ are invariants (respectively relative invariants) for the $G$-action given by $\rho$ as for the $G$-action given by $\rho\circ\alpha$. Moreover, if $f\in K[(V,\rho)]^{G,\chi}$, $g\in G$ and $v\in V$, then $f((\rho\circ\alpha)(g)(v))=f(\rho(\alpha(g))(v)=\chi(\alpha(g))f(v)=(\chi\circ\alpha)(g)f(v)$. So if a relative invariant $f\in K[V]$ has weight $\chi$ with respect to $\rho$, then it has weight $\chi\circ\alpha$ with respect to $\rho\circ\alpha$. This explains both statements.
(iii): For each $g\in G$ we have
$(f\circ T)(gv)=f((T\circ\rho(g))(v))=f((\rho\circ\alpha)(g)(T(v))=\chi(\alpha(g))f(T(v))=(\chi\circ\alpha)(g)(f\circ T)(v).$
4. Groups with a cyclic subgroup of index two
Denote by $\mathrm{C}_{n}$ the cyclic group of order $n$. We use the notation for semidirect products
$\mathrm{C}_{m}\rtimes_{d}\mathrm{C}_{n}=\langle a,b\mid a^{m}=1,b^{n}=1,bab^{-1}=a^{d}\rangle\quad\text{ where }d\in\mathbb{N}\mbox{ is coprime to }m.$
We need the following classification of finite groups with a cyclic subgroup of index $2$ (see [12, Section 10] for details and references):
Proposition 4.1.
Any finite group containing a cyclic subgroup of index two is isomorphic to
(11) $\mathrm{C}_{s}\times(\mathrm{C}_{r}\rtimes_{-1}H)$
where $r,s$ are coprime odd integers, $\mathrm{C}_{r}\rtimes_{-1}H$ is the semidirect product where the kernel of the action of $H$ on $\mathrm{C}_{r}$ contains an index two cyclic subgroup $\langle a\rangle$ of $H$, $b\in H\setminus\langle a\rangle$ acts via inversion on $\mathrm{C}_{r}$ (i.e. $bxb^{-1}=x^{-1}$ for all $x\in\mathrm{C}_{r}$), and $H$ is one of the following $2$-groups:
  • (i)
    $\mathrm{C}_{2^{n}}$  ($n\geq 1$);
  • (ii)
    $\mathrm{C}_{2^{n-1}}\times\mathrm{C}_{2}$  ($n\geq 2$);
  • (iii)
    $\mathrm{M}_{2^{n}}:=\mathrm{C}_{2^{n-1}}\rtimes_{d}\mathrm{C}_{2},\qquad d={2^{n-2}+1}$  ($n\geq 3$);
  • (iv)
    $\mathrm{D}_{2^{n}}:=\mathrm{C}_{2^{n-1}}\rtimes_{-1}\mathrm{C}_{2}$  ($n\geq 4$);
  • (v)
    $\mathrm{SD}_{2^{n}}:=\mathrm{C}_{2^{n-1}}\rtimes_{d}\mathrm{C}_{2},\qquad d={2^{n-2}-1}$  ($n\geq 4$);
  • (vi)
    $\mathrm{Dic}_{2^{n}}:=\langle a,b\mid a^{2^{n-1}}=1,b^{2}=a^{2^{n-2}},bab^{-1}=a^{-1}\rangle$  ($n\geq 3$).
Among the groups listed in Proposition 4.1 are the dicyclic groups. For a positive integer $m>1$, the dicyclic group is
$\mathrm{Dic}_{4m}=\begin{cases}\mathrm{C}_{r}\rtimes_{-1}\mathrm{Dic}_{2^{n}}&\text{ for }m=2^{n-2}r\text{ even }\\ \mathrm{C}_{m}\rtimes_{-1}\mathrm{C}_{4}&\text{ for }m>1\text{ odd. }\end{cases}$
Note that $\mathrm{Dic}_{8}$ is the quaternion group of order $8$.
Proposition 4.2.
Let $G$ be a non-cyclic group with a cyclic subgroup of index two. Assume that $K$ has an element of multiplicative order $|G|$. Then
$\beta_{\mathrm{sep}}^{K}(G)\geq\frac{1}{2}|G|+\begin{cases}2&\text{ if }G=\mathrm{Dic}_{4m},\text{ $m>1$};\\ 1&\text{ otherwise. }\end{cases}$
Proof.
Let $G$ be a group as in (11) in Proposition 4.2, where $H$ is a semidirect product of a cyclic group and the two-element group (so $H$ is $\mathrm{C}_{2}$ or $H$ is of type (ii), (iii), (iv), or (v) from Proposition 4.2). Then $G$ has a matrix representation generated by the matrices
$\ \begin{bmatrix}0&1&0\\ 1&0&0\\ 0&0&-1\end{bmatrix},\quad\begin{bmatrix}\xi&0&0\\ 0&\xi^{k}&0\\ 0&0&1\end{bmatrix},\quad\begin{bmatrix}\rho&0&0\\ 0&\rho^{-1}&0\\ 0&0&1\end{bmatrix},\quad\begin{bmatrix}\varepsilon&0&0\\ 0&\varepsilon&0\\ 0&0&1\end{bmatrix},$
where $k$ is some positive integer coprime to $2^{n-1}$ depending on the type of $H$, and $\xi$, $\rho$, $\varepsilon$ are roots of $1$ of multiplicative order $2^{n-1}$, $r$, $s$. Consider the corresponding algebra $K[x_{1},x_{2},t]^{G}$, where $x_{1},x_{2},t$ are the coordinate functions on the vector space $K^{3}$, on which the group $G$ acts via the given matrix representation. Setting $m:=sr2^{n-1}$, the monomials $x_{1}^{m}$ and $x_{2}^{m}$ are fixed by the cyclic index two subgroup $A$ of $G$ generated by the second, third, and fourth matrices above. Moreover, they are interchanged by the first matrix. It follows that $(x_{1}^{m}-x_{2}^{m})t$ is a $G$-invariant. This degree $m+1$ invariant takes different values on the vectors $v:=[1,0,1]^{T}$ and $v^{\prime}:=[1,0,-1]^{T}$ in $K^{3}$.
Note that our representation is the direct sum of a $2$-dimensional and a $1$-dimensional representation. So by Lemma 3.5 it is sufficient to show that $v$ and $v^{\prime}$ can not be separated by a multihomogeneous invariant (cf. Section 3.3) of degree at most $m$. Suppose for contradiction that $f(v)\neq f(v^{\prime})$, where $f$ is a multihomogeneous invariant of degree at most $m$. So $f=t^{d}f_{1}(x_{1},x_{2})$. Since the projections of these two vectors in the $x_{1},x_{2}$ coordinate plane are the same, no invariant depending only on $x_{1}$ and $x_{2}$ can separate them. Thus we have $d>0$. The invariants depending only on $t$ are the polynomials in $t^{2}$, which takes the same value (namely $1$) on these vectors, hence $\deg(f_{1})>0$. Since $t$ is $A$-invariant, $f_{1}(x_{1},x_{2})\in K[x_{1},x_{2}]^{A}$, so $f_{1}$ is a $K$-linear combination of $A$-invariant monomials in $x_{1}$ and $x_{2}$. The monomials having positive degree in $x_{2}$ vanish on both $v$ and $v^{\prime}$. It follows that $f_{1}$ must have an $A$-invariant monomial depending only on $x_{1}$. The only such monomials are the powers of $x_{1}^{m}$. Thus $\deg(f)\geq m+d>m$, a contradiction.
Assume next that $G$ is a dicyclic group $\mathrm{Dic}_{4m}$. It is isomorphic to the matrix group generated by
$\begin{bmatrix}0&\mathrm{i}\\ \mathrm{i}&0\\ \end{bmatrix},\quad\begin{bmatrix}\rho&0\\ 0&\rho^{-1}\\ \end{bmatrix}$
where $\mathrm{i}$ is an element of $K$ with multiplicative order $4$ (so $\mathrm{i}^{2}=-1$) and $\rho$ is an element of $K$ of multiplicative order $2m$. It is an easy exercise to show that the corresponding algebra of invariants is generated by $x_{1}^{2m}-x_{2}^{2m}$, $(x_{1}x_{2})^{2}$, $(x_{1}^{2m}+x_{2}^{2m})x_{1}x_{2}$ when $m$ is odd, and by $x_{1}^{2m}+x_{2}^{2m}$, $(x_{1}x_{2})^{2}$, $(x_{1}^{2m}-x_{2}^{2m})x_{1}x_{2}$ when $m$ is even. If $m$ is odd, the vectors $[1,1]^{T}$ and $[1,-1]^{T}$ are separated by the degree $22m+$ 2 generator, but all the smaller degree generators agree on them. By assumption $K$ has an element $\nu$ of multiplicative order $4m=|G|$. For even $m$ the vectors $[\nu,1]^{T}$ and $[\nu,-1]^{T}$ are separated by the degree $22m+$ 2 generator, whereas the smaller degree generators coincide on them.
A matrix representation of $\mathrm{C}_{s}\times\mathrm{Dic}_{4m}$ (where $s>1$ is odd and is co-prime to $m$) is generated by
$\begin{bmatrix}0&\mathrm{i}&0\\ \mathrm{i}&0&0\\ 0&0&-1\end{bmatrix},\quad\begin{bmatrix}\rho&0&0\\ 0&\rho^{-1}&0\\ 0&0&1\end{bmatrix},\quad\begin{bmatrix}\varepsilon&0&0\\ 0&\varepsilon&0\\ 0&0&1\end{bmatrix},$
where $\rho$ and $\varepsilon$ are roots of $1$ with multiplicative order $2m$ and $s$. The vectors $v:=[1,0,1]^{T}$ and $v^{\prime}:=[1,0,-1]^{T}$ are separated by the invariant $(x_{1}^{2ms}-x_{2}^{2ms})t$ when $m$ is odd and by $(x_{1}^{2ms}+x_{2}^{2ms})t$ when $m$ is even. On the other hand, any invariant of degree at most $2ms$ takes the same value on $v$ and $v^{\prime}$ (this can be seen similarly to the argument in the second paragraph of the proof, using Lemma 3.5).
It remains to deal with the case when $G\cong\mathrm{C}_{s}\times(\mathrm{C}_{r}\rtimes_{-1}\mathrm{C}_{2^{n}})$ where $n\geq 3$. A matrix representation of this group is generated by
$\begin{bmatrix}0&\omega&0\\ \omega&0&0\\ 0&0&-1\end{bmatrix},\quad\begin{bmatrix}\rho&0&0\\ 0&\rho^{-1}&0\\ 0&0&1\end{bmatrix},\quad\begin{bmatrix}\varepsilon&0&0\\ 0&\varepsilon&0\\ 0&0&1\end{bmatrix},$
where $\omega$, $\rho$ and $\varepsilon$ are roots of $1$ with multiplicative order $2^{n}$, $r$, and $s$. The square of the first matrix together with the second and the third generate a cyclic subgroup $A$ of $G$ of order $m:=2^{n-1}rs$, acting on the line spanned by $[1,0,0]^{T}$ via a character of order $m$. The invariant $(x_{1}^{m}+x_{2}^{m})t$ separates $v:=[1,0,1]^{T}$ and $v^{\prime}:=[1,0,-1]^{T}$, and any invariant of degree at most $m$ takes the same value on $v$ and $v^{\prime}$ (this can be seen similarly to the argument in the second paragraph of the proof, using Lemma 3.5). ∎
Proof of Theorem 2.1. The Noether number $\beta^{K}(G)$ was computed for all groups with a cyclic subgroup of index $2$ in [12, Theorem 10.3], where (as a special case of a statement on the so-called generalized Noether number) it was shown that if $G$ is a non-cyclic group with a cyclic subgroup of index two and $\mathrm{char}(K)$ is not a divisor of $|G|$, then
(12) $\beta^{K}(G)=\frac{1}{2}|G|+\begin{cases}2&\text{ if }G=\mathrm{Dic}_{4m},\text{ $m>1$};\\ 1&\text{ otherwise. }\end{cases}$
Therefore the desired result on the separating Nother number is an immediate consequence of Proposition 4.2, the obvious inequality $\beta_{\mathrm{sep}}^{K}(G)\leq\beta^{K}(G)$, and (12). $\square$
4.1. The groups $\mathrm{D}_{2n}\times\mathrm{C}_{2}$
In this section $n\geq 2$ is even, and
$G=\mathrm{D}_{2n}\times\mathrm{C}_{2}=\langle a,b,c\mid a^{n}=b^{2}=c^{2}=1,\ bab=a^{-1},\ ac=ca,\ bc=cb\rangle.$
Proposition 4.3.
Assume that $K$ contains an element of multiplicative order $n$. Then we have the inequality $\beta_{\mathrm{sep}}^{K}(G)\geq n+2$.
Proof.
Let $\rho$ denote an element of $K$ of multiplicative order $n$. Take the direct sum of the representations
$a\mapsto\begin{bmatrix}\rho&0\\ 0&\rho^{-1}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix},\quad c\mapsto\begin{bmatrix}1&0\\ 0&1\\ \end{bmatrix},$
$a\mapsto 1,\quad b\mapsto-1,\quad c\mapsto-1\qquad\text{ and }\qquad a\mapsto 1,\quad b\mapsto 1,\quad c\mapsto-1.$
Recall that $x_{1},x_{2}$ stand for the standard coordinate functions on the $2$-dimensional summand, and $t_{1},t_{2}$ stand for the coordinate functions on the $1$-dimensional summands. The invariant $(x_{1}^{n}-x_{2}^{n})t_{1}t_{2}$ separates the points $v:=([1,0]^{T},1,1)$ and $v^{\prime}:=([1,0]^{T},1,-1)$ in $K^{4}$. It is sufficient to show that $v$ and $v^{\prime}$ can not be separated by an invariant of degree at most $n+1$. Suppose for contradiction that $h(v)\neq h(v^{\prime})$ for some homogeneous $h\in K[x_{1},x_{2},t_{1},t_{2}]^{G}$ of degree at most $n+1$. We may assume by Lemma 3.5 that $h$ is multihomogeneous (so it is homogeneous both in $t_{1}$ and $t_{2}$), and has minimal possible total degree. So $h=t_{1}^{k_{1}}t_{2}^{k_{2}}h_{1}(x_{1},x_{2})$. Since the $(x_{1},x_{2},t_{1})$-coordinates of $v$ and $v^{\prime}$ are the same, we have $k_{2}>0$. Since $h_{1}$ is $\langle c\rangle$-invariant, $k_{1}+k_{2}$ must be even. Moreover, $K[t_{1},t_{2}]^{G}=K[t_{1}^{2},t_{2}^{2}]$, and $t_{1}^{2}(v)=1=t_{1}^{2}(v^{\prime})$, $t_{2}^{2}(v)=1=t_{2}^{2}(v^{\prime})$, so $\deg(h_{1})>0$, and (since $h$ has minimal possible degree) $h=t_{1}t_{2}h_{1}$. Thus $h_{1}$ is a relative $G$-invariant of degree at most $n-1$, on which $G$ acts via the character $a\mapsto 1,$ $b\mapsto-1$, $c\mapsto 1$. The action of $G$ on $K^{2}$ factors through the action of the dihedral group $\mathrm{D}_{2n}\cong G/\langle c\rangle$, and it is well known (and easy to see) that the smallest degree relative invariant in $K[x_{1},x_{2}]$ with the above weight is $x_{1}^{n}-x_{2}^{n}$. We reached the desired contradiction. ∎
Proof of Theorem 2.2. We have the isomorphism
$\mathrm{D}_{2n}\times\mathrm{C}_{2}\cong(\mathrm{C}_{n}\times\mathrm{C}_{2})\rtimes_{-1}\mathrm{C}_{2}.$
Therefore by [12, Corollary 5.5] we have
$\beta^{K}(\mathrm{D}_{2n}\times\mathrm{C}_{2})=\mathsf{D}(\mathrm{C}_{n}\times\mathrm{C}_{2})+1=n+2,$
implying in turn that $\beta_{\mathrm{sep}}^{K}(\mathrm{D}_{2n}\times\mathrm{C}_{2})\leq n+2$. The reverse inequality is given in Proposition 4.3. $\square$
5. Further groups $G$ with $\beta_{\mathrm{sep}}^{K}(G)=\beta^{K}(G)$
5.1. The alternating group $\mathrm{A}_{4}$
Proposition 5.1.
For $n\geq 4$ we have the inequality $\beta_{\mathrm{sep}}^{K}(\mathrm{A}_{n})\geq n(n-1)/2$.
Proof.
Our running assumption that $\mathrm{char}(K)$ does not divide $|G|$ forces for $G=\mathrm{A}_{n}$ that $|K|\geq n$ if $n\geq 4$. Therefore $K$ has $n$ distinct elements $v_{1},\dots,v_{n}$. The vectors $v:=[v_{1},v_{2},v_{3},\dots,v_{n}]^{T}$ and $v^{\prime}:=[v_{2},v_{1},v_{3},v_{4},\dots,v_{n}]^{T}$ belong to the same orbit under the symmetric group $\mathrm{S}_{n}$ acting on $K^{n}$ via permuting the coordinates, but $v$ and $v^{\prime}$ have different orbits with respect to the alternating subgroup $\mathrm{A}_{n}$. It is well known that the corresponding algebra of invariants is
$K[x_{1},\dots,x_{n}]^{\mathrm{A}_{n}}=K[x_{1},\dots,x_{n}]^{\mathrm{S}_{n}}\oplus\Delta K[x_{1},\dots,x_{n}]^{\mathrm{S}_{n}},$
where $\Delta:=\sum_{1\leq i<j\leq n}(x_{i}-x_{j})$. This shows that all $\mathrm{A}_{n}$-invariants of degree less than $n(n-1)/2$ are in fact $\mathrm{S}_{n}$-invariants. So they can not separate $v$ and $v^{\prime}$, which have different $\mathrm{A}_{n}$-orbits. ∎
It is known that $\beta^{K}(\mathrm{A}_{4})=6$ (see [11, Theorem 3.4]), hence Proposition 5.1 has the following consequence:
Theorem 5.2.
We have the equality $\beta_{\mathrm{sep}}^{K}(\mathrm{A}_{4})=6$.
5.2. The binary tetrahedral group
In this section $G:=\widetilde{\mathrm{A}}_{4}$ is the binary tetrahedral group: the special orthogonal group $\mathrm{SO}(3,\mathbb{R})$ has a subgroup isomorphic to $\mathrm{A}_{4}$ (the group of rotations mapping a regular tetrahedron into itself), and $G$ is isomorphic to the preimage of $\mathrm{A}_{4}$ under the classical two-fold covering $\mathrm{SU}(2,\mathbb{C})\to\mathrm{SO}(3,\mathbb{R})$. So $G$ can be embedded as a subgroup into the special unitary group $\mathrm{SU}(2,\mathbb{C})$, acting via its defining representation on $V:=\mathbb{C}^{2}$. Thus $V$ is a $G$-module, and it is well known that $\mathbb{C}[V]^{G}$ is generated by three homogeneous invariants $f_{6}$, $f_{8}$, $f_{12}$, having degree $6$, $8$, and $12$ (see [31, Lemma 4.1] or [41, Section 0.13]). Since $G$ (as a subgroup of $\mathrm{GL}(V)$) is not generated by pseudo-reflections, $\mathbb{C}[V]^{G}$ does not contain an algebraically independent separating set by [24], so the invariants $f_{6}$ and $f_{8}$ can not separate all the $G$-orbits in $V$. This implies that $\beta_{\mathrm{sep}}^{\mathbb{C}}(G,V)\geq 12$. A second proof valid for more general base fields is given below.
Theorem 5.3.
Suppose in addition that $K$ contains an element $\xi$ of multiplicative order $8$. Then we have $\beta_{\mathrm{sep}}^{K}(\widetilde{\mathrm{A}}_{4})=12$.
Proof.
An alternative way to think of $G=\widetilde{\mathrm{A}_{4}}$ is that $G\cong\mathrm{Dic}_{8}\rtimes\mathrm{C}_{3}$, the non-abelian semidirect product of the quaternion group of order $8$ and the $3$-element group. Denote by $a,b$ generators of $G$, where $a^{4}=1$, $b^{3}=1$, and $\langle a,bab^{-1}\rangle\cong\mathrm{Dic}_{8}$, such that $a,bab^{-1},b^{2}ab^{-2}$ correspond to the elements of the quaternion group traditionally denoted by $\mathrm{i}$, $\mathrm{j}$, $\mathrm{k}$. Then $G$ has the representation
$\psi:a\mapsto\begin{bmatrix}\xi^{2}&0\\ 0&-\xi^{2}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}\frac{1}{2}(-1-\xi^{2})&\frac{1}{2}(1+\xi^{2})\\ \frac{1}{2}(-1+\xi^{2})&\frac{1}{2}(-1+\xi^{2})\\ \end{bmatrix}$
(indeed, $\psi(b)^{3}$ is the identity matrix, and $\psi(b)\psi(a)\psi(b)^{-1}=\begin{bmatrix}0&-1\\ 1&0\\ \end{bmatrix}$ together with $\psi(a)$ generate a subgroup in $\mathrm{GL}(K^{2})$ isomorphic to the quaternion group). Write $V$ for $K^{2}$ endowed with the representation $\psi$. It is easy to see that $K[V]^{\mathrm{Dic}_{8}}$ is generated by $f_{1}:=(x_{1}x_{2})^{2}$, $f_{2}:=(x_{1}^{2}+x_{2}^{2})^{2}$, $f_{3}:=x_{1}x_{2}(x_{1}^{4}-x_{2}^{4})$. The generator $f_{3}$ is fixed by $b$ as well, hence $f_{3}$ is $G$-invariant, whereas $f_{1}$ and $f_{2}$ span a $\langle b\rangle$-invariant subspace in $K[x_{1},x_{2}]$. The matrix of the linear transformation $b$ on the space $\mathrm{Span}_{K}\{f_{1},f_{2}\}$ with respect to the basis $4f_{1},f_{2}$ is $\begin{bmatrix}0&1\\ -1&-1\\ \end{bmatrix}.$ Endow the $3$-variable polynomial algebra $K[s_{1},s_{2},s_{3}]$ with the action of $\langle b\rangle\cong\mathrm{C}_{3}$ given by $s_{1}\mapsto-s_{2}$, $s_{2}\mapsto s_{1}-s_{2}$, $s_{3}\mapsto s_{3}$. The $K$-algebra homomorphism given by $s_{1}\mapsto 4f_{1}$, $s_{2}\mapsto f_{2}$, $s_{3}\mapsto f_{3}$ restricts to a surjective $K$-algebra homomorphism from $K[s_{1},s_{2},s_{3}]^{\langle b\rangle}$ onto $K[V]^{G}$. Knowing that $K[s_{1},s_{2},s_{3}]^{\langle b\rangle}$ is generated in degree at most $\mathsf{D}(\mathrm{C}_{3})=3$, an easy direct computation yields
$K[s_{1},s_{2},s_{3}]^{\langle b\rangle}=K[s_{1}^{2}-s_{1}s_{2}+s_{2}^{2},s_{1}^{2}s_{2}-s_{1}s_{2}^{2},s_{1}^{3}-3s_{1}s_{2}^{2}+s_{2}^{3},s_{3}].$
This shows that $K[V]^{G}$ is generated by
$h_{8}$ $:=(4f_{1})^{2}-4f_{1}f_{2}+f_{2}^{2}$
$h_{6}$ $:=x_{1}x_{2}(x_{1}^{2}+x_{2}^{2})(x_{1}^{2}-x_{2}^{2})$
$h_{12}^{(1)}$ $:=(4f_{1})^{2}f_{2}-4f_{1}f_{2}^{2}$
$h_{12}^{(2)}$ $:=(4f_{1})^{3}-12f_{1}f_{2}^{2}+f_{2}^{3}$
having degree $4,6,12,12$. In fact $h_{12}^{(1)}$ can be omitted, because we have $h_{12}^{(1)}=-4h_{6}^{2}$. Set $v:=[1,\xi^{2}]^{T}$ and $v^{\prime}:=\xi v$. Then $(x_{1}^{2}+x_{2}^{2})(v)=0=(x_{1}^{2}+x_{2}^{2})(v^{\prime})$, hence $f_{2}(v)=0=f_{2}(v^{\prime})$. It follows that $h_{6}(v)=0=h_{6}(v^{\prime})$. Moreover, $h_{8}(v)=16(\xi^{4})^{2}=16=16(\xi^{8})^{2}=h_{8}(v^{\prime})$. Thus no invariants of degree less than $12$ separate $v$ and $v^{\prime}$. However, they have different $G$-orbit, since $h_{12}^{(2)}(v)=4^{3}(\xi^{4})^{3}=-64$, whereas $h_{12}^{(2)}(v^{\prime})=4^{3}(\xi^{8})^{3}=64$. Thus we proved the inequality $\beta_{\mathrm{sep}}(G,V)\geq 12$, implying in turn $\beta_{\mathrm{sep}}^{K}(G)\geq 12$. On the other hand, by [11, Corollary 3.6] we know that $\beta^{K}(G)=12$, implying the reverse inequality $\beta_{\mathrm{sep}}^{K}(G)\leq 12$. ∎
5.3. The symmetric group $\mathrm{S}_{4}$ of degree $4$.
Theorem 5.4.
Assume that $K$ has an element of multiplicative order $24$ or $\mathrm{char}(K)\neq 5$. Then we have $\beta_{\mathrm{sep}}^{K}(\mathrm{S}_{4})=9$.
Proof.
We have $\beta^{K}(\mathrm{S}_{4})=9$ by [14, Example 5.3], implying the inequality $\beta_{\mathrm{sep}}^{K}(\mathrm{S}_{4})\leq 9$.
In order to show the reverse inequality consider $V=K^{4}$ endowed with the tensor product of the sign representation of $\mathrm{S}_{4}$ and the standard action of $\mathrm{S}_{4}$ on $K^{4}$ via permuting the coordinates. Thus $\pi\in\mathrm{S}_{4}$ maps the variable $x_{j}$ to $\mathrm{sign}(\pi)x_{\pi(j)}$. Denote by $\sigma_{j}$ ($j=1,2,3,4)$ the elementary symmetric polynomial of degree $j$, and set $\Delta:=\prod_{1\leq i<j\leq 4}(x_{i}-x_{j})$. The algebra $K[V]^{\mathrm{S}_{4}}$ is minimally generated by $\sigma_{1}^{2}$, $\sigma_{2}$, $\sigma_{1}\sigma_{3}$, $\sigma_{4}$, $\sigma_{3}^{2}$, $\sigma_{1}\Delta$, $\sigma_{3}\Delta$ (see [14, Example 5.3]).
We shall show that there exists a $v\in V$ with
(13) $\sigma_{1}(v)=0,\quad\Delta(v)\neq 0\quad\text{ and }\sigma_{3}(v)\neq 0.$
Then setting $v^{\prime}=-v$ we have that all the even degree invariants agree on $v$ and $v^{\prime}$. The only odd degree invariant of degree less than $9$ is $\sigma_{1}\Delta$ (up to non-zero scalar multiples), and it vanishes both on $v$ and $v^{\prime}$ by $\sigma_{1}(v)=0=\sigma_{1}(v^{\prime})$. On the other hand, $\Delta(v)=\Delta(v^{\prime})\neq 0$, $\sigma_{3}(v)=-\sigma_{3}(v^{\prime})\neq 0$ imply that $(\sigma_{3}\Delta)(v)\neq(\sigma_{3}\Delta)(v^{\prime})$. This shows that $\beta_{\mathrm{sep}}^{K}(\mathrm{S}_{4})\geq 9$.
It remains to prove the existence of $v$ satisfying (13). If $\mathrm{char}(K)\neq 5$, then $v:=[0,1,2,-3]^{T}\in V$ satisfies (13) (recall that by the running assumption $\mathrm{char}(K)\notin\{2,3\}$). If $\mathrm{char}(K)=5$ and $K$ contains an element of multiplicative order $24$, then $K$ contains $\mathbb{F}_{25}$, the field of $25$ elements. Let $v_{1},v_{2}$ be the roots of the polynomial $x^{2}+x+1$ in $\mathbb{F}_{25}\subseteq K$, and let $v_{3},v_{4}$ be the roots of $x^{2}-x+2$ in $\mathbb{F}_{25}\subseteq K$. Set $v:=[v_{1},v_{2},v_{3},v_{4}]^{T}$. Note that
$(x^{2}+x+1)(x^{2}-x+2)=x^{4}+2x^{2}+x+2\in K[x],$
showing that $\sigma_{1}(v)=0$, $\sigma_{3}(v)=-1$. Moreover, $v_{1},v_{2},v_{3},v_{4}$ are distinct, so $\Delta(v)\neq 0$, and thus (13) holds. ∎
5.4. The Pauli group
In this section $G$ stands for the Pauli group $G$ of order $16$. Assume the $K$ has an element $\mathrm{i}$ of multiplicative order $4$. Then $G$ is generated by the matrices
$\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix},\quad\begin{bmatrix}0&-\mathrm{i}\\ \mathrm{i}&0\\ \end{bmatrix},\quad\begin{bmatrix}1&0\\ 0&-1\\ \end{bmatrix}$
in $\mathrm{GL}(K^{2})$.
Theorem 5.5.
Assume that $K$ has an element of multiplicative order $4$. We have the equality $\beta_{\mathrm{sep}}^{K}(G)=7$.
Proof.
Since $G$ is generated by reflections, by the Sheppard-Todd-Chevalley Theorem (see e.g. [3, Theorem 7.2.1] or [16, Section 3.9.4]), $K[x_{1},x_{2}]^{G}$ is a polynomial algebra with two generators. In fact it is generated by $x_{1}^{4}+x_{2}^{4}$ and $(x_{1}x_{2})^{2}$. Moreover, the factor of $K[x_{1},x_{2}]$ modulo the Hilbert ideal (the ideal generated by $x_{1}^{4}+x_{2}^{4}$ and $(x_{1}x_{2})^{2}$) is isomorphic to the regular representation of $G$ (see [8]). The element $f:=(x_{1}^{4}-x_{2}^{4})x_{1}x_{2}$ has the property that $g\cdot f=-f$ where $g$ is any of the above three generators of $G$. One can easily see that $f$ does not belong to the Hilbert ideal. Take an extra indeterminate $t$ and extend the action of $G$ to $K[x_{1},x_{2},t]$ by setting $g\cdot t=-t$ for any of the above three generators of $G$. Then $ft$ is a degree $7$ $G$-invariant separating $v:=[1,1,1]^{T}$ and $v^{\prime}:=[1,1,-1]^{T}$. We claim that $v$ and $v^{\prime}$ can not be separated by an invariant of degree at most $6$. Suppose for contradiction that $h$ is a multihomogeneous $G$-invariant of degree at most $6$ with $h(v)\neq h(v^{\prime})$ (cf. Lemma 3.5). Then $h$ is homogeneous in $t$. Since $v$ and $v^{\prime}$ have the same $(x_{1},x_{2})$ coordinates, any element of $K[x_{1},x_{2}]$ takes the same value on them, so $h$ has positive degree in $t$. Moreover, since $t^{2}$ is $G$-invariant, we may assume that $h=th_{1}$ for some $h_{1}\in K[x_{1},x_{2}]$, where $h_{1}$ is homogeneous of degree at most $5$. By $G$-invariance of $h$ we must have that $g\cdot h_{1}=-h_{1}$ for any of the above three generators $g$ of $G$. However, this $1$-dimensional representation of $G$ occurs only once as a summand in the factor of $K[x_{1},x_{2}]$ modulo the Hilbert ideal, and we found already one occurrence in the degree $6$ homogeneous component. This contradiction proves the inequality $\beta_{\mathrm{sep}}^{K}(G)\geq 7$. The reverse inequality $\beta_{\mathrm{sep}}^{K}(G)\leq 7$ follows from the equality $\beta^{K}(G)=7$ proved in [14, Example 5.4]. ∎
5.5. The Heisenberg group $\mathrm{H}_{27}$
In this section
$G=\mathrm{H}_{27}=\langle a,b,c\mid a^{3}=b^{3}=c^{3}=1,\ a^{-1}b^{-1}ab=c,\ ac=ca,\ bc=cb\rangle$
is the Heisenberg group of order $27$. It is generated by $a,b$, and its commutator subgroup $\langle c\rangle\cong\mathrm{C}_{3}$, whereas $G/G^{\prime}\cong\mathrm{C}_{3}\times\mathrm{C}_{3}$.
Assume that $K$ has an element $\omega$ of multiplicative order $3$, and consider the following irreducible $3$-dimensional representation of $G$:
$\psi:a\mapsto\begin{bmatrix}1&0&0\\ 0&\omega&0\\ 0&0&\omega^{2}\end{bmatrix},\quad b\mapsto\begin{bmatrix}0&0&1\\ 1&0&0\\ 0&1&0\end{bmatrix},\quad c\mapsto\begin{bmatrix}\omega&0&0\\ 0&\omega&0\\ 0&0&\omega\end{bmatrix}.$
Write $V$ for the vector space $K^{3}$ endowed with the representation $\psi$.
Proposition 5.6.
We have $\beta_{\mathrm{sep}}(G,V)\geq 9$ if $|K|\neq 4$, whereas $\beta_{\mathrm{sep}}(G,V)=6$ if $|K|=4$.
Proof.
Denote by $H$ the subgroup of $G$ generated by $a$ and $c$. The algebra $K[V]^{H}$ is generated by the monomials $x_{1}^{3}$, $x_{2}^{3}$, $x_{3}^{3}$, $x_{1}x_{2}x_{3}$. It follows that each homogeneous $G$-invariant has degree divisible by $3$. The element $b$ fixes $x_{1}x_{2}x_{3}$, and permutes cyclically $x_{1}^{3}$, $x_{2}^{3}$, $x_{3}^{3}$. So $K[V]^{G}$ is generated by $x_{1}x_{2}x_{3}$ and the generators of $K[x_{1}^{3},x_{2}^{3},x_{3}^{3}]^{\langle b\rangle}$. One easily deduces that the $\langle b\rangle$-invariants in $K[V]^{H}$ (i.e. the $G$-invariants in $K[V]$) of degree at most $6$ are contained in the subalgebra generated by $f_{1}:=x_{1}x_{2}x_{3}$, $f_{2}:=x_{1}^{3}+x_{2}^{3}+x_{3}^{3}$, $f_{3}:=x_{1}^{3}x_{2}^{3}+x_{1}^{3}x_{3}^{3}+x_{2}^{3}x_{3}^{3}$. If $|K|\neq 4$ then there exists a non-zero $\lambda\in K$ with $\lambda^{3}\neq 1$. Now $f_{1}$, $f_{2}$, and $f_{3}$ are all symmetric in the variables $x_{1},x_{2},x_{3}$, therefore they agree on $v:=[1,\lambda,0]^{T}$ and $v^{\prime}:=[\lambda,1,0]^{T}$. On the other hand, $v$ and $v^{\prime}$ have different $G$-orbits. Indeed, the algebra $K[V]^{G}$ contains also $f_{4}:=x_{1}^{3}x_{2}^{6}+x_{1}^{6}x_{3}^{3}+x_{2}^{3}x_{3}^{6}$. We have $f_{4}(v)=\lambda^{6}$, whereas $f_{4}(v^{\prime})=\lambda^{3}$, so $f_{4}(v)\neq f_{4}(v^{\prime})$. Thus $G\cdot v\neq G\cdot v^{\prime}$, and $v$, $v^{\prime}$ can not be separated by invariants of degree $<9$.
If $|K|=4$, then for any positive integer $n$, the polynomial $x_{j}^{3n}$ induces the same function on $V$ as $x_{j}^{3}$ ($j=1,2,3$). Therefore all functions on $V$ that are induced by some element of $K[x_{1}^{3},x_{2}^{3},x_{3}^{3}]$ can be induced by a polynomial in
$\mathrm{Span}_{K}\{1,x_{1}^{3},x_{2}^{3},x_{3}^{3},x_{1}^{3}x_{2}^{3},x_{1}^{3}x_{3}^{3},x_{2}^{3}x_{3}^{3},x_{1}^{3}x_{2}^{3}x_{3}^{3}\}.$
Consequently, $x_{1}x_{2}x_{3}$ together with the $G$-invariants of degree at most $6$ in the above space of polynomials form a separating set in $K[V]^{G}$. In particular, we have $\beta_{\mathrm{sep}}(G,V)\leq 6$. To see the reverse inequality, note that the degree $3$ invariants are linear combinations of $x_{1}x_{2}x_{3}$ and $x_{1}^{3}+x_{2}^{3}+x_{3}^{3}$, and they agree on $v:=[0,0,0]^{T}$ and $v^{\prime}:=[1,1,0]^{T}$. However, $v$ and $v^{\prime}$ have different $G$-orbits, since $f_{3}(v)=0\neq 1=f_{3}(v^{\prime})$. Thus $\beta_{\mathrm{sep}}(G,V)=6$ when $|K|=4$. ∎
Theorem 5.7.
Assume that $K$ contains an element of multiplicative order $3$ and $|K|\neq 4$. Then we have $\beta_{\mathrm{sep}}^{K}(\mathrm{H}_{27})=9$.
Proof.
By [10, Corollary 18] we have $\beta^{K}(G)=9$, therefore the result follows by Proposition 5.6. ∎
5.6. The non-abelian group $(\mathrm{C}_{2}\times\mathrm{C}_{2})\rtimes\mathrm{C}_{4}$
In this section
$G:=\langle a,b,c\mid a^{2}=b^{2}=c^{4}=1,\quad ab=ba,\quad cac^{-1}=b,\quad cbc^{-1}=a\rangle.$
Theorem 5.8.
Assume that $K$ has an element $\mathrm{i}$ of multiplicative order $4$. Then we have the equality $\beta_{\mathrm{sep}}^{K}((\mathrm{C}_{2}\times\mathrm{C}_{2})\rtimes\mathrm{C}_{4})=6$.
Proof.
The equality $\beta^{K}(G)=6$ is proved in [15, Proposition 3.10], so it is sufficient to prove that $\beta_{\mathrm{sep}}^{K}(G)\geq 6$. The element $c^{2}$ belongs to the center of $G$, and the factor group $G/\langle c^{2}\rangle$ is isomorphic to the dihedral group $\mathrm{D}_{8}$. The representation
$a\mapsto\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}0&-1\\ -1&0\\ \end{bmatrix},\quad c\mapsto\begin{bmatrix}1&0\\ 0&-1\\ \end{bmatrix}$
has $c^{2}$ in its kernel and gives the defining representation of $\mathrm{D}_{8}$ as the group of symmetries of a square in the euclidean plane. The element $ab$ also belongs to the center of $G$, and the factor group $G/\langle ab\rangle\cong\mathrm{C}_{4}\times\mathrm{C}_{2}$, generated by the coset of $c$ and the coset of $a$. Thus $G$ has the $1$-dimensional representations
$a\mapsto 1,\quad b\mapsto 1,\quad c\mapsto\mathrm{i}\qquad\text{ and }\qquad a\mapsto-1,\quad b\mapsto-1,\quad c\mapsto\mathrm{i}.$
Take the direct sum of the above $2$-dimensional and $1$-dimensional representations, and denote by $x_{1},x_{2}$ the standard coordinate functions on the $2$-dimensional summand, and by $t_{1},t_{2}$ the coordinate functions on the $1$-dimensional summands. The corresponding ring $K[x_{1},x_{2},t_{1},t_{2}]^{G}$ of invariants contains the invariant $h:=(x_{1}^{2}-x_{2}^{2})t_{1}^{3}t_{2}$, separating the points $v:=([1,0]^{T},1,1)$ and $v^{\prime}:=([1,0]^{T},1,-1)$ in $K^{4}$. By Lemma 3.5, in order to prove $\beta_{\mathrm{sep}}^{K}(G)\geq 6$, it is sufficient to show that $v$ and $v^{\prime}$ can not be separated by a multihomogeneous invariant (cf. Section 3.3) of degree at most $5$. Suppose for contradiction that $f(v)\neq f(v^{\prime})$, where $f$ is a multihomogeneous invariant of degree at most $5$. It has positive degree in $t_{2}$, because the $(x_{1},x_{2},t_{1})$ coordinates of $v$ and $v^{\prime}$ are the same. Any $G$-invariant depending only on $t_{1}$ and $t_{2}$ is a polynomial in $t_{1}^{4}$, $t_{1}^{2}t_{2}^{2}$, $t_{2}^{4}$, and all these polynomials take the value $1$ both at $v$ and $v^{\prime}$. So $f=t_{1}^{k_{1}}t_{2}^{k_{2}}f_{1}(x_{1},x_{2})$, where $0<k_{1}+k_{2}\leq 4$, $\deg(f_{1})>0$, and $k_{2}$ is odd. Note that $t_{1}^{k_{1}}t_{2}^{k_{2}}$ is a relative $G$-invariant, hence $f_{1}(x_{1},x_{2})$ must be a relative $G$-invariant, on which $G$ acts by the inverse character. Since $c^{2}$ fixes $f_{1}$, it must fix also $t_{1}^{k_{1}}t_{2}^{k_{2}}$, so $k_{1}+k_{2}$ is even. Consequently, $(k_{1},k_{2})\in\{(1,1),(3,1),(1,3)\}$, and $f_{1}\in K[x_{1},x_{2}]$ is a relative $\mathrm{D}_{8}\cong G/\langle c^{2}\rangle$-invariant of degree at most $5-(k_{1}+k_{2})$. There is no degree $1$ relative $G$-invariant in $K[x_{1},x_{2}]$, hence $t_{1}^{k_{1}}t_{2}^{k_{2}}=t_{1}t_{2}$, and so $f$ is a relative invariant of degree at most $3$. It is easy to see that up to non-zero scalar multiples, the only relative $G$-invariants in $K[x_{1},x_{2}]_{\leq 3}$ with non-trivial weight are $x_{1}x_{2}$ and $x_{1}^{2}-x_{2}^{2}$, but none of $x_{1}x_{2}t_{1}t_{2}$ or $(x_{1}^{2}-x_{2}^{2})t_{1}t_{2}$ is a $G$-invariant. We arrived at the desired contradiction, finishing the proof. ∎
5.7. The group $(\mathrm{C}_{3}\times\mathrm{C}_{3})\rtimes_{-1}\mathrm{C}_{2}$
In this section
$G:=(\mathrm{C}_{3}\times\mathrm{C}_{3})\rtimes_{-1}\mathrm{C}_{2}=\langle a,b,c\mid 1=a^{3}=b^{3}=c^{2},\ ab=ba,\ cac^{-1}=a^{-1},\ cbc^{-1}=b^{-1}\rangle.$
Assume that $K$ contains an element $\xi$ of multiplicative order $6$. So $\omega:=\xi^{2}$ has multiplicative order $3$, and consider the following irreducible $2$-dimensional representations of $G$:
$\psi_{1}:a\mapsto\begin{bmatrix}\omega&0\\ 0&\omega^{2}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}1&0\\ 0&1\\ \end{bmatrix},\quad c\mapsto\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix},$
$\psi_{2}:a\mapsto\begin{bmatrix}1&0\\ 0&1\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}\omega&0\\ 0&\omega^{2}\\ \end{bmatrix},\quad c\mapsto\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix}.$
Denote by $W_{j}$ the vector space $K^{2}$ endowed with the representation $\psi_{j}$ for $j=1,2$ (see Section 3.6).
Proposition 5.9.
We have $\beta_{\mathrm{sep}}(G,W_{1}\oplus W_{2})\geq 6$.
Proof.
Note that $\ker(\psi_{1})=\langle b\rangle$ and $\ker(\psi_{2})=\langle a\rangle$. We have $G/\ker(\psi_{1})\cong\mathrm{S}_{3}$ and $G/\ker(\psi_{2})\cong\mathrm{S}_{3}$. Both these representations factor through the irreducible $2$-dimensional representation of $\mathrm{S}_{3}$. It is well known and easy to see that the corresponding algebra of invariants is generated by a degree $2$ and a degree $3$ invariant. Moreover, $K[W_{j}]/\mathcal{H}(G,W_{j})$ as an $\mathrm{S}_{3}=G/\ker(\psi_{j})$-module is isomorphic to the regular representation of $\mathrm{S}_{3}$. Thus $x_{1}^{3}-x_{2}^{3}$ spans the only minimal non-trivial $G$-invariant subspace in a $G$-module direct complement of $\mathcal{H}(G,W_{1})$ in $K[W_{1}]$ on which $\ker(\psi_{2})$ acts trivially, and $y_{1}^{3}-y_{2}^{3}$ spans the only minimal non-trivial $G$-invariant subspace in a $G$-module direct complement of $\mathcal{H}(G,W_{2})$ in $K[W_{2}]$ on which $\ker(\psi_{1})$ acts trivially. It follows that $K[W_{1}\oplus W_{2}]^{G}$ is generated by $K[W_{1}]^{G}$, $K[W_{2}]^{G}$, and $f:=(x_{1}^{3}-x_{2}^{3})(y_{1}^{3}-y_{2}^{3})$. Consider the points $v=(w_{1},w_{2})=([1,0]^{T},[1,0]^{T})$ and $v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2})=([1,0]^{T},[0,1]^{T})$. We claim that all invariants of degree at most $5$ agree on $v$ and on $v^{\prime}$, but $f(v)\neq f(v^{\prime})$. Indeed, $w_{1}=w^{\prime}_{1}$ and $c\cdot w_{2}=w^{\prime}_{2}$. Thus the invariants in $K[W_{1}]^{G}$ and in $K[W_{2}]^{G}$ agree on $v$ and on $v^{\prime}$. On the other hand, $f(v)=1$ and $f(v^{\prime})=-1$. The proof is finished. ∎
Theorem 5.10.
Assume that $K$ has an element of multiplicative order $6$. Then we have $\beta_{\mathrm{sep}}^{K}((\mathrm{C}_{3}\times\mathrm{C}_{3})\rtimes_{-1}\mathrm{C}_{2})=6$.
Proof.
By [12, Corollary 5.5] we have $\beta^{K}(G)=6$. Thus the result follows from Proposition 5.9. ∎
5.8. The group $\mathrm{C}_{3}\rtimes\mathrm{D}_{8}\cong(\mathrm{C}_{6}\times\mathrm{C}_{2})\rtimes\mathrm{C}_{2}$.
In this section
$G=\mathrm{C}_{3}\rtimes\mathrm{D}_{8}=(\mathrm{C}_{6}\times\mathrm{C}_{2})\rtimes\mathrm{C}_{2}=\langle a,b,c\mid a^{6}=b^{2}=c^{2}=1,\ ba=ab,\ cac=a^{-1},\ cbc=a^{3}b\rangle.$
Assume that $K$ contains an element $\omega$ of multiplicative order $6$, and consider the following irreducible $2$-dimensional representation of $G$:
$\psi:a\mapsto\begin{bmatrix}\omega&0\\ 0&\omega^{-1}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}1&0\\ 0&-1\\ \end{bmatrix},\quad c\mapsto\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix}.$
The commutator subgroup of $G$ is $\langle a\rangle$. Denote by $\chi$ the $1$-dimensional representation of $G$ given by
$\chi:a\mapsto 1,\ b\mapsto-1,\ c\mapsto-1$
Denote by $W$ the vector space $K^{2}$ endowed with the representation $\psi$, and denote by $U_{\chi}$ the vector space $K$ endowed with the representation $\chi$.
Proposition 5.11.
The algebra $K[W\oplus U_{\chi}]^{G}$ is minimally generated by $(x_{1}x_{2})^{2}$, $x_{1}^{6}+x_{2}^{6}$, $t^{2}$, $(x_{1}^{6}-x_{2}^{6})x_{1}x_{2}t$.
Proof.
The abelian subgroup $\langle a,b\rangle$ preserves the $1$-dimensional subspaces $Kx_{1}$, $Kx_{2}$, $Kt$ in $K[W\oplus U_{\chi}]$, hence $K[W\oplus U_{\chi}]^{\langle a,b\rangle}$ is generated by $\langle a,b\rangle$-invariant monomials (see Lemma 3.9). So $K[W\oplus U_{\chi}]^{\langle a,b\rangle}$ is generated by $x_{1}^{6}$, $x_{2}^{6}$, $(x_{1}x_{2})^{2}$, $t^{2}$, $x_{1}x_{2}t$. The element $c\in G$ fixes $(x_{1}x_{2})^{2}$ and $t^{2}$, interchanges $x_{1}^{6}$ and $x_{2}^{6}$, and $c\cdot(x_{1}x_{2}t)=-x_{1}x_{2}t$. It follows that $K[W\oplus U_{\chi}]^{G}$ is generated by $(x_{1}x_{2})^{2}$, $t^{2}$, $x_{1}^{6}+x_{2}^{6}$, $(x_{1}^{6}-x_{2}^{6})x_{1}x_{2}t$, and $(x_{1}^{6}-x_{2}^{6})^{2}$. The latter generator can be omitted, since we have $(x_{1}^{6}-x_{2}^{6})^{2}=(x_{1}^{6}+x_{2}^{6})^{2}-4(x_{1}^{2}x_{2}^{2})^{3}$. ∎
Theorem 5.12.
Assume that $K$ contains an element $\xi$ of multiplicative order $12$. Then we have $\beta_{\mathrm{sep}}^{K}((\mathrm{C}_{6}\times\mathrm{C}_{2})\rtimes\mathrm{C}_{2})=9$.
Proof.
Set $v:=([1,\xi]^{T},1)\in W\oplus U_{\chi}$ and $v^{\prime}:=([1,\xi]^{T},-1)\in W\oplus U_{\chi}$. The invariant $(x_{1}^{6}-x_{2}^{6})x_{1}x_{2}t$ has different values on $v$ and $v^{\prime}$. By Proposition 5.11 we see that all $G$-invariants of degree less than $9$ agree on $v$ and $v^{\prime}$. This shows that $\beta_{\mathrm{sep}}^{K}((\mathrm{C}_{6}\times\mathrm{C}_{2})\rtimes\mathrm{C}_{2})\geq 9$. On the other hand, $\beta^{K}((\mathrm{C}_{6}\times\mathrm{C}_{2})\rtimes\mathrm{C}_{2})=9$ by [15, Proposition 3.5], implying the reverse inequality $\beta_{\mathrm{sep}}^{K}((\mathrm{C}_{6}\times\mathrm{C}_{2})\rtimes\mathrm{C}_{2})\leq 9$. ∎
6. Some groups $G=H\times\mathrm{C}_{2}$ with $\beta_{\mathrm{sep}}^{K}(G)=\beta_{\mathrm{sep}}^{K}(H)$
6.1. The group $\mathrm{Dic}_{8}\times\mathrm{C}_{2}$
In this section
$G=\mathrm{Dic}_{8}\times\mathrm{C}_{2}=\langle a,b\mid a^{4}=1,\ b^{2}=a^{2},\ ba=a^{3}b\rangle\times\langle c\mid c^{2}=1\rangle$
is the direct product of the quaternion group of order $8$ and the group of order $2$. Assume that $K$ has an element $\mathrm{i}$ of multiplicative order $4$. Then $G$ has two irreducible $2$-dimensional representations (up to isomorphism), namely
$\psi_{1}:a\mapsto\begin{bmatrix}\mathrm{i}&0\\ 0&-\mathrm{i}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}0&-1\\ 1&0\\ \end{bmatrix},\quad c\mapsto\begin{bmatrix}1&0\\ 0&1\\ \end{bmatrix}$
(this is the only $2$ dimensional representation of $\mathrm{Dic}_{8}$ lifted to $G$ by the projection from $G$ to its direct factor $\mathrm{C}_{2}$), and
$\psi_{2}:a\mapsto\begin{bmatrix}\mathrm{i}&0\\ 0&-\mathrm{i}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}0&-1\\ 1&0\\ \end{bmatrix},\quad c\mapsto\begin{bmatrix}-1&0\\ 0&-1\\ \end{bmatrix},$
the representation $\psi_{1}$ tensored with the non-trivial representation $c\mapsto-1$ of $\mathrm{C}_{2}$. The other irreducible representations of $G$ are $1$-dimensional, and can be labeled by the elements of the subgroup $\widehat{G}:=\{\pm 1\}\times\{\pm 1\}\times\{\pm 1\}$ of $K^{\times}\times K^{\times}\times K^{\times}$ as follows: identify $\chi=(\chi_{1},\chi_{2},\chi_{3})\in\widehat{G}$ with the representation $\chi:G\to K^{\times}$ given by
$\chi:a\mapsto\chi_{1},\quad b\mapsto\chi_{2},\quad c\mapsto\chi_{3}.$
Denote by $W_{j}$ the vector space $K^{2}$ endowed with the representation $\psi_{j}$ for $j=1,2$, and for $\chi\in\widehat{G}$ denote by $U_{\chi}$ the vector space $K$ endowed with the representation $\chi$. Set
$U:=\bigoplus_{\chi\in\widehat{G}}U_{\chi}.$
The coordinate functions on $W_{1}$ are denoted by $x_{1},x_{2}$, the coordinate functions on $W_{2}$ are $y_{1},y_{2}$, and the coordinate function on $U_{\chi}$ is $t_{\chi}$ for $\chi\in\widehat{G}$ (see Section 3.6).
Proposition 6.1.
For $i=1,2$ we have $\beta(G,W_{i}\oplus U)=6$.
Proof.
First we deal with the case $i=1$. It is well known (see for example the proof of Proposition 4.2) that
(14) $K[W_{1}]^{G}=K[(x_{1}x_{2})^{2},\ x_{1}^{4}+x_{2}^{4},\ x_{1}x_{2}(x_{1}^{4}-x_{2}^{4})].$
It is easy to verify that a Gröbner basis with respect to the lexicographic term order induced by $x_{1}>x_{2}$ of the Hilbert ideal $\mathcal{H}(G,W_{1})$ is $x_{1}^{2}x_{2}^{2}$, $x_{1}^{4}+x_{2}^{4}$, $x_{2}^{6}$, $x_{1}x_{2}^{5}$. Using this Gröbner basis one can show that setting
$s_{(1,-1,1)}:=x_{1}x_{2},\quad s_{(-1,1,1)}:=x_{1}^{2}+x_{2}^{2},\quad s_{(-1,-1,1)}:=x_{1}^{2}-x_{2}^{2},$
a $G$-module direct complement in $K[x_{1},x_{2}]$ of the Hilbert ideal $\mathcal{H}(G,W_{1})$ is
$K\oplus\mathrm{Span}_{K}\{x_{1},x_{2}\}\oplus Ks_{(1,-1,1)}\oplus Ks_{(-1,1,1)}\oplus Ks_{(-1,-1,1)}\oplus\mathrm{Span}_{K}\{x_{1}^{3},x_{2}^{3}\}$
$\oplus\mathrm{Span}_{K}\{x_{1}^{2}x_{2},x_{1}x_{2}^{2}\}\oplus Ks_{(1,-1,1)}s_{(-1,1,1)}\oplus Ks_{(1,-1,1)}s_{(-1,-1,1)}$
$\oplus Ks_{(-1,1,1)}s_{(-1,-1,1)}\oplus\mathrm{Span}_{K}\{x_{1}^{5},x_{2}^{5}\}.$
The direct summands above are minimal $G$-invariant subspaces. Moreover, $G$ acts on $Ks_{\chi}$ via the character $\chi$. It follows by Lemma 3.9 that $K[W_{1}\oplus U]^{G}$ is generated over $K[W_{1}]^{G}$ by products of elements from $\{s_{\chi},\quad t_{\chi^{\prime}}\mid\chi\in\{(1,-1,1),(-1,1,1),(-1,-1,1)\},\ \chi^{\prime}\in\widehat{G}\}$, with at most two factors of the form $s_{\chi}$, and no $G$-invariant proper subproduct. Since the Davenport constant of the group $\widehat{G}$ is $4$, it is sufficient to consider products with at most $4$ factors. All of them have degree at most $6$.
The case $i=2$ is essentially the same. Indeed, the map $x_{1}\mapsto y_{1}$, $x_{2}\mapsto y_{2}$ extends to a $K$-algebra isomorphism between $K[W_{1}]$ and $K[W_{2}]$, which is also a $\mathrm{Dic}_{8}$-module isomorphism. The subgroup $\langle c\rangle$ acts trivially on $K[W_{1}]$, and acts by multiplication by $-1$ of the variables $y_{1},y_{2}$ in $K[W_{2}]$. Therefore replacing $x_{1}$ by $y_{1}$ and $x_{2}$ by $y_{2}$ in the corresponding formulae above, we get the generators of $K[W_{2}]^{G}$, as well as a direct sum decomposition into minimal $G$-invariant subspaces of a $G$-module direct complement in $K[W_{2}]$ of the Hilbert ideal. So in the same way as in the above paragraph, we get the conclusion $\beta(W_{2}\oplus U)=6$. ∎
Proposition 6.2.
We have $\beta(G,W_{1}\oplus W_{2})=6$.
Proof.
The element $a^{2}\in G$ multiplies each of the indeterminates $x_{1},x_{2},y_{1},y_{2}$ by $-1$. Since $G$-invariants are $a^{2}$-invariant, they must involve only monomials of even degree. Thus $\beta(G,W_{1}\oplus W_{2})$ is even. On the other hand, $\beta^{K}(G)=7$ by [15, Proposition 3.2]. We conclude that $\beta(G,W_{1}\oplus W_{2})\leq 6$. Note also that $\beta(G,W_{1}\oplus W_{2})\geq\beta(G,W_{1})$, and we saw in the course of the proof of Proposition 6.1 that $\beta(G,W_{1})=6$ (the third generator in (14) is not symmetric in $x_{1}$ and $x_{2}$, hence can not be expressed by the first two). ∎
Proposition 6.3.
Let $v:=(w_{1},w_{2})\in W_{1}\oplus W_{2}$ with $w_{1}\neq 0$ and $w_{2}\neq 0$. Then for any $\chi\in\widehat{G}$ there exists a relative invariant $f$ of weight $\chi$ such that $\deg(f)\leq 4$ and $f(v)\neq 0$.
Proof.
$\begin{array}[]{c|c}\chi&\text{relative invariants}\\ \hline\cr\hline\cr(-1,1,1)&f_{1}:=x_{1}^{2}+x_{2}^{2},\quad f_{2}:=x_{1}x_{2}(x_{1}^{2}-x_{2}^{2})\\ \hline\cr(-1,1,-1)&f_{1}:=x_{1}y_{1}+x_{2}y_{2},\quad f_{2}:=x_{1}x_{2}(x_{1}y_{1}-x_{2}y_{2}),\quad f_{3}:=x_{2}^{3}y_{1}-x_{1}^{3}y_{2}\\ \hline\cr(1,1,-1)&f_{1}:=x_{1}y_{2}-x_{2}y_{1},\quad f_{2}:=x_{1}x_{2}(x_{1}y_{2}+x_{2}y_{1}),\quad f_{3}:=x_{2}^{3}y_{2}+x_{1}^{3}y_{1}\end{array}$
A row in the above table contains a character $\chi\in\widehat{G}$ and relative invariants of weight $\chi$ of degree at most $4$, such that the common zero-set of these relative invariants is contained in the union of $W_{1}\oplus\{0\}$ and $W_{2}\oplus\{0\}$. For the other non-trivial weights the result can be deduced without further computation using Lemma 3.11. Indeed, the automorphism group of $G$ contains the automorphism group of its subgroup $\mathrm{Dic}_{8}$ as a subgroup. The automorphism group of $\mathrm{Dic}_{8}$ acts (on the right) therefore on $\widehat{G}$ (an automorphism $\alpha$ sends $\chi$ to $\chi\circ\alpha$), and any non-trivial $\chi\in\widehat{G}$ is in the orbit of one of the three weights in the above table. Moreover, for an automorphism $\alpha$ of $\mathrm{Dic}_{8}$ (viewed as an automorphism of $G$) we have that $\psi_{1}\circ\alpha\cong\psi_{1}$ and $\psi_{2}\circ\alpha\cong\psi_{2}$. Observe that for $(w_{1},w_{2})\in W_{1}\oplus W_{2}$, the condition that none of $w_{i}$ ($i=1,2)$ is zero is equivalent to the condition that $\{g\cdot(w_{1},w_{2})\mid g\in G\}$ spans $W_{1}\oplus W_{2}$ as a $K$-vector space. Therefore by Lemma 3.11 (iii), no $(w_{1},w_{2})$ with $w_{1}\neq 0$, $w_{2}\neq 0$ is contained in the common zero locus $\mathcal{V}(f\mid f\in K[W_{1}\oplus W_{2}]^{G,\chi\circ\alpha},\ \deg(f)\leq 4)$ if no such element of $W_{1}\oplus W_{2}$ is contained in $\mathcal{V}(f\mid f\in K[W_{1}\oplus W_{2}]^{G,\chi},\ \deg(f)\leq 4)$.
Note finally that for the trivial weight $(1,1,1)\in\widehat{G}$ we may take $f=1\in K[W_{1}\oplus W_{2}]^{G}$. ∎
Theorem 6.4.
Assume that $K$ contains an element of multiplicative order $8$. Then we have $\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{8}\times\mathrm{C}_{2})=6$.
Proof.
By Theorem 2.1 we know that $\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{8})=6$, and since $\mathrm{Dic}_{8}$ is a direct factor of $G$, we have $\beta_{\mathrm{sep}}^{K}(G)\geq\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{8})=6$.
Now we turn to the reverse inequality. By Lemma 3.1 (iii) it is sufficient to deal with the case when $|K|$ is large enough so that we can apply Lemma 3.6. Therefore it remains to prove that $\beta_{\mathrm{sep}}(G,V)\leq 6$ where $V=W_{1}\oplus W_{2}\oplus U$. That is, we have to show that if all invariants of degree at most $6$ take the same value on $v,v^{\prime}\in V$, then $G\cdot v=G\cdot v^{\prime}$. So $v=(w_{1},w_{2},u_{\chi}\mid\chi\in\widehat{G})$ and $v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2},u^{\prime}_{\chi}\mid\chi\in\widehat{G})$. It follows from Proposition 6.2 that replacing $v^{\prime}$ by an appropriate element in its $G$-orbit, we may assume that $w_{1}=w^{\prime}_{1}$ and $w_{2}=w^{\prime}_{2}$. If $w_{1}=0$ or $w_{2}=0$, then $v$ and $v^{\prime}$ belong to the same orbit by Proposition 6.1. For any $\chi\in\widehat{G}$ there exists a relative invariant $f\in K[W_{1}\oplus W_{2}]^{G,\chi^{-1}}$ with $\deg(f)\leq 4$ and $f(w_{1},w_{2})\neq 0$ by Proposition 6.3. Consider the invariant $ft_{\chi}\in K[V]^{G}$: it has degree at most $5$, and so from $(ft_{\chi})(v)=(ft_{\chi})(v^{\prime})$ we deduce $u_{\chi}=u^{\prime}_{\chi}$. This holds for all $\chi\in\widehat{G}$, thus we showed $v=v^{\prime}$, as claimed. ∎
6.2. The group $\mathrm{Dic}_{12}\times\mathrm{C}_{2}$.
In this section
$G=\mathrm{Dic}_{12}\times\mathrm{C}_{2}=\langle a,b\mid a^{6}=1,\ b^{2}=a^{3},\ ba=a^{-1}b\rangle\times\langle c\mid c^{2}=1\rangle.$
Assume that $K$ has an element $\xi$ of multiplicative order $12$; set $\omega:=\xi^{2}$ and $\mathrm{i}:=\xi^{3}$, so $\omega$ has multiplicative order $6$ and $\mathrm{i}$ has multiplicative order $4$. Consider the following irreducible $2$-dimensional representations of $G$:
$\psi_{1}:a\mapsto\begin{bmatrix}\omega&0\\ 0&\omega^{-1}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}0&1\\ -1&0\\ \end{bmatrix},\quad c\mapsto\begin{bmatrix}-1&0\\ 0&-1\\ \end{bmatrix}$
$\psi_{2}:a\mapsto\begin{bmatrix}\omega&0\\ 0&\omega^{-1}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}0&1\\ -1&0\\ \end{bmatrix},\quad c\mapsto\begin{bmatrix}1&0\\ 0&1\\ \end{bmatrix}$
$\psi_{3}:a\mapsto\begin{bmatrix}\omega^{2}&0\\ 0&\omega^{-2}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix},\quad c\mapsto\begin{bmatrix}-1&0\\ 0&-1\\ \end{bmatrix}$
$\psi_{4}:a\mapsto\begin{bmatrix}\omega^{2}&0\\ 0&\omega^{-2}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix},\quad c\mapsto\begin{bmatrix}1&0\\ 0&1\\ \end{bmatrix}.$
The other irreducible representations of $G$ are $1$-dimensional, and can be labelled by the group $\widehat{G}=\{\pm\mathrm{i},\pm 1\}\times\{\pm 1\}\leq K^{\times}\times K^{\times}$, where $\chi=(\chi_{1},\chi_{2})\in\widehat{G}$ is identified with the representation
$\chi:a\mapsto\chi_{1}^{2},\ b\mapsto\chi_{1},\ c\mapsto\chi_{2}$
(note that $\langle a^{2}\rangle$ is the commutator subgroup of $G$, so $a^{2}$ is in the kernel of any $1$-dimensional representation of $G$). For $j=1,2,3,4$ denote by $W_{j}$ the vector space $K^{2}$ endowed with the representation $\psi_{j}$, and for $\chi\in\widehat{G}$ denote by $U_{\chi}$ the vector space $K$ endowed with the representation $\chi$, and set
$U:=\bigoplus_{\chi\in\widehat{G}}U_{\chi}.$
The following result can be easily obtained using the CoCalc platform [43]:
Proposition 6.5.
Assume that $K$ has characteristic 00. Then we have the equality $\beta(G,V)=8$ if $V$ is any of the $G$-modules
$W_{1}\oplus W_{4}\oplus U,\ W_{2}\oplus W_{4}\oplus U,\ W_{3}\oplus W_{4}\oplus U,$
$\ W_{1}\oplus W_{2}\oplus W_{4},\ W_{1}\oplus W_{3}\oplus W_{4},\ W_{2}\oplus W_{3}\oplus W_{4}.$
Remark 6.6.
Proposition 6.5 implies in particular that $\beta(W_{i}\oplus W_{j})\leq 8$ for all $1\leq i<j\leq 3$. On the other hand, also using computer we found that $\beta(G,W_{1}\oplus W_{2}\oplus W_{3})=9$, and Proposition 6.5 can not yield immediately a better upper bound for $\beta_{\mathrm{sep}}(G,W_{1}\oplus W_{2}\oplus W_{3})$. Indeed, consider $v=(w_{1},w_{2},w_{3}):=([1,0]^{T},[1,0]^{T},[1,1]^{T})\in W_{1}\oplus W_{2}\oplus W_{3}$ and $v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2},w^{\prime}_{3}):=([1,0]^{T},[1,0]^{T},[-1,-1]^{T})\in W_{1}\oplus W_{2}\oplus W_{3}$. Then $(w_{1},w_{2})=(w^{\prime}_{1},w^{\prime}_{2})$, $b^{2}c\cdot(w_{1},w_{3})=(w^{\prime}_{1},w^{\prime}_{3})$, $c\cdot(w_{2},w_{3})=(w^{\prime}_{2},w^{\prime}_{3})$. So all pairs $(w_{i},w_{j})$ and $(w^{\prime}_{i},w^{\prime}_{j})$ have the same $G$-orbit. However, $v$ and $v^{\prime}$ have different $G$-orbits, because the $G$-invariant $x_{2}y_{2}z_{1}+x_{1}y_{1}z_{2}$ separates them.
Proposition 6.7.
Assume that $K$ has characteristic 00. For $V:=W_{1}\oplus W_{2}\oplus W_{3}$ we have the inequality $\beta_{\mathrm{sep}}(G,V)\leq 8$.
Proof.
Assume that for $v=(w_{1},w_{2},w_{3})\in V$, $v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2},w^{\prime}_{3})\in V$ we have that $f(v)=f(v^{\prime})$ for all $f\in K[V]^{G}$ with $\deg(f)\leq 8$. We need to show that $G\cdot v=G\cdot v^{\prime}$. By Proposition 6.5 we know that $\beta(G,W_{i}\oplus W_{j})\leq 8$ for all $i,j\in\{1,2,3,4\}$. In particular, $G\cdot(w_{1},w_{2})=G\cdot(w^{\prime}_{1},w^{\prime}_{2})$. So replacing $v^{\prime}$ by an appropriate element in its orbit we may assume that $w^{\prime}_{1}=w_{1}$, $w^{\prime}_{2}=w_{2}$, and moreover, both $w_{1}$ and $w_{2}$ are non-zero. We shall show that necessarily $w_{3}=w^{\prime}_{3}$. Consider the following $G$-invariants in $K[V]^{G}$:
$f_{1}:=x_{2}y_{2}z_{1}+x_{1}y_{1}z_{2},\quad f_{2}:=y_{1}y_{2}(x_{2}y_{2}z_{1}-x_{1}y_{1}z_{2}),\quad f_{3}:=x_{1}^{3}y_{1}z_{1}+x_{2}^{3}y_{2}z_{2},$
$\ f_{4}:=x_{2}^{3}y_{1}z_{1}-x_{1}^{3}y_{2}z_{2},\qquad f_{5}:=x_{1}y_{2}^{3}z_{1}-x_{2}y_{1}^{3}z_{2}.$
For $w\in V$ consider the matrices
$M_{1}(w)$ $:=\begin{bmatrix}x_{2}(w)y_{2}(w)&x_{1}(w)y_{1}(w)\\ x_{2}(w)y_{1}(w)y_{2}^{2}(w)&-x_{1}(w)y_{1}^{2}(w)y_{2}(w)\end{bmatrix}$
$M_{2}(w)$ $:=\begin{bmatrix}x_{2}(w)y_{2}(w)&x_{1}(w)y_{1}(w)\\ x_{1}^{3}(w)y_{1}(w)&x_{2}^{3}(w)y_{2}(w)\end{bmatrix}$
$M_{3}(w)$ $:=\begin{bmatrix}x_{2}^{3}(w)y_{1}(w)&-x_{1}^{3}(w)y_{2}(w)\\ x_{1}(w)y_{2}^{3}(w)&-x_{2}(w)y_{1}^{3}(w)\end{bmatrix}.$
The definition of the $f_{j}$ implies that for any $w\in V$ we have the matrix equalities
$M_{1}(w)\cdot\begin{bmatrix}z_{1}(w)\\ z_{2}(w)\end{bmatrix}$ $=\begin{bmatrix}f_{1}(w)\\ f_{2}(w)\end{bmatrix}$
$M_{2}(w)\cdot\begin{bmatrix}z_{1}(w)\\ z_{2}(w)\end{bmatrix}$ $=\begin{bmatrix}f_{1}(w)\\ f_{3}(w)\end{bmatrix}$
$M_{3}(w)\cdot\begin{bmatrix}z_{1}(w)\\ z_{2}(w)\end{bmatrix}$ $=\begin{bmatrix}f_{4}(w)\\ f_{5}(w)\end{bmatrix}$
Note that $M_{j}(v)=M_{j}(v^{\prime})$ for $j=1,2,3$ (because $(w_{1},w_{2})=(w^{\prime}_{1},w^{\prime}_{2})$), and since $\deg(f_{j})\leq 5$, by assumption, we have $f_{j}(v)=f_{j}(v^{\prime})$ for $j=1,2,3,4,5$. By basic linear algebra we conclude that $\begin{bmatrix}z_{1}(v)\\ z_{2}(v)\end{bmatrix}=\begin{bmatrix}z_{1}(v^{\prime})\\ z_{2}(v^{\prime})\end{bmatrix}$, unless the matrix $M_{j}(v)$ has zero determinant for all $j\in\{1,2,3\}$. We claim that this is not the case. Suppose to the contrary that $\det M_{j}(v)=0$ for $j=1,2,3$. Then $\det M_{1}(v)=0$ says that one of $x_{1}(v)$, $x_{2}(v)$, $y_{1}(v)$, $y_{2}(v)$ equals zero. Suppose for example that $x_{1}(v)=0$. Then $x_{2}(v)\neq 0$ (as $w_{2}\neq 0$), and $\det M_{2}(v)=0$ yields $y_{2}(v)=0$, implying in turn that $y_{1}(v)\neq 0$. Then $\det M_{3}(v)\neq 0$, a contradiction. The cases when $x_{2}(v)$, $y_{1}(v)$, or $y_{2}(v)$ is zero can be dealt with similarly. ∎
Proposition 6.8.
Assume that $K$ has characteristic 00. Let $i,j\in\{1,2,3\}$, $i\neq j$, and $\chi\in\widehat{G}$. Then $\beta_{\mathrm{sep}}(G,W_{i}\oplus W_{j}\oplus U_{\chi})\leq 8$.
Proof.
Using the CoCalc platform [43] we verified that for a $G$-module of the form $V=W_{i}\oplus W_{j}\oplus U_{\chi}$ we have $\beta(G,V)>8$ if and only if $V$ is as in the table below:
$\begin{array}[]{c|c}V&tf_{1},tf_{2},tf_{3}\in K[V]^{G}\\ \hline\cr W_{2}\oplus W_{3}\oplus U_{(\mathrm{i},-1)}&(y_{1}z_{1}+\mathrm{i}y_{2}z_{2})t,\ y_{1}y_{2}(y_{1}z_{1}-\mathrm{i}y_{2}z_{2})t,\ (y_{2}z_{1}^{5}-\mathrm{i}y_{1}z_{2}^{5})t\\ W_{2}\oplus W_{3}\oplus U_{(-\mathrm{i},-1)}&(y_{1}z_{1}-\mathrm{i}y_{2}z_{2})t,\ y_{1}y_{2}(y_{1}z_{1}+\mathrm{i}y_{2}z_{2})t,\ (y_{2}z_{1}^{5}+\mathrm{i}y_{1}z_{2}^{5})t\\ W_{1}\oplus W_{3}\oplus U_{(\mathrm{i},1)}&(x_{1}z_{1}+\mathrm{i}x_{2}z_{2})t,\ x_{1}x_{2}(x_{1}z_{1}-\mathrm{i}x_{2}z_{2})t,\ (x_{2}z_{1}^{5}-\mathrm{i}x_{1}z_{2}^{5})t\\ W_{1}\oplus W_{3}\oplus U_{(-\mathrm{i},1)}&(x_{1}z_{1}-\mathrm{i}x_{2}z_{2})t,\ x_{1}x_{2}(x_{1}z_{1}+\mathrm{i}x_{2}z_{2})t,\ (x_{2}z_{1}^{5}+\mathrm{i}x_{1}z_{2}^{5})t\\ W_{1}\oplus W_{2}\oplus U_{(-1,-1)}&(x_{2}y_{1}+x_{1}y_{2})t,\ x_{1}x_{2}(x_{2}y_{1}-x_{1}y_{2})t,\ (x_{1}y_{1}^{5}-x_{2}y_{2}^{5})t\\ W_{1}\oplus W_{2}\oplus U_{(1,-1)}&(x_{2}y_{1}-x_{1}y_{2})t,\ x_{1}x_{2}(x_{2}y_{1}+x_{1}y_{2})t,\ (x_{1}y_{1}^{5}+x_{2}y_{2}^{5})t\end{array}$
We shall show that $\beta_{\mathrm{sep}}(G,V)\leq 8$ for $V=W_{2}\oplus W_{3}\oplus U_{(\mathrm{i},-1)}$; the argument for the other $V$ in the table above is the same. Take $v=(w_{2},w_{3},u)\in V$ and $v^{\prime}=(w^{\prime}_{2},w^{\prime}_{3},u^{\prime})\in V$, and assume that $f(v)=f(v^{\prime})$ holds for all $f\in K[V]^{G}$ with $\deg(f)\leq 8$. We need to show that then $G\cdot v=G\cdot v^{\prime}$. Recall that the coordinate functions on $W_{2}$ are denoted by $y_{1},y_{2}$, on $W_{3}$ by $z_{1},z_{2}$, and on $U_{(\mathrm{i},-1)}$ by $t$. By Proposition 6.5 we know that if $w_{2}=0$, then $w^{\prime}_{2}=0$, and $(w_{3},u)$ and $(w^{\prime}_{3},u^{\prime})$ have the same $G$-orbit, Similarly, if $w_{3}=0$, then $G\cdot v=G\cdot v^{\prime}$. So we may assume that $w_{2}\neq 0$ and $w_{3}\neq 0$, and again by Proposition 6.5, by replacing $v^{\prime}$ by an appropriate element in its $G$-orbit, we may assume that $w_{2}=w^{\prime}_{2}$ and $w_{3}=w^{\prime}_{3}$. It remains to show that $u=u^{\prime}$. Denote by $tf_{1}$, $tf_{2}$, $tf_{3}$ the $G$-invariants on $V$ given in the table. All have degree less than $8$, so $(tf_{j})(v)=(tf_{j})(v^{\prime})$ holds for $j=1,2,3$. One can easily deduce from $w_{2}\neq 0$ and $w_{3}\neq 0$ that $f_{j}(w_{2},w_{3})\neq 0$ for some $j\in\{1,2,3\}$, hence $t(v)=t(v^{\prime})$, i.e. $u=u^{\prime}$. ∎
Theorem 6.9.
Assume that $K$ has characteristic 00 and it has an element of multiplictive order $12$. Then we have the equality $\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{12}\times\mathrm{C}_{2})=8$.
Proof.
Since $\mathrm{Dic}_{12}$ is a homomorphic image of $G$, we have the obvious inequality $\beta_{\mathrm{sep}}^{K}(G)\geq\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{12})$, and by Theorem 2.1 we have $\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{12})=8$. By Lemma 3.6 it remains to prove that $\beta_{\mathrm{sep}}(G,V)\leq 8$, where $V$ is the $G$-module $V:=W_{1}\oplus W_{2}\oplus W_{3}\oplus W_{4}\oplus U$. Let $v=(w_{1},w_{2},w_{3},w_{4},u)\in V$, $v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2},w^{\prime}_{3},w^{\prime}_{4},u^{\prime})\in V$, and assume that
(15) $f(v)=f(v^{\prime})\text{ for all }f\in K[V]^{G}\text{ with }\deg(f)\leq 8.$
We shall show that $v$ and $v^{\prime}$ have the same $G$-orbit.
Case 1: There exists some $i,j\in\{1,2,3\}$, $i\neq j$, such that $w_{i}\neq 0$ and $w_{j}\neq 0$.
If $w_{1}\neq 0$ then $\mathrm{Stab}_{G}(w_{1})=\ker(\psi_{1})=\langle b^{2}c\rangle$ has order $2$. If $w_{2}\neq 0$ then $\mathrm{Stab}_{G}(w_{2})=\ker(\psi_{2})=\langle c\rangle$ has order $2$. Since $\langle b^{2}c\rangle\cap\langle c\rangle=\{1_{G}\}$ and no non-zero element of $W_{3}$ is fixed by $b^{2}c$ or $c$, we have that $\mathrm{Stab}_{G}(w_{i},w_{j})$ is trivial. By Proposition 6.5 we know that $G\cdot(w_{i},w_{j})=G\cdot(w^{\prime}_{i},w^{\prime}_{j})$, so replacing $v^{\prime}$ by an appropriate element in its $G$-orbit, we may assume that $w^{\prime}_{i}=w_{i}$ and $w^{\prime}_{j}=w_{j}$. Take a component $w\in\{w_{1},w_{2},w_{3},w_{4},u_{\chi}\mid\chi\in\widehat{G}\}$ of $v$ different from $w_{i}$ and $w_{j}$, and let $w^{\prime}$ be the corresponding component of $v^{\prime}$. By Proposition 6.5, Proposition 6.7 and Proposition 6.8 we conclude that $G\cdot(w_{i},w_{j},w)=G\cdot(w_{i},w_{j},w^{\prime})$, so $w^{\prime}$ is in the orbit of $w$ with respect to the stabilizer of $(w_{i},w_{j})$, implying in turn that $w=w^{\prime}$. Since this holds for all components $w$ and $w^{\prime}$ of $v$ and $v^{\prime}$, we have $v=v^{\prime}$.
Case 2: At most one of $w_{1},w_{2},w_{3}$ is non-zero, so there exists an $i\in\{1,2,3\}$ such that $w_{j}=0$ for each $j\in\{1,2,3\}\setminus\{i\}$. As $\beta_{\mathrm{sep}}(W_{j})\leq 8$ for all $j\in\{1,2,3\}\setminus\{i\}$, we conclude that also $w^{\prime}_{j}=0$ for all $j\in\{1,2,3\}\setminus\{i\}$. Thus $v$ and $v^{\prime}$ belong to a submodule $X$ of $V$ for which $\beta(G,X)\leq 8$ by Proposition 6.5, and hence (15) implies $G\cdot v=G\cdot v^{\prime}$. ∎
6.3. The group $\mathrm{A}_{4}\times\mathrm{C}_{2}$.
In this section
$G=\mathrm{A}_{4}\times\mathrm{C}_{2}=\langle a,b,c\mid a^{2}=b^{2}=c^{3}=1,\ ab=ba,\ cac^{-1}=b,\ cbc^{-1}=ab\rangle\times\langle d\mid d^{2}=1\rangle.$
Consider the following irreducible $3$-dimensional representations of $G$:
$\psi_{1}:a\mapsto\begin{bmatrix}-1&0&0\\ 0&-1&0\\ 0&0&1\end{bmatrix},\quad b\mapsto\begin{bmatrix}1&0&0\\ 0&-1&0\\ 0&0&-1\end{bmatrix},\quad c\mapsto\begin{bmatrix}0&0&1\\ 1&0&0\\ 0&1&0\end{bmatrix},\quad d\mapsto\begin{bmatrix}-1&0&0\\ 0&-1&0\\ 0&0&-1\end{bmatrix}$
$\psi_{2}:a\mapsto\begin{bmatrix}-1&0&0\\ 0&-1&0\\ 0&0&1\end{bmatrix},\quad b\mapsto\begin{bmatrix}1&0&0\\ 0&-1&0\\ 0&0&-1\end{bmatrix},\quad c\mapsto\begin{bmatrix}0&0&1\\ 1&0&0\\ 0&1&0\end{bmatrix},\quad d\mapsto\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}.$
For $j=1,2$ denote by $W_{j}$ the vector space $K^{3}$ endowed with the representation $\psi_{j}$. Assume in addition that $K$ contains an element $\omega$ of multiplicative order $3$. Then the other irreducible representations of $G$ are $1$-dimensional, and can be labelled by the group $\widehat{G}=\{\omega,\omega^{2},1\}\times\{\pm 1\}\leq K^{\times}\times K^{\times}$, where $\chi=(\chi_{1},\chi_{2})\in\widehat{G}$ is identified with the representation
$\chi:a\mapsto 1,\ b\mapsto 1,\ c\mapsto\chi_{1},\ d\mapsto\chi_{2}$
(note that $\langle a,b\rangle$ is the commutator subgroup of $G$). For $\chi\in\widehat{G}$ denote by $U_{\chi}$ the vector space $K$ endowed with the representation $\chi$, and set $U:=\bigoplus_{\chi\in\widehat{G}}U_{\chi}$.
Proposition 6.10.
We have the equalities
  • (i)
    $\beta(G,W_{1})=6$.
  • (ii)
    $\beta(G,W_{2}\oplus U)=6$.
Proof.
(i) The algebra $K[W_{1}]^{\langle a,b,d\rangle}$ is generated by the monomials $x_{1}^{2}$, $x_{2}^{2}$, $x_{3}^{2}$. The group element $c$ permutes cyclically these monomials, hence the algebra $K[W_{1}]^{G}$ is generated by $x_{1}^{2}+x_{2}^{2}+x_{3}^{2}$, $x_{1}^{2}x_{2}^{2}+x_{2}^{2}x_{3}^{2}+x_{1}^{2}x_{3}^{2}$, $x_{1}^{2}x_{2}^{2}x_{3}^{2}$, $x_{1}^{4}x_{2}^{2}+x_{2}^{4}x_{3}^{2}+x_{1}^{2}x_{3}^{4}$, and since the elements with degree less than $6$ are symmetric, they can not form a generating set. Therefore we have (i).
(ii) By a similar reasoning, $K[W_{2}]^{G}$ is generated by $y_{1}^{2}+y_{2}^{2}+y_{3}^{2}$, $y_{1}y_{2}y_{3}$, $y_{1}^{2}y_{2}^{2}+y_{2}^{2}y_{3}^{2}+y_{1}^{2}y_{3}^{2}$, $y_{1}^{4}y_{2}^{2}+y_{2}^{4}y_{3}^{2}+y_{3}^{4}y_{1}^{2}$, hence $\beta(G,W_{2})=6$. The factor group $G/G^{\prime}=G/\langle a,b\rangle$ is cyclic of order $6$, hence its Davenport constant is $6$, implying that $\beta(G,U)=6$. Following the notation of Lemma 3.9, write $A_{\chi}$ for a set of homogeneous elements in $K[W_{2}]^{G,\chi}$ that span a vector space direct complement in $K[W_{2}]^{G,\chi}$ of $K[W_{2}]^{G,\chi}\cap\mathcal{H}(G,W_{2})$. By Lemma 3.9 it is sufficient to show that (for some choice of the sets $A_{\chi}$) the polynomials in the following set have degree at most $6$:
$C:=\{ht_{\chi^{(1)}}\cdots t_{\chi^{(k)}}\mid\chi^{(1)},\dots\chi^{(k)}$ is a product-one free sequence over $\widehat{G}$,
$h\in A_{\chi}$ where $\chi^{-1}=\chi^{(1)}\cdots\chi^{(k)}\}.$
As $d\in\ker(\psi_{2})$, the space $K[W_{2}]^{G,\chi}$ is non-zero only if $\chi_{2}=1$. Moreover, $a,b\in\ker(\chi)$ for any $\chi\in\widehat{G}$, hence $K[W_{2}]^{G,\chi}$ is contained in $K[W_{2}]^{\langle a,b\rangle}$. It is easy to see that $K[W_{2}]^{\langle a,b\rangle}$ is generated by $y_{1}^{2},y_{2}^{2},y_{3}^{2},y_{1}y_{2}y_{3}$, and $K[W_{2}]^{G}=K[y_{1}^{2},y_{2}^{2},y_{3}^{2},y_{1}y_{2}y_{3}]^{\langle c\rangle}$ is generated by $y_{1}^{2}+y_{2}^{2}+y_{3}^{2}$, $y_{1}^{2}y_{2}^{2}+y_{2}^{2}y_{3}^{2}+y_{1}^{2}y_{3}^{2}$, $y_{1}^{4}y_{2}^{2}+y_{2}^{4}y_{3}^{2}+y_{1}^{2}y_{3}^{4}$, $y_{1}y_{2}y_{3}$. Set
$s_{\omega}:=y_{1}^{2}+\omega y_{2}^{2}+\omega^{2}y_{3}^{2},\quad s_{\omega^{2}}:=y_{1}^{2}+\omega^{2}y_{2}^{2}+\omega y_{3}^{2}.$
We may take $A_{(\omega,1)}:=\{s_{\omega},s^{2}_{\omega^{2}}\}$ and $A_{(\omega^{2},1)}:=\{s_{\omega^{2}},s^{2}_{\omega}\}$ (this can be verified by computing a Gröbner basis in $\mathcal{H}(G,W_{2})$). Now consider an element $f=ht_{\chi^{(1)}}\cdots t_{\chi^{(k)}}\in C$. Here $\deg(h)=2$, or $\deg(h)=4$ and $h$ is the product of two $\langle a,b,d\rangle$-invariants. After a possible renumbering we may assume that $\chi^{(i)}_{2}=-1$ for each $i=1,...,\ell$ and $\chi^{(j)}_{2}=1$ for each $j=\ell+1,...,k$. Since $h$ is a $\langle d\rangle$-invariant, $\ell=2\ell^{\prime}$ must be even. We have $f=h\prod_{i=1}^{\ell^{\prime}}(t_{\chi^{(i)}}t_{\chi^{(i+\ell^{\prime})}})\prod_{j=\ell+1}^{k}t_{\chi^{(j)}}$. So the $G$-invariant $f$ is written as a product of relative invariants of weight $\chi\in\langle(\omega,1)\rangle\cong\mathrm{C}_{3}$, where the number of factors is $k-\ell^{\prime}+1$ or $k-\ell^{\prime}+2$, depending on whether $\deg(h)=2$ or $\deg(h)=4$. On the other hand, the number of factors is at most $\mathsf{D}(\mathrm{C}_{3})=3$, since $\chi^{(1)},\dots\chi^{(k)}$ is a product-one free sequence over $\widehat{G}$. Since $\deg(f)=\deg(h)+k$, we conclude in both cases that $\deg(f)\leq 6$, and we are done. ∎
Lemma 6.11.
For a non-zero $v\in W_{1}$ we have
$|\mathrm{Stab}_{G}(v)|=\begin{cases}3&\text{ if }x_{1}^{2}(v)=x_{2}^{2}(v)=x_{3}^{2}(v)\\ 2&\text{ if }x_{j}(v)=0\text{ for a unique }j\in\{1,2,3\}\\ 4&\text{ if }x_{j}(v)\neq 0\text{ for a unique }j\in\{1,2,3\}\\ 1&\text{ otherwise.}\end{cases}.$
Proof.
One can easily find the non-zero elements of $W_{1}$ that occur as an eigenvector with eigenvalue $1$ for some non-identity element of $G$. ∎
Proposition 6.12.
We have the inequality $\beta_{\mathrm{sep}}(G,W_{1}\oplus W_{2})\leq 6$.
Proof.
Take $v=(w_{1},w_{2}),\ v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2})\in W_{1}\oplus W_{2}$ such that $f(v)=f(v^{\prime})$ for all $f\in K[W_{1}\oplus W_{2}]^{G}$ with $\deg(f)\leq 6$. We need to show that $G\cdot v=G\cdot v^{\prime}$. By Proposition 6.10, $w_{1}$ and $w^{\prime}_{1}$ have the same $G$-orbit, so we assume that $w_{1}=w^{\prime}_{1}$ (i.e. $x_{1}(v)=x_{1}(v^{\prime})$, $x_{2}(v)=x_{2}(v^{\prime})$, $x_{3}(v)=x_{3}(v^{\prime})$). Moreover, we may assume that $G\cdot w_{2}=G\cdot w^{\prime}_{2}$, and it is sufficient to deal with the case when both $w_{1}$ and $w_{2}$ are non-zero. By Lemma 6.11, the stabilizer of $w_{1}$ in $G$ is non-trivial if and only if $(x_{1}x_{2}x_{3})(v)=0$ or $x_{1}^{2}(v)=x_{2}^{2}(v)=x_{3}^{2}(v)$. This dictates the distinction of several cases below.
Case I: $(x_{1}x_{2}x_{3})(v)\neq 0$ and $x_{i}^{2}(v)\neq x_{j}^{2}(v)$ if $i\neq j$. Consider the $G$-invariants
$f_{1}:=y_{1}^{2}+y_{2}^{2}+y_{3}^{2},\qquad f_{2}:=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{3}^{2}y_{3}^{2},\qquad f_{3}:=x_{1}^{4}y_{1}^{2}+x_{2}^{4}y_{2}^{2}+x_{3}^{4}y_{3}^{2}.$
For $w\in V$ set
$M(w):=\begin{bmatrix}1&1&1\\ x_{1}^{2}(w)&x_{2}^{2}(w)&x_{3}^{2}(w)\\ x_{1}^{4}(w)&x_{2}^{4}(w)&x_{3}^{4}(w)\end{bmatrix}.$
We have $\det M(w)=(x_{1}^{2}(w)-x_{2}^{2}(w))(x_{1}^{2}(w)-x_{3}^{2}(w))(x_{2}^{2}(w)-x_{3}^{2}(w))$, so $\det M(v)\neq 0$ by assumption, and as $M(v)$ depends only on $w_{1}=w^{\prime}_{1}$, we have $M(v)=M(v^{\prime})$. Using that $f_{i}(v)=f_{i}(v^{\prime})$ for $i=1,2,3$ (since $\deg(f_{i})\leq 6$), we conclude
$\begin{bmatrix}y_{1}^{2}(v)\\ y_{2}^{2}(v)\\ y_{3}^{2}(v)\end{bmatrix}=M(v)^{-1}\cdot\begin{bmatrix}f_{1}(v)\\ f_{2}(v)\\ f_{3}(v)\end{bmatrix}=M(v^{\prime})^{-1}\cdot\begin{bmatrix}f_{1}(v^{\prime})\\ f_{2}(v^{\prime})\\ f_{3}(v^{\prime})\end{bmatrix}=\begin{bmatrix}y_{1}^{2}(v^{\prime})\\ y_{2}^{2}(v^{\prime})\\ y_{3}^{2}(v^{\prime})\end{bmatrix}.$
Thus $y_{i}(v)=\pm y_{i}(v^{\prime})$ for $i=1,2,3$. Taking into account that $G\cdot w_{2}=G\cdot w^{\prime}_{2}$, one can easily see that either $y_{i}(v)=y_{i}(v^{\prime})$ for $i=1,2,3$, so $v=v^{\prime}$, and we are done, or for some $j\in\{1,2,3\}$ we have that $y_{j}(v)=y_{j}(v^{\prime})$ and $y_{i}(v)=-y_{i}(v^{\prime})$ for all $i\in\{1,2,3\}\setminus\{j\}$. By symmetry we may assume that
(16) $y_{3}(v)=y_{3}(v^{\prime}),\qquad y_{1}(v)=-y_{1}(v^{\prime}),\qquad y_{2}(v)=-y_{2}(v^{\prime}).$
Consider the $G$-invariants
$f_{4}:=x_{2}x_{3}y_{1}+x_{1}x_{3}y_{2}+x_{1}x_{2}y_{3},\qquad f_{5}:=x_{1}x_{2}x_{3}(x_{1}y_{1}+x_{2}y_{2}+x_{2}y_{3}).$
Then $f_{4}$ and $f_{5}$ have degree less than $6$, hence $f_{4}(v)=f_{4}(v^{\prime})$ and $f_{5}(v)=f_{5}(v^{\prime})$, implying by $y_{3}(v)=y_{3}(v^{\prime})$ (see (16)) that
$\begin{bmatrix}(x_{2}x_{3})(v)&(x_{1}x_{3})(v)\\ x_{1}(v)&x_{2}(v)\end{bmatrix}\cdot\begin{bmatrix}y_{1}(v)\\ y_{2}(v)\end{bmatrix}=\begin{bmatrix}(x_{2}x_{3})(v)&(x_{1}x_{3})(v)\\ x_{1}(v)&x_{2}(v)\end{bmatrix}\cdot\begin{bmatrix}y_{1}(v^{\prime})\\ y_{2}(v^{\prime})\end{bmatrix}.$
The assumptions for Case I guarantee that the determinant of the $2\times 2$ coefficient matrix above is non-zero, hence $\begin{bmatrix}y_{1}(v)\\ y_{2}(v)\end{bmatrix}=\begin{bmatrix}y_{1}(v^{\prime})\\ y_{2}(v^{\prime})\end{bmatrix}$, and so $v=v^{\prime}$.
Case II: $(x_{1}x_{2}x_{3})(v)\neq 0$ and |$|\{x_{1}^{2}(v),x_{2}^{2}(v),x_{3}^{2}(v)\}|=2$, say $x_{1}^{2}(v)=x_{2}^{2}(v)\neq x_{3}^{2}(v)$. Similarly to Case I, by basic linear algebra we conclude from $f_{1}(v)=f_{1}(v^{\prime})$ and $f_{2}(v)=f_{2}(v^{\prime})$ that
(17) $(y_{1}^{2}+y_{2}^{2})(v)=(y_{1}^{2}+y_{2}^{2})(v^{\prime})\text{ and }y_{3}^{2}(v)=y_{3}^{2}(v^{\prime}).$
Consider the $G$-invariant
$f_{6}:=x_{3}^{2}y_{1}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{3}^{2}.$
Now (17), $f_{6}(v)=f_{6}(v^{\prime})$ imply that $y_{1}^{2}(v)=y_{1}^{2}(v^{\prime})$ and $y_{2}^{2}(v)=y_{2}^{2}(v^{\prime})$. So we have that
$y_{1}(v)=\pm y_{1}(v^{\prime}),\quad y_{2}(v)=\pm y_{2}(v^{\prime}),\quad y_{3}(v)=\pm y_{3}(v^{\prime}).$
Taking into account that $G\cdot w_{2}=G\cdot w^{\prime}_{2}$ we conclude that either $w_{2}=w^{\prime}_{2}$, and we are done, or for some $j\in\{1,2,3\}$ we have that $y_{j}(v)=y_{j}(v^{\prime})$ and $y_{i}(v)=-y_{i}(v^{\prime})$ for all $i\in\{1,2,3\}\setminus\{j\}$. So we have to deal with the cases II.a, II.b, II.c below:
Case II.a: $y_{1}(v)=-y_{1}(v^{\prime})$, $y_{2}(v)=-y_{2}(v^{\prime})$, $y_{3}(v)=y_{3}(v^{\prime})$. Consider the invariant
$f_{7}:=x_{2}x_{3}^{3}y_{1}+x_{1}^{3}x_{3}y_{2}+x_{1}x_{2}^{3}y_{3}.$
It has degree less than $6$, and from $f_{4}(v)=f_{4}(v^{\prime})$, $f_{7}(v)=f_{7}(v^{\prime})$ we conclude
$\begin{bmatrix}(x_{2}x_{3})(v)&(x_{1}x_{3})(v)\\ (x_{2}x_{3}^{3})(v)&(x_{1}^{3}x_{3})(v)\end{bmatrix}\cdot\begin{bmatrix}y_{1}(v)\\ y_{2}(v)\end{bmatrix}=\begin{bmatrix}0\\ 0\end{bmatrix}.$
The determinant of the $2\times 2$ matrix above is $x_{1}(v)x_{2}(v)x_{3}(v)^{2}(x_{1}(v)^{2}-x_{3}(v)^{2})$, which is non-zero by the assumptions for Case II. Consequently, $y_{1}(v)=y_{2}(v)=0$, implying in turn that $v=v^{\prime}$.
Case II.b: $y_{1}(v)=-y_{1}(v^{\prime})$, $y_{2}(v)=y_{2}(v^{\prime})$, $y_{3}(v)=-y_{3}(v^{\prime})$. Similar to Case II.a, using invariants $f_{4}$ and $f_{7}$.
Case II.c: $y_{1}(v)=y_{1}(v^{\prime})$, $y_{2}(v)=-y_{2}(v^{\prime})$, $y_{3}(v)=-y_{3}(v^{\prime})$. Similar to Case II.a, but instead of $f_{7}$ we have to use the invariant $x_{2}^{3}x_{3}y_{1}+x_{1}x_{3}^{3}y_{2}+x_{1}^{3}x_{2}y_{3}$.
Case III.a: Two of $x_{1}(v)$, $x_{2}(v)$, $x_{3}(v)$ are zero, say $x_{1}(v)=x_{2}(v)=0$, $x_{3}(v)\neq 0$. So $\mathrm{Stab}_{G}(w_{1})$ has order $4$ by Lemma 6.11, in fact $\mathrm{Stab}_{G}(w_{1})=\langle a,bd\rangle$. Then $f_{2}(v)=f_{2}(v^{\prime})$ implies $y_{3}(v)^{2}=y_{3}(v^{\prime})^{2}$, $f_{6}(v)=f_{6}(v^{\prime})$ implies $y_{1}(v)^{2}=y_{1}(v^{\prime})^{2}$, and then $f_{1}(v)=f_{1}(v^{\prime})$ implies $y_{2}(v)^{2}=y_{2}(v^{\prime})^{2}$. Taking into account that $G\cdot w_{2}=G\cdot w^{\prime}_{2}$ we conclude that either $w_{2}=w^{\prime}_{2}$, and we are done, or for some $j\in\{1,2,3\}$ we have that $y_{j}(v)=y_{j}(v^{\prime})$ and $y_{i}(v)=-y_{i}(v^{\prime})$ for all $i\in\{1,2,3\}\setminus\{j\}$. In the latter case we have $w_{2}$ and $w^{\prime}_{2}$ belong to the same orbit under $\langle a,bd\rangle=\mathrm{Stab}_{G}(w_{1})$, hence $G\cdot v=G\cdot v^{\prime}$.
Case III.b: $x_{i}(v)=0$ for a unique $i\in\{1,2,3\}$, say $x_{3}(v)=0$. Then $f_{4}(v)=f_{4}(v^{\prime})$ implies $y_{3}(v)=y_{3}(v^{\prime})$. The $G$-invariant
$f_{8}:=x_{1}^{2}x_{3}^{2}y_{1}^{2}+x_{1}^{2}x_{2}^{2}y_{2}^{2}+x_{2}^{2}x_{3}^{2}y_{3}^{2}$
shows that $y_{2}^{2}(v)=y_{2}^{2}(v^{\prime})$, hence by $f_{1}(v)=f_{1}(v^{\prime})$ we get $y_{1}^{2}(v)=y_{1}^{2}(v^{\prime})$. Recall that $G\cdot w_{2}=G\cdot w^{\prime}_{2}$, hence either $v=v^{\prime}$ and we are done, or $y_{1}(v)=-y_{1}(v^{\prime})$ and $y_{2}(v)=-y_{2}(v^{\prime})$. Then $ad\cdot v=v^{\prime}$, so $v$ and $v^{\prime}$ have the same $G$-orbit.
Case IV: $x_{1}^{2}(v)=x_{2}^{2}(v)=x_{3}^{2}(v)$. Applying an element of $G$ and a rescaling on $W_{1}$ (see Section 3.3) we may assume that $w_{1}=[1,1,1]^{T}$. We have the $G$-invariants
$f_{9}:=x_{1}x_{2}y_{1}y_{2}+x_{1}x_{3}y_{1}y_{3}+x_{2}x_{3}y_{2}y_{3},\qquad f_{10}:=y_{1}y_{2}y_{3}.$
The multiset $\{f_{4}(v)=f_{4}(v^{\prime}),f_{9}(v)=f_{9}(v^{\prime}),f_{10}(v)=f_{10}(v^{\prime})\}$ consists of the elementary symmetric polynomials of $y_{1}(v),y_{2}(v),y_{3}(v)$ (respectively of $y_{1}(v^{\prime}),y_{2}(v^{\prime}),y_{3}(v^{\prime})$). Thus $w^{\prime}_{2}$ is obtained from $w_{2}$ by permuting its coordinates. In fact $w_{2}$ can be taken to $w^{\prime}_{2}$ by an even permutation of the coordinates, because $f_{11}(v)=f_{11}(v^{\prime})$, where $f_{11}$ is the $G$-invariant
$f_{11}:=x_{2}x_{3}y_{1}y_{2}^{2}+x_{1}x_{2}y_{1}^{2}y_{3}+x_{1}x_{3}y_{2}y_{3}^{2}.$
It means that $w^{\prime}_{2}$ belongs to the $\langle c\rangle$-orbit of $w_{2}$. Since $\mathrm{Stab}_{G}(w_{1})=\langle c\rangle$, we conclude that $G\cdot v=G\cdot v^{\prime}$. ∎
Lemma 6.13.
Let $w_{1}$ be a non-zero element in $W_{1}$, and $w_{2}\in W_{2}$.
  • (i)
    If $|\mathrm{Stab}_{G}(w_{1})|\neq 3$, then for any $\chi\in\{(\omega,1),(\omega^{2},1),(1,1)\}$ there exists a homogeneous $f\in K[W_{1}]^{G,\chi}$ with $\deg(f)\leq 4$ such that $f(w_{1})\neq 0$.
  • (ii)
    If $|\mathrm{Stab}_{G}(w_{1})|\notin\{2,4\}$, then for any $\chi\in\{(1,-1),(1,1)\}$ there exists a homogeneous $f\in K[W_{1}]^{G,\chi}$ with $\deg(f)\leq 3$ and $f(w_{1})\neq 0$.
  • (iii)
    If there exist $i,j\in\{1,2,3\}$ with $y_{i}^{2}(w_{2})\neq y_{j}^{2}(w_{2})$, then for any $\chi\in\{(\omega,1),(\omega^{2},1),(1,1)\}$ there exists a homogeneous $f\in K[W_{2}]^{G,\chi}$ with $\deg(f)\leq 4$ such that $f(w_{2})\neq 0$.
Proof.
Set
$r_{(\omega,1)}:=x_{1}^{2}+\omega x_{2}^{2}+\omega^{2}x_{3}^{2}$ $\qquad\qquad s_{(\omega,1)}:=y_{1}^{2}+\omega y_{2}^{2}+\omega^{2}y_{3}^{2}$
$r_{(\omega^{2},1)}:=x_{1}^{2}+\omega^{2}x_{2}^{2}+\omega x_{3}^{2}$ $\qquad\qquad s_{(\omega^{2},1)}:=y_{1}^{2}+\omega^{2}y_{2}^{2}+\omega y_{3}^{2}$
$r_{(1,-1)}$ $:=x_{1}x_{2}x_{3}.$
Then $r_{\chi}$ and $s_{\chi}$ are relative $G$-invariants of weight $\chi$.
(i) Take first $\chi=(\omega,1)$. Then $r_{(\omega,1)}$, $(r_{(\omega^{2},1)})^{2}$ are both homogeneous relative invariants of degree at most $4$ and weight $\chi$, and their common zero locus in $W_{1}$ is $\{w\in W_{1}\mid x_{1}^{2}(w)=x_{2}^{2}(w)=x_{3}^{2}(w)\}$. Lemma 6.11 implies that $w_{1}$ does not belong to this common zero locus by the assumption on its stabilizer. The proof for $\chi=(\omega^{2},1)$ is similar, one uses the relative invariants $r_{(\omega^{2},1)}$, $(r_{(\omega,1)})^{2}$. For $\chi=(1,1)$ we can take $f=1$.
(ii) For $\chi=(1,-1)$ we can take $f=r_{(1,-1)}$, because Lemma 6.11 implies $(x_{1}x_{2}x_{3})(v_{1})\neq 0$ by the assumption on the stabilizer of $w_{1}$. For $\chi=(1,1)$ we can take $f=1$.
(iii) Similarly to the proof of (i), the common zero locus of the weight $(\omega,1)$ relative invariants $s_{(\omega,1)}$, $(s_{(\omega^{2},1)})^{2}$ is $\{w\in W_{2}\mid y_{1}^{2}(w)=y_{2}^{2}(w)=y_{3}^{2}(w)\}$. Now $w_{2}$ does not belong to this common zero locus by assumption. The case of $\chi=(\omega^{2},1)$ is settled using the relative invariants $s_{(\omega^{2},1)}$, $(s_{(\omega,1)})^{2}$. ∎
Theorem 6.14.
Assume that $K$ contains an element of multiplicative order $6$. Then we have the equality $\beta_{\mathrm{sep}}^{K}(\mathrm{A}_{4}\times\mathrm{C}_{2})=6$.
Proof.
By (3) we have $\beta_{\mathrm{sep}}^{K}(G)\geq\beta_{\mathrm{sep}}^{K}(\mathrm{A}_{4})$, and by Proposition 5.1 we have $\beta_{\mathrm{sep}}^{K}(\mathrm{A}_{4})\geq 6$. By Lemma 3.1 (iii), to prove the reverse inequality $\beta_{\mathrm{sep}}^{K}(G)\leq 6$ it is sufficient to show that $\beta_{\mathrm{sep}}^{L}(G)\leq 6$ for some extension $L$ of $K$. Therefore we may assume that $K$ is large enough to apply Lemma 3.6, and so it is sufficient to prove $\beta_{\mathrm{sep}}(G,V)\leq 6$ for $V:=W_{1}\oplus W_{2}\oplus U$. Take $v=(w_{1},w_{2},u)\in V$, $v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2},u^{\prime})\in V$, and assume that $f(v)=f(v^{\prime})$ for all $f\in K[V]^{G}_{\leq 6}$. We need to show that then $G\cdot v=G\cdot v^{\prime}$. By Proposition 6.12 we know that $G\cdot(w_{1},w_{2})=G\cdot(w^{\prime}_{1},w^{\prime}_{2})$, so replacing $v^{\prime}$ by an appropriate element in its $G$-orbit we may assume that $w^{\prime}_{1}=w_{1}$, $w^{\prime}_{2}=w_{2}$. By Proposition 6.10 (ii) it is sufficient to deal with the case $w_{1}\neq 0$. The Davenport constant of $G/G^{\prime}$ is $6$, therefore $G\cdot u=G\cdot u^{\prime}$. By Proposition 6.12 it is sufficient to deal with the case when $u\neq 0$, moreover, when $u$ has a non-zero component $u_{\chi}$ for some non-identity character $\chi\in\widehat{G}$.
Case I: $\mathrm{Stab}_{G}(w_{1})$ is trivial. We claim that $u_{\chi}=u^{\prime}_{\chi}$ for all $\chi\in\widehat{G}$, so $v=v^{\prime}$. Since $\chi^{2}\in\langle(\omega,1)\rangle$ (respectively, $\chi^{3}\in\langle(1,-1)\rangle$), by Lemma 6.13 there exists a relative invariant $f\in K[W_{1}]^{G,\chi^{-2}}$ (respectively, $f\in K[W_{1}]^{G,\chi^{-3}}$) such that $ft_{\chi}^{2}$ (respectively, $ft_{\chi}^{3}$) is a homogeneous $G$-invariant of degree at most $6$, and $f(v)=f(w_{1})\neq 0$. From $(ft_{\chi}^{2})(v)=(ft_{\chi}^{2})(v^{\prime})$ (respectively, $(ft_{\chi}^{3})(v)=(ft_{\chi}^{3})(v^{\prime})$) we conclude that both $t_{\chi}^{2}(u)=t_{\chi}^{2}(u^{\prime})$ and $t_{\chi}^{3}(u)=t_{\chi}^{3}(u^{\prime})$. It follows that $t_{\chi}(u)=t_{\chi}(u^{\prime})$, i.e. $u_{\chi}=u^{\prime}_{\chi}$.
Case II.a: $|\mathrm{Stab}_{G}(w_{1})|=3$ (consequently, by Lemma 6.11 we have $x_{1}^{2}(v)=x_{2}^{2}(v)=x_{3}^{2}(v)\neq 0$) and there exist $i,j\in\{1,2,3\}$ with $y_{i}^{2}(v)\neq y_{j}^{2}(v)$. By Lemma 6.13 (iii) for $\chi\in\langle(\omega,1)\rangle$ we conclude the existence of a homogeneous $f\in K[W_{2}]^{G,\chi}$ of degree at most $4$ with $f(w_{2})\neq 0$. Moreover, $x_{1}x_{2}x_{3}\in K[W_{1}]^{G,(1,-1)}$ does not vanish at $v$. In the same way as in Case I we conclude that for all $\chi\in\widehat{G}$ we have $t_{\chi}^{2}(v)=t_{\chi}^{2}(v^{\prime})$ and $t_{\chi}^{3}(v)=t_{\chi}^{3}(v^{\prime})$, implying in turn $t_{\chi}(v)=t_{\chi}(v^{\prime})$, i.e. $u_{\chi}=u^{\prime}_{\chi}$. This holds for all $\chi$, thus $v=v^{\prime}$ in this case.
Case II.b: $|\mathrm{Stab}_{G}(w_{1})|=3$ (consequently, by Lemma 6.11 we have $x_{1}^{2}(v)=x_{2}^{2}(v)=x_{3}^{2}(v)\neq 0$, hence in particular, $(x_{1}x_{2}x_{3})(v)\neq 0$) and $\mathrm{Stab}_{G}(w_{1})\subseteq\mathrm{Stab}_{G}(w_{2})$. Write $H:=\mathrm{Stab}_{G}(w_{1})$. Then $H\cong HG^{\prime}/G^{\prime}$ (since $|H|=3$ is coprime to $|G^{\prime}|=4$). It follows that an $H$-invariant monomial $h$ in $K[U]$ is either $G$-invariant or is a relative $G$-invariant with weight $(1,-1)$. Hence $h$ or $x_{1}x_{2}x_{3}h$ is a $G$-invariant, implying by $(x_{2}x_{2}x_{3})(v)\neq 0$ that $h(u)=h(u^{\prime})$ holds for all $H$-invariant monomials in $K[U]$ of degree at most $3=\mathsf{D}(\mathrm{C}_{3})=\mathsf{D}(H)=\beta_{\mathrm{sep}}^{K}(H)$. Consequently, there exists an element $g\in H$ with $g\cdot u=u^{\prime}$. Now $g\cdot(w_{1},w_{2},u)=(w_{1},w_{2},u^{\prime})$, so $G\cdot v=G\cdot v^{\prime}$.
Case II.c: $|\mathrm{Stab}_{G}(w_{1})|=3$ (consequently, by Lemma 6.11 we have $x_{1}^{2}(v)=x_{2}^{2}(v)=x_{3}^{2}(v)\neq 0$) and $y_{1}^{2}(v)=y_{2}^{2}(v)=y_{2}^{2}(v)\neq 0$. Replacing the pair $(v,v^{\prime})$ by an appropriate element in its $G$-orbit we may assume that $w_{1}$ is a non-zero scalar multiple of $[1,1,1]^{T}$. If $w_{2}$ is also a scalar multiple of $[1,1,1]^{T}$, then we are in Case II.b. So we may assume that $w_{2}$ is not a scalar multiple of $[1,1,1]^{T}$. Then necessarily $w_{2}$ is a non-zero scalar multiple of one of $[-1,1,1]^{T}$, $[1,-1,1]^{T}$, $[1,1,-1]^{T}$. Then the relative invariants
$x_{1}y_{1}+x_{2}y_{2}+x_{3}y_{3}$ $\in K[W_{1}\oplus W_{2}]^{G,(1,-1)},$
$x_{1}y_{1}+\omega x_{2}y_{2}+\omega^{2}x_{3}y_{3}$ $\in K[W_{1}\oplus W_{2}]^{G,(\omega,-1)},$
$x_{1}y_{1}+\omega^{2}x_{2}y_{2}+\omega x_{3}y_{3}$ $\in K[W_{1}\oplus W_{2}]^{G,(\omega^{2},-1)}$
do not vanish at $(w_{1},w_{2})$. Multiplying them by $x_{1}x_{2}x_{3}$ we get degree $5$ relative invariants of weight $(1,1)$, $(\omega,1)$, $(\omega^{2},1)$. It follows that for any $\chi\in\widehat{G}$ there exists a homogeneous $f\in K[W_{1}\oplus W_{2}]^{G,\chi^{-1}}$ of degree at most $5$ with $f(w_{1},w_{2})\neq 0$; then $(ft_{\chi})(v)=(ft_{\chi})(v^{\prime})$, implying $u_{\chi}=u^{\prime}_{\chi}$. This holds for all $\chi$, so $u=u^{\prime}$ and thus $v=v^{\prime}$ in this case.
Case III: $|\mathrm{Stab}_{G}(w_{1})|=4$. By Lemma 6.11 exactly one of $x_{1}(v)$, $x_{2}(v)$, $x_{3}(v)$ is non-zero; by symmetry we may assume that $x_{1}(v)\neq 0$ and $x_{2}(v)=x_{3}(v)=0$. By Lemma 6.13 (i) for any $\chi\in\langle(\omega,1)\rangle$ there exists a homogeneous $f\in K[W_{1}\oplus W_{2}]^{G,\chi}$ with $\deg(f)\leq 4$ and $f(w_{1},w_{2})\neq 0$. If we can find a homogeneous $h\in K[W_{1}\oplus W_{2}]^{G,(1,-1)}$ with $\deg(h)\leq 3$ and $h(w_{1},w_{2})\neq 0$ then we are done as in Case I. This happens in the following case:
Case III.a: $y_{1}(v)\neq 0$ or $(y_{2}y_{3})(v)\neq 0$.
Then we can take $h:=x_{1}y_{1}+x_{2}y_{2}+x_{3}y_{3}$ or $h:=x_{1}y_{2}y_{3}+x_{2}y_{3}y_{1}+x_{3}y_{1}y_{2}$.
Otherwise we are in the following case:
Case III.b: $y_{1}(v)=0$ and $(y_{2}y_{3})(v)=0$. If $y_{3}(v)=0$, then $H:=\langle ad\rangle\subseteq\mathrm{Stab}_{G}(w_{1},w_{2})$. Note that $H\cong HG^{\prime}/G^{\prime}$. A monomial $m\in K[U]^{H}$ is a relative $G$-invariant with weight in $\langle(\omega,1)\rangle$. If $\deg(m)\leq 2$, then by Lemma 6.13 (iii) we have a $G$-invariant $mf$ with $\deg(mf)\leq 6$ and $f(w_{1},w_{2})\neq 0$. It follows that $m(u)=m(u^{\prime})$ for all $H$-invariant monomials $m\in K[U]$ with $\deg(m)\leq 2=\mathsf{D}(H)=\beta_{\mathrm{sep}}(H)$. Consequently, $H\cdot u=H\cdot u^{\prime}$, and as $H$ stabilizes $(w_{1},w_{2})$, we conclude $G\cdot v=G\cdot v^{\prime}$. The case $y_{2}(v)=0$ is similar, we just need to take $H:=\langle abd\rangle\subseteq\mathrm{Stab}_{G}(w_{1},w_{2})$ in the above argument.
Case IV: $|\mathrm{Stab}_{G}(w_{1})|=2$. By Lemma 6.11 exactly one of $x_{1}(v)$, $x_{2}(v)$, $x_{3}(v)$ is zero; by symmetry we may assume that $x_{1}(v)\neq 0$, $x_{2}(v)\neq 0$, and $x_{3}(v)=0$. Note that this implies that $H:=\mathrm{Stab}_{G}(w_{1})=\langle ad\rangle$ is not contained in $G^{\prime}$, hence $H\cong HG^{\prime}/G^{\prime}$. By Lemma 6.13 (i) for any $\chi\in\langle(\omega,1)\rangle$ there exists a homogeneous $f\in K[W_{1}\oplus W_{2}]^{G,\chi}$ with $\deg(f)\leq 4$ and $f(w_{1},w_{2})\neq 0$. If we can find a homogeneous $h\in K[W_{1}\oplus W_{2}]^{G,(1,-1)}$ with $\deg(h)\leq 3$ and $h(w_{1},w_{2})\neq 0$ then we are done as in Case I, whereas if $H\subseteq\mathrm{Stab}_{G}(w_{2})$ then we can finish as in Case III.b.
Case IV.a: $y_{1}(v)=y_{2}(v)=0$. Then $ad\cdot w_{2}=w_{2}$, hence $H\subseteq\mathrm{Stab}_{G}(w_{2})$, so we are done, as we pointed out above.
Case IV.b: exactly one of $y_{1}(v)$, $y_{2}(v)$ is zero. Then $h:=x_{1}y_{1}+x_{2}y_{2}+x_{3}y_{3}$ is a degree $\leq 3$ relative $G$-invariant with weight $(1,-1)$ and not vanishing at $(w_{1},w_{2})$, so we are done, as we explained in the beginning of Case IV.
Case IV.c: Both of $y_{1}(v)$, $y_{2}(v)$ are non-zero. We claim that for any non-zero weight $\chi\in\widehat{G}$ there exists a a relative invariant $f\in K[W_{1}\oplus W_{2}]^{G,\chi^{-1}}$ with $\deg(f)\leq 4$ such that $f(w_{1},w_{2})\neq 0$. Then $ft_{\chi}$ is a $G$-invariant of degree $\leq 5$, so $(ft_{\chi})(v)=(ft_{\chi})(v^{\prime})$ implies $u_{\chi}=u^{\prime}_{\chi}$ for all non-trivial $\chi\in\widehat{G}$, thus $v=v^{\prime}$. For the weights $\chi\in\langle(\omega,1)\rangle$ the claim follows from Lemma 6.13 (i). For $\chi=(1,-1)$ we may take $f:=x_{1}^{2}x_{2}y_{2}+x_{2}^{2}x_{3}y_{3}+x_{3}^{2}x_{1}y_{1}$. For $\chi=(\omega,-1)$ consider the relative invariants
$r_{(\omega,-1)}^{(1)}:=x_{1}y_{1}+\omega x_{2}y_{2}+\omega^{2}x_{3}y_{3}$
$r_{(\omega,-1)}^{(2)}:=x_{1}y_{1}^{3}+\omega x_{2}y_{2}^{3}+\omega^{2}x_{3}y_{3}^{3}$
$r_{(\omega,-1)}^{(3)}:=x_{1}^{3}y_{1}+\omega x_{2}^{3}y_{2}+\omega^{2}x_{3}^{3}y_{3}$
in $K[W_{1}\oplus W_{2}]^{G,\chi}$. One of them does not vanish at $(w_{1},w_{2})$. Indeed, suppose for contradiction that all of them vanish at $(w_{1},w_{2})$. From $r_{(\omega,-1)}^{(1)}(w_{1},w_{2})=0=r_{(\omega,-1)}^{(2)}(w_{1},w_{2})$ we deduce $y_{1}^{2}(w_{2})=y_{2}^{2}(w_{2})$. Thus $[y_{1}(v),y_{2}(v)]$ is a non-zero scalar multiple of $[1,1]$ or $[1,-1]$. After that, from $r_{(\omega,-1)}^{(1)}(w_{1},w_{2})=0=r_{(\omega,-1)}^{(3)}(w_{1},w_{2})$ we conclude $x_{1}^{2}(v)=x_{2}^{2}(v)$ and hence $x_{1}(v)=\pm x_{2}(v)$, which clearly contradicts to the assumption that $r_{(\omega,-1)}^{(1)}(w_{1},w_{2})=0$. The case of the weight $\chi=(\omega^{2},-1)$ can be settled similarly, using the relative invariants $x_{1}y_{1}+\omega^{2}x_{2}y_{2}+\omega x_{3}y_{3}$, $x_{1}y_{1}^{3}+\omega^{2}x_{2}y_{2}^{3}+\omega x_{3}y_{3}^{3}$, $x_{1}^{3}y_{1}+\omega^{2}x_{2}^{3}y_{2}+\omega x_{3}^{3}y_{3}$. ∎
7. The group $\mathrm{S}_{3}\times\mathrm{C}_{3}$
In this section
$G=\mathrm{S}_{3}\times\mathrm{C}_{3}=\langle a,b\mid a^{3}=b^{2}=1,\ ba=a^{2}b\rangle\times\langle c\mid c^{3}=1\rangle.$
Assume in addition that $K$ contains an element $\omega$ of multiplicative order $3$, and consider the following irreducible $2$-dimensional representations of $G$:
$\psi_{1}:a\mapsto\begin{bmatrix}\omega&0\\ 0&\omega^{2}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix},\quad c\mapsto\begin{bmatrix}1&0\\ 0&1\\ \end{bmatrix},$
$\psi_{2}:a\mapsto\begin{bmatrix}\omega&0\\ 0&\omega^{2}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix},\quad c\mapsto\begin{bmatrix}\omega&0\\ 0&\omega\\ \end{bmatrix},$
$\psi_{3}:a\mapsto\begin{bmatrix}\omega&0\\ 0&\omega^{2}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix},\quad c\mapsto\begin{bmatrix}\omega^{2}&0\\ 0&\omega^{2}\\ \end{bmatrix}.$
The other irreducible representations of $G$ are $1$-dimensional, and can be labelled by the group $\widehat{G}=\{\pm 1\}\times\{1,\omega,\omega^{2}\}$, where $\chi=(\chi_{1},\chi_{2})\in\widehat{G}$ is identified with the representation
$\chi:a\mapsto 1,\ b\mapsto\chi_{1},\ c\mapsto\chi_{2}$
(note that $\langle a\rangle$ is the commutator subgroup of $G$, so $a$ is in the kernel of any $1$-dimensional representation of $G$). For $j=1,2,3$ denote by $W_{j}$ the vector space $K^{2}$ endowed with the representation $\psi_{j}$, and for $\chi\in\widehat{G}$ denote by $U_{\chi}$ the vector space $K$ endowed with the representation $\chi$, and set $U:=\bigoplus_{\chi\in\widehat{G}}U_{\chi}$.
For $\xi,\eta,\zeta\in\{x,y,z\}$ set
$q_{\xi\eta}:=\frac{1}{2}(\xi_{1}\eta_{2}+\xi_{2}\eta_{1})\quad\text{ and }\quad p_{\xi\eta\nu}:=\xi_{1}\eta_{1}\zeta_{1}+\xi_{2}\eta_{2}\zeta_{2}.$
For example, $q_{xx}=x_{1}x_{2}$, $q_{yz}=\frac{1}{2}(y_{1}z_{2}+y_{2}z_{1})$, $p_{xxx}=x_{1}^{3}+x_{2}^{3}$, $p_{xxy}=x_{1}^{2}y_{1}+x_{2}^{2}y_{2}$. It is well known (see for example [32, Theorem 4.1]) that the elements $q_{\xi\eta}$, $p_{\xi\eta\zeta}$ minimally generate $K[W_{1}\oplus W_{2}\oplus W_{3}]^{\mathrm{S}_{3}}$, where $\mathrm{S}_{3}$ is identified with the subgroup $\langle a,b\rangle$ of $G$. Set
$A_{(1,1)}$ :=$:=\{q_{xx},q_{yz},p_{xxx},p_{yyy},p_{zzz},p_{xyz}\}$
$A_{(1,\omega)}$ :=$:=\{q_{xy},q_{zz},p_{xxy},p_{xzz},p_{yyz}\}$
$A_{(1,\omega^{2})}$ :=$:=\{q_{xz},q_{yy},p_{xxz},p_{xyy},p_{yzz}\}.$
For $\chi\in\{(1,1),(1,\omega),(1,\omega^{2})\}\subset\widehat{G}$ the elements of $A_{\chi}$ are relative $G$-invariants of weight $\chi$.
Proposition 7.1.
A homogeneous generating system of the algebra $K[W_{1}\oplus W_{2}\oplus W_{3}]^{G}$ is
(18) $B:=A_{(1,1)}\cup\{f_{1}f_{2}\mid f_{1}\in A_{(1,\omega)},f_{2}\in A_{(1,\omega^{2})}\}\cup\{q_{xy}^{3},q_{xy}^{2}q_{zz},q_{xy}q_{zz}^{2},q_{zz}^{3}\}$
$\cup\{q_{xz}^{3},q_{xz}^{2}q_{yy},q_{xz}q_{yy}^{2},q_{yy}^{3}\}\cup\{q_{zz}^{2}p_{xzz},\ q_{yy}^{2}p_{xyy}\}.$
Proof.
If $\chi^{(1)},\dots,\chi^{(k)}$ is an irreducible product-one sequence over the subgroup $\langle(1,\omega)\rangle$ of $\widehat{G}$ and $f_{j}\in A_{\chi^{(j)}}$ for $j=1,\dots k$, then $f_{1}\dots f_{k}$ is a $G$-invariant, and the invariants of this form constitute a homogeneous generating system of $K[W_{1}\oplus W_{2}\oplus W_{3}]^{G}$ (see the discussion in Section 3.7 about the relevance of the Davenport constant for our study). It follows that $K[W_{1}\oplus W_{2}\oplus W_{3}]^{G}$ is generated by
(19) $A:=A_{(1,1)}\cup\{f_{1}f_{2}\mid f_{1}\in A_{(1,\omega)},f_{2}\in A_{(1,\omega^{2})}\}$
$\cup\{f_{1}f_{2}f_{3}\mid f_{1},f_{2},f_{3}\in A_{(1,\omega)}\text{ or }f_{1},f_{2},f_{3}\in A_{(1,\omega^{2})}\}.$
By [15, Proposition 3.3] we have the inequality $\beta^{K}(G)=8$, so we can omit the degree $9$ elements from the generating system $A$. Therefore in order to prove that $B$ generates $K[W_{1}\oplus W_{2}\oplus W_{3}]^{G}$, it is sufficient to show that all the elements of $A\setminus B$ of degree $7$ or $8$ are contained in the ideal of $K[W_{1}\oplus W_{2}\oplus W_{3}]$ generated by the invariants of degree at most $6$. All elements of degree $7$ or $8$ in $A\setminus B$ are products of the form $f_{1}f_{2}f_{3}$ where $\deg(f_{1})=2$, $\deg(f_{2})=3$, and both of $f_{1}$ and $f_{2}$ belong either to $A_{(1,\omega)}$ or to $A_{(1,\omega^{2})}$. The equalities
$2q_{xy}p_{xxy}$ $=p_{xxx}q_{yy}+q_{xx}p_{xyy}$
$q_{xy}p_{xzz}$ $=q_{xx}p_{yzz}+q_{yz}p_{xxz}-p_{xyz}q_{xz}$
$q_{xy}p_{yyz}$ $=p_{yyy}q_{xz}-q_{yz}p_{xyy}+p_{xyz}q_{yy}$
$q_{zz}p_{xxy}$ $=2p_{xyz}q_{xz}-q_{xx}p_{yzz}$
$q_{zz}p_{yyz}$ $=2q_{yz}p_{yzz}-p_{zzz}q_{yy}$
show that with the only exception of $q_{zz}p_{xzz}$, all products $f_{1}f_{2}$ with $f_{1},f_{2}\in A_{(1,\omega)}$, $\deg(f_{1})=2$, $\deg(f_{2})=3$ are contained in the ideal generated by the $G$-invariants of degree at most $3$. Furthermore, we have
$p_{xzz}^{2}=4q_{xx}q_{zz}^{2}-4q_{xz}^{2}q_{zz}+p_{zzz}p_{xxz},$
showing that $q_{zz}p_{xzz}p_{xzz}$ is contained in the ideal generated by the $G$-invariants of degree at most $4$. Thus the elements of $A\setminus B$ of degree $7$ or $8$ that are the product of three factors from $A_{(1,\omega)}$ are contained in the subalgebra of $G$-invariants of degree at most $6$. Similar considerations work for products of the form $f_{1}f_{2}f_{3}$ with $f_{i}\in A_{(1,\omega^{2})}$, one just needs to interchange the roles of $y$ and $z$. This shows that $B$ is indeed a homogeneous generating system of $K[W_{1}\oplus W_{2}\oplus W_{3}]^{G}$. ∎
Proposition 7.2.
We have the inequality $\beta_{\mathrm{sep}}(G,W_{1}\oplus W_{2}\oplus W_{3})\leq 6$.
Proof.
Let $v=(w_{1},w_{2},w_{3})$, $v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2},w^{\prime}_{3})\in W_{1}\oplus W_{2}\oplus W_{3}$, and suppose that $f(v)=f(v^{\prime})$ for all $f\in K[W_{1}\oplus W_{2}\oplus W_{3}]^{G}$ with $\deg(f)\leq 6$. By Proposition 7.1 it is sufficient to show that
(20) $(q_{zz}^{2}p_{xzz})(v)$ $=(q_{zz}^{2}p_{xzz})(v^{\prime})$
(21) $(q_{yy}^{2}p_{xyy})(v)$ $=(q_{yy}^{2}p_{xyy})(v^{\prime}).$
We show (20), the argument for (21) is obtained by interchanging the roles of $y$ and $z$. By Proposition 7.1 we know that $K[W_{3}]^{G}$ is generated in degree $\leq 6$. So by assumption $w_{3}$ and $w^{\prime}_{3}$ belong to the same $G$-orbit. Replacing $w$ by an appropriate element in its orbit we may assume that $w_{3}=w^{\prime}_{3}$. Proposition 7.1 also shows that $K[W_{1}]^{G}$ is generated in degree at most $3$, so $G\cdot w_{1}=G\cdot w^{\prime}_{1}$. If $w_{1}=0$, then $w^{\prime}_{1}=0$, and both sides of (20) are zero. From now on we assume that $w_{1}\neq 0$, and so equivalently, $w^{\prime}_{1}\neq 0$. If $z_{1}(w_{3})=0$ or $z_{2}(w_{3})=0$, then $q_{zz}(v)=0=q_{zz}(v^{\prime})$, hence $(q_{zz}^{2}p_{xzz})(v)=0=(q_{zz}^{2}p_{xzz})(v^{\prime})$. In particular, (20) holds. From now on we assume in addition that
(22) $z_{1}(w_{3})\neq 0\text{ and }z_{2}(w_{3})\neq 0.$
By assumption we have
$(q_{zz}q_{xz})(v)=(q_{zz}q_{xz})(v^{\prime}).$
hence
$q_{xz}(v)=q_{xz}(v^{\prime}).$
Again by assumption we have
$(p_{xzz}q_{xz})(v)=(p_{xzz}q_{xz})(v^{\prime}).$
Therefore if $q_{xz}(v)\neq 0$, then we have
$p_{xzz}(v)=p_{xzz}(v^{\prime}),$
implying in turn the desired equality (20). It remains to deal with the case
$q_{xz}(v)=0=q_{xz}(v^{\prime}).$
Now $q_{xz}(v)=0$ implies that
(23) $w_{1},w^{\prime}_{1}\in K^{\times}[-z_{1}(w_{3}),z_{2}(w_{3})]^{T}.$
Consequently, $w^{\prime}_{1}$ is a non-zero scalar multiple of $w_{1}$. On the other hand, $w^{\prime}_{1}\in G\cdot w_{1}$, thus $w_{1}$ is an eigenvector of some matrix in $\{\psi_{1}(g)\mid g\in G\}=\psi_{1}(G)$. If the corresponding eigenvalue is $1$, then $(w_{1},w_{3})=(w^{\prime}_{1},w^{\prime}_{3})$ and (20) holds. If the eigenvalue is not $1$, then $w_{1}$ is the $-1$-eigenvector of some non-diagonal element of $\psi_{1}(G)$ (since the eigenvectors with eigenvalue different from $1$ of the diagonal elements in $\psi_{1}(G)$ are scalar multiples of $[1,0]^{T}$ or $[0,1]^{T}$, and $w_{1}$ is not of this form by (23) and (22)). After rescaling $W_{1}$ and $W_{3}$ (see Section 3.3), we have that
$(w_{1},w_{3})=([1,-1]^{T},[1,1]^{T})\text{ and }(w^{\prime}_{1},w^{\prime}_{3})=([-1,1]^{T},[1,1]^{T}),\text{ or }$
$(w_{1},w_{3})=([\omega,-1]^{T},[\omega,1]^{T})\text{ and }(w^{\prime}_{1},w^{\prime}_{3})=([-\omega,1]^{T},[\omega,1]^{T}),\text{ or }$
$(w_{1},w_{3})=([\omega^{2},-1]^{T},[\omega^{2},1]^{T})\text{ and }(w^{\prime}_{1},w^{\prime}_{3})=([-\omega^{2},1]^{T},[\omega^{2},1]^{T})$
Now one can directly check that $p_{xzz}(v)=0=p_{xzz}(v^{\prime})$ in each of these cases (in fact, $(w^{\prime}_{1},w^{\prime}_{3})=g\cdot(w_{1},w_{3})$ for some $g\in\{b,ab,a^{2}b\}$ in these cases), and the desired (20) holds. ∎
For $\xi,\eta,\zeta\in\{x,y,z\}$ set
$q^{-}_{\xi\eta}:=\frac{1}{2}(\xi_{1}\eta_{2}-\xi_{2}\eta_{1})\quad\text{ and }\quad p^{-}_{\xi\eta\nu}:=\xi_{1}\eta_{1}\zeta_{1}-\xi_{2}\eta_{2}\zeta_{2},$
and define
$A_{(-1,1)}$ :=$:=\{q^{-}_{yz},p^{-}_{xxx},p^{-}_{yyy},p^{-}_{zzz},p^{-}_{xyz}\}$
$A_{(-1,\omega)}$ :=$:=\{q^{-}_{xy},p^{-}_{xxy},p^{-}_{xzz},p^{-}_{yyz}\}$
$A_{(-1,\omega^{2})}$ :=$:=\{q^{-}_{xz},p^{-}_{xxz},p^{-}_{xyy},p^{-}_{yzz}\}.$
The elements of $A_{\chi}$ are relative invariants of weight $\chi$. (In fact every relative $G$-invariant in $K[W_{1}\oplus W_{2}\oplus W_{3}]$ is a linear combination of products of elements from $\bigcup_{\chi\in\widehat{G}}A_{\chi}$, but we do not need this.)
Proposition 7.3.
We have the equality $\beta(G,U)=5$.
Proof.
The action of $G$ on $U$ factors through the regular representation of $\mathrm{C}_{3}\times\mathrm{C}_{3}\cong G/G^{\prime}$, the factor of $G$ modulo its commutator subgroup. Therefore $\beta(G,U)=\mathsf{D}(\mathrm{C}_{3}\times\mathrm{C}_{3})=5$. ∎
Proposition 7.4.
We have the inequality $\beta(W_{1}\oplus U)\leq 6$.
Proof.
We have $K[W_{1}]^{G,(-1,1)}=Kp^{-}_{xxx}\oplus(\mathcal{H}(G,W_{1})\cap K[W_{1}]^{G,(-1,1)})$, and $K[W_{1}]^{G,\chi}=\{0\}$ for all $\chi\in\widehat{G}\setminus\{(\pm 1,1)\}$. Therefore by Lemma 3.9, $K[W_{1}\oplus U]^{G}$ is generated by $K[W_{1}]^{G}$, $K[U]^{G}$, and
$\{p^{-}_{xxx}t_{\chi^{(1)}}\cdots t_{\chi^{(k)}}\mid\chi^{(1)},\dots,\chi^{(k)}\text{ is a product-one free sequence over }\widehat{G},$
$\chi^{(1)}\cdots\chi^{(k)}=(-1,1)\}.$
Since $\mathsf{D}(\mathrm{C}_{6}/\mathrm{C}_{2})=\mathsf{D}(\mathrm{C}_{3})=3$, for $k\geq 3$ the sequence $\chi^{(1)},\dots\chi^{(k)}$ necessarily has a subsequence with product (±$(\pm 1,1)$, so the above generators of the form $p^{-}_{xxx}t_{\chi^{(1)}}\cdots t_{\chi^{(k)}}$ have degree at most $3+3$. We have $K[W_{1}]^{G}=K[q_{xx},p_{xxx}]$, and $\beta(G,U)=\mathsf{D}(\mathrm{C}_{6})=6$. Thus each generator of $K[W_{1}\oplus U]^{G}$ has degree at most $6$. ∎
Lemma 7.5.
For a non-zero $w_{1}\in W_{1}$ we have $|\mathrm{Stab}_{G}(w_{1})/\langle c\rangle|\in\{1,2\}$, and
$|\mathrm{Stab}_{G}(w_{1})/\langle c\rangle|=2\iff p^{-}_{xxx}(w_{1})=0\iff\mathrm{Stab}_{G}(w_{1})\in\{\langle b,c\rangle,\langle ab,c\rangle,\langle a^{2}b,c\rangle\}.$
For a non-zero $w_{j}\in W_{j}$ ($j\in\{2,3\}$) we have $|\mathrm{Stab}_{G}(w_{j})|\in\{1,2,3\}$, and
  • (i)
    for $j=2$:
    $|\mathrm{Stab}_{G}(w_{2})|=2\iff p^{-}_{yyy}(w_{2})=0\iff\mathrm{Stab}_{G}(w_{2})\in\{\langle b\rangle,\langle ab\rangle,\langle a^{2}b\rangle\}$
    $|\mathrm{Stab}_{G}(w_{2})|=3\iff q_{yy}(w_{2})=0\iff\mathrm{Stab}_{G}(w_{2})\in\{\langle ac\rangle,\langle ac^{2}\rangle\}.$
  • (ii)
    for $j=3$:
    $|\mathrm{Stab}_{G}(w_{3})|=2\iff p^{-}_{zzz}(w_{3})=0\iff\mathrm{Stab}_{G}(w_{3})\in\{\langle b\rangle,\langle ab\rangle,\langle a^{2}b\rangle\}$
    $|\mathrm{Stab}_{G}(w_{3})|=3\iff q_{zz}(w_{3})=0\iff\mathrm{Stab}_{G}(w_{3})\in\{\langle ac\rangle,\langle ac^{2}\rangle\}.$
Proof.
The statement can be verified by straightforward direct computation. ∎
Lemma 7.6.
Let $W$ stand for $W_{2}$ or $W_{3}$. Let $w\in W$, $u,u^{\prime}\in U$ such that $f(w,u)=f(w,u^{\prime})$ holds for all $f\in K[W\oplus U]^{G}$ with $\deg(f)\leq 6$. Then $u$ and $u^{\prime}$ belong to the same $\mathrm{Stab}_{G}(w)$-orbit.
Proof.
Consider first the case $W=W_{2}$. Assume that $\mathrm{Stab}_{G}(w)=\{1_{G}\}$. Then $q_{yy}(w)\neq 0$ and $p^{-}_{yyy}(w)\neq 0$ by Lemma 7.5. If $\chi\in\{(1,\omega),(-1,1),(1,\omega^{2}),(-1,\omega)\}$, then there is an $f\in\{q_{yy},p^{-}_{yyy},q_{yy}^{2},q_{yy}p^{-}_{yyy}\}$ such that $ft_{\chi}$ is a $G$-invariant of degree at most $6$, hence $t_{\chi}(u)=t_{\chi}(u^{\prime})$. If $\chi=(-1,\omega^{2})$, then $q_{yy}t_{\chi}^{2}$ and $p^{-}_{yyy}t_{\chi}^{3}$ are $G$-invariants showing that $t_{\chi}(u)=t_{\chi}(u^{\prime})$. Finally, if $\chi=(1,1)$, then $t_{\chi}$ is a $G$-invariant of degree $1$, so $t_{\chi}(u)=t_{\chi}(u^{\prime})$. Summarizing, we have $u=u^{\prime}$.
Assume next that $|\mathrm{Stab}_{G}(w)|=2$; this means that $H:=\mathrm{Stab}_{G}(w)$ is $\langle b\rangle$, $\langle ab\rangle$, or $\langle a^{2}b\rangle$, and $q_{yy}(w)\neq 0$. For any $H$-invariant monomial $T:=t_{\chi}\cdots t_{\chi^{\prime}}$ with $\deg(T)\leq 2$, either $T$, $q_{yy}T$, or $q_{yy}^{2}T$ is a $G$-invariant of degree at most $6$. It follows that $T(u)=T(u^{\prime})$ for any $H$-invariant of degree at most $2$. Therefore there exists an $h\in H$ with $h\cdot u=u^{\prime}$ (note that $\beta_{\mathrm{sep}}^{K}(H)=\beta^{K}(H)=2$).
Assume next that $|\mathrm{Stab}_{G}(w)|=3$; this means that $H:=\mathrm{Stab}_{G}(w)$ is $\langle ac\rangle$ or $\langle ac^{2}\rangle$, and $p^{-}_{yyy}(w)\neq 0$. For any $H$-invariant monomial $T:=t_{\chi}\cdots t_{\chi^{\prime}}$ with $\deg(T)\leq 3$, either $T$ or $p^{-}_{yyy}T$ is a $G$-invariant of degree at most $6$. It follows that $T(u)=T(u^{\prime})$ for any $H$-invariant of degree at most $3$. Therefore there exists an $h\in H$ with $h\cdot u=u^{\prime}$ (note that $\beta_{\mathrm{sep}}^{K}(H)=\beta^{K}(H)=3$).
By Lemma 7.5, the only remaining case is when $w=0$, and then the result follows from Proposition 7.3.
The case $W=W_{3}$ follows by Lemma 3.11. Indeed, the group $G$ has the automorphism $\alpha:a\mapsto a$, $b\mapsto b$, $c\mapsto c^{-1}$. We have $\psi_{2}=\psi_{3}\circ\alpha$, so $\alpha$ interchanges $W_{2}$ and $W_{3}$, and permutes the $1$-dimensional representations. ∎
Lemma 7.7.
Let $v=(w_{i},w_{j})\in V=W_{i}\oplus W_{j}$, where $(i,j)\in\{(1,2),(1,3),(2,3)\}$. Suppose that $\mathrm{Stab}_{G}(w_{i})$, $\mathrm{Stab}_{G}(w_{j})$ are both nontrivial, and $\mathrm{Stab}_{G}(v)=\{1_{G}\}$. Then for any $\chi\in\widehat{G}$ there exists a relative invariant $f\in K[V]^{G,\chi}$ with $\deg(f)\leq 4$ and $f(v)\neq 0$.
Proof.
For the trivial character $(1,1)$ we may take $f=1\in K[V]^{G}$.
Case 1: $V=W_{2}\oplus W_{3}$. Then the assumptions on the stabilizers imply by Lemma 7.5 that the stabilizers of $w_{2}$ and $w_{3}$ both have order $2$ or $3$, moreover, exactly one of $q_{yy}(w_{2})$ and $p^{-}_{yyy}(w_{2})$ is zero, and exactly one of $q_{zz}(w_{3})$ and $p^{-}_{zzz}(w_{3})$ is zero.
Case 1.a: $q_{yy}(w_{2})=0=q_{zz}(w_{3})$. Then both $w_{2}$ and $w_{3}$ are non-zero scalar multiples of $[1,0]^{T}$ or $[0,1]^{T}$. There is no harm to assume that $w_{2}=[1,0]^{T}$ (the case $w_{2}\in K^{\times}[0,1]^{T}$ is similar). Then $\mathrm{Stab}_{G}(w_{2})=\langle ac^{2}\rangle$, and $w_{3}$ can not be a scalar multiple of $[0,1]^{T}$ (since the stabilizer of $[0,1]^{T}\in W_{3}$ is also $\langle ac^{2}\rangle$). This means that $w_{3}$ is a non-zero scalar multiple of $[1,0]^{T}$ as well, and so none of $\{p_{yyz},p_{yzz},p^{-}_{yyz},p^{-}_{yzz},p^{-}_{zzz}\}$ vanishes on $v=(w_{2},w_{3})$. These are relative invariants of degree $3$, and their weights represent all the five non-trivial characters.
Case 1.b: $q_{yy}(w_{2})=0=p^{-}_{zzz}(w_{3})$. Then $w_{2}$ is a non-zero scalar multiple of $[1,0]^{T}$ or $[0,1]^{T}$, and $w_{3}$ is a non-zero scalar multiple of $[1,1]^{T}$, $[1,\omega]^{T}$, $[1,\omega^{2}]^{T}$. It follows that none of $\{p_{yyz},p_{yzz},p^{-}_{yyz},p^{-}_{yzz},p^{-}_{yyy}\}$ vanishes on $v$. These are relative invariants of degree $3$, and their weights represent all the five non-trivial characters.
Case 1.c: $p^{-}_{yyy}(w_{2})=0=q_{zz}(w_{3})$: same as Case 1.b, we just need to interchange the roles of $W_{2}$ and $W_{3}$.
Case 1.d: $p^{-}_{yyy}(w_{2})=0=p^{-}_{zzz}(w_{3})$. Then both of $w_{2},w_{3}$ are non-zero scalar multiples of $[1,1]^{T}$, $[1,\omega]^{T}$, or $[1,\omega^{2}]^{T}$. Moreover, $w_{2}$ and $w_{3}$ are linearly independent (otherwise they would have the same stabilizer). Thus none of $\{q_{yy},q_{zz},q^{-}_{yz},q_{yy}q^{-}_{yz},q_{zz}q^{-}_{yz}\}$ vanishes on $v$. These are relative invariants of degree at most $4$, and their weights represent all the five non-trivial characters.
Case 2: $V=W_{1}\oplus W_{2}$. Then the assumptions on the stabilizers imply by Lemma 7.5 that exactly one of $q_{yy}(w_{2})$ and $p^{-}_{yyy}(w_{2})$ is zero.
Case 2.a: $q_{yy}(w_{2})=0$ and $p^{-}_{yyy}(w_{2})\neq 0$. Then $w_{2}$ is a non-zero scalar multiple of $[1,0]^{T}$ or $[0,1]^{T}$, say $w_{2}$ is a non-zero scalar multiple of $[1,0]^{T}$. If $w_{1}$ is a scalar multiple of $[0,1]^{T}$, then none of $q_{xy}$, $q^{-}_{xy}$, $q_{xy}q^{-}_{xy}$, $q_{xy}^{2}$, $p^{-}_{yyy}$ vanishes on $v=(w_{1},w_{2})$, and these relative invariants on $V$ of degree at most $4$ exhaust all the five non-trivial weights, so we are done. Otherwise the first coordinate of $w_{1}$ is non-zero, and thus $p_{xxy}$, $p_{xyy}$, $p^{-}_{xxy}$, $p^{-}_{xyy}$ are relative invariants on $V$ of degree $3$ not vanishing on $v$ (just like $p^{-}_{yyy}$). So we have exhausted all the five non-trivial weights, and we are done.
Case 2.b: $p^{-}_{yyy}(w_{2})=0$ and $q_{yy}(w_{2})\neq 0$. Then $w_{2}$ is a non-zero scalar multiple of $[1,1]^{T}$, $[1,\omega]^{T}$, or $[1,\omega^{2}]^{T}$. Moreover, $w_{1}$ is not a scalar multiple of $w_{2}$ (since otherwise its stabilizer would contain the stabilizer of $w_{2}$). Then $q_{yy}$, $q^{-}_{xy}$, $q_{yy}^{2}$, $q^{-}_{xy}q_{yy}$, $p^{-}_{xyy}$ are non-zero on $v=(w_{1},w_{2})$, and these relative invariants on $W_{1}\oplus W_{2}$ of degree at most $4$ exhaust all the five non-trivial weights.
Case 3: $V=W_{1}\oplus W_{3}$. The group $G$ has the automorphism $\alpha:a\mapsto a$, $b\mapsto b$, $c\mapsto c^{2}$. We have $\psi_{3}=\psi_{2}\circ\alpha$, so $\alpha$ interchanges $W_{2}$ and $W_{3}$, and permutes the $1$-dimensional representations. So Case 3 follows from Case 2 by Lemma 3.11. ∎
Theorem 7.8.
Assume that $K$ has an element of multiplicative order $6$. Then we have $\beta_{\mathrm{sep}}^{K}(\mathrm{S}_{3}\times\mathrm{C}_{3})=6$.
Proof.
The cyclic group $\mathrm{C}_{6}$ is a homomorphic image of $G$, therefore $\beta_{\mathrm{sep}}^{K}(G)\geq\beta_{\mathrm{sep}}^{K}(\mathrm{C}_{6})=6$.
Let us turn to the proof of the reverse inequality $\beta_{\mathrm{sep}}^{K}(G)\leq 6$. By Lemma 3.1 we may assume that $K$ is large enough so that Lemma 3.6 applies and we can reduce to the study of multiplicity-free representations. Set $V:=W_{1}\oplus W_{2}\oplus W_{3}\oplus U$, take $v:=(w_{1},w_{2},w_{3},u)\in V$ and $v^{\prime}:=(w_{1}^{\prime},w_{2}^{\prime},w_{3}^{\prime},u^{\prime})\in V$, and assume that $f(v)=f(v^{\prime})$ for all $f\in K[V]^{G}$ with $\deg(f)\leq 6$. We need to show that $G\cdot v=G\cdot v^{\prime}$. By Proposition 7.2, $(w_{1},w_{2},w_{3})$ and $(w_{1}^{\prime},w_{2}^{\prime},w_{3}^{\prime})$ have the same $G$-orbit in $W_{1}\oplus W_{2}\oplus W_{3}$. So replacing $v^{\prime}$ by an appropriate element in its $G$-orbit, we may assume that $w_{1}=w_{1}^{\prime}$, $w_{2}=w_{2}^{\prime}$, $w_{3}=w_{3}^{\prime}$.
If $w_{2}=w_{3}=0$, then $v$ and $v^{\prime}$ are in the same $G$-orbit by Proposition 7.4. Otherwise there exists a $k\in\{2,3\}$ such that $w_{k}\neq 0$, and thus $|\mathrm{Stab}_{G}(w_{k})|\in\{1,2,3\}$. If $\mathrm{Stab}_{G}(w_{2})=\{1_{G}\}$ or $\mathrm{Stab}_{G}(w_{3})=\{1_{G}\}$, then $u=u^{\prime}$ (and hence $v=v^{\prime}$) by Lemma 7.6. Similarly, if $\mathrm{Stab}_{G}(w_{l})\supseteq\mathrm{Stab}_{G}(w_{k})$ for both $l\in\{1,2,3\}\setminus\{k\}$, then $\mathrm{Stab}_{G}(w_{k})\subseteq\mathrm{Stab}_{G}(w_{1},w_{2},w_{3})$. By Lemma 7.6 we have a $g\in\mathrm{Stab}_{G}(w_{k})$ with $g\cdot u=u^{\prime}$. Then $g\cdot v=v^{\prime}$, and we are done.
It remains to deal with the case when there exists a pair $(i,j)\in\{(1,2),(1,3),(2,3)\}$ such that $(w_{i},w_{j})$ satisfies the assumptions of Lemma 7.7. Therefore for any $\chi\in\widehat{G}$ there exists a relative invariant $f\in K[W_{1}\oplus W_{2}\oplus W_{3}]^{G,\chi^{-1}}$ with $f(w_{1},w_{2},w_{3})\neq 0$ and $\deg(f)\leq 4$. Then $ft_{\chi}\in K[V]^{G}$ has degree at most $5$, thus $(ft_{\chi})(v)=(ft_{\chi})(v^{\prime})$, implying in turn that
$t_{\chi}(u)=\frac{(ft)(v)}{f(w_{1},w_{2},w_{3})}=\frac{(ft)(v)}{f(w_{1},w_{2},w_{3})}=t_{\chi}(u^{\prime}).$
This holds for all $\chi\in\widehat{G}$, so $u=u^{\prime}$ and hence $v=v^{\prime}$. ∎
8. Some groups $G=\mathrm{C}_{n}\rtimes\mathrm{C}_{k}$ with $\beta_{\mathrm{sep}}(G)<\beta^{K}(G)$
8.1. The group $\mathrm{C}_{4}\rtimes\mathrm{C}_{4}$
In this section
$G=\mathrm{C}_{4}\rtimes\mathrm{C}_{4}=\langle a,b\mid a^{4}=1=b^{4},\quad bab^{-1}=a^{3}\rangle.$
Assume that $K$ has an element $\mathrm{i}$ of multiplicative order $4$. Then $G$ has two irreducible $2$-dimensional representations (up to isomorphism), namely
$\psi_{1}:a\mapsto\begin{bmatrix}\mathrm{i}&0\\ 0&-\mathrm{i}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix},\qquad\psi_{2}:a\mapsto\begin{bmatrix}\mathrm{i}&0\\ 0&-\mathrm{i}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}0&-1\\ 1&0\\ \end{bmatrix}.$
The other irreducible representations of $G$ are $1$-dimensional, and can be labeled by the elements of the subgroup $\widehat{G}:=\{\pm 1\}\times\{\pm 1,\pm\mathrm{i}\}$ of $K^{\times}\times K^{\times}$ as follows: identify $\chi=(\chi_{1},\chi_{2})\in\widehat{G}$ with the representation $\chi:G\to K^{\times}$ given by
$\chi:a\mapsto\chi_{1},\quad b\mapsto\chi_{2}.$
Denote by $W_{j}$ the vector space $K^{2}$ endowed with the representation $\psi_{j}$ for $j=1,2$, and for $\chi\in\widehat{G}$ denote by $U_{\chi}$ the vector space $K$ endowed with the representation $\chi$. Set $U:=\bigoplus_{\chi\in\widehat{G}}U_{\chi}$. Note that none of $\psi_{1}$ or $\psi_{2}$ is faithful: $\ker(\psi_{1})=\langle b^{2}\rangle$, $G/\ker(\psi_{1})\cong\mathrm{D}_{8}$, and $\psi_{1}$ is the lift to $G$ of the only $2$-dimensional representation of $\mathrm{D}_{8}$. On the other hand, $\ker(\psi_{2})=\langle a^{2}b^{2}\rangle$, $G/\ker(\psi_{2})\cong\mathrm{Dic}_{8}$, and $\psi_{2}$ is the lift to $G$ of the only $2$-dimensional irreducible representation of the quaternion group. The coordinate functions on $W_{1}$ are denoted by $x_{1},x_{2}$, the coordinate functions on $W_{2}$ are $y_{1},y_{2}$, and the coordinate function on $U_{\chi}$ is $t_{\chi}$ for $\chi\in\widehat{G}$. It is well known and easy to see that
$K[x_{1},x_{2}]^{G}$ $=K[x_{1}x_{2},x_{1}^{4}+x_{2}^{4}],$
$K[y_{1},y_{2}]^{G}$ $=K[(y_{1}y_{2})^{2},\ y_{1}^{4}+y_{2}^{4},\ y_{1}y_{2}(y_{1}^{4}-y_{2}^{4})].$
Consider the following relative invariants (their index indicates their weight):
$r_{(-1,1)}:=x_{1}^{2}+x_{2}^{2},\quad r_{(-1,-1)}:=x_{1}^{2}-x_{2}^{2},\quad r_{(1,-1)}:=r_{(-1,1)}r_{(-1,-1)}=x_{1}^{4}-x_{2}^{4}$
$s_{(1,-1)}:=y_{1}y_{2},\quad s_{(-1,1)}:=y_{1}^{2}+y_{2}^{2},\quad s_{(-1,-1)}:=y_{1}^{2}-y_{2}^{2}.$
Using the above observations, one can prove by arguments similar to the proof of Proposition 6.1 the following:
Proposition 8.1.
For $i=1,2$ we have $\beta(G,W_{i}\oplus U)=6$.
Proposition 8.2.
We have $\beta(G,W_{1}\oplus W_{2})\leq 6$.
Proof.
We know that $\beta^{K}(G)\leq 7$ by [15, Proposition 3.1]. On the other hand, the element $a^{2}\in G$ acts on $W_{1}\oplus W_{2}$ as multiplication by the scalar $-1$, hence any homogeneous element of $K[W_{1}\oplus W_{2}]^{G}$ has even degree. It follows that $\beta(G,W_{1}\oplus W_{2})$ is an even number, less than or equal to $7$. Consequently, $\beta(G,W_{1}\oplus W_{2})\leq 6$. ∎
Proposition 8.3.
We have $\beta_{\mathrm{sep}}(W_{1}\oplus W_{2}\oplus U_{(1,\mathrm{i})})\leq 6$.
Proof.
Let $v=(w_{1},w_{2},u)\in V$ and $v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2},u^{\prime})\in V$ such that $f(v)=f(v^{\prime})$ for all $f\in K[V]^{G}$ with $\deg(f)\leq 6$ (here $w_{i},\ w^{\prime}_{i}\in W_{i}$ and $u,u^{\prime}\in U$). We need to show that $v$ and $v^{\prime}$ have the same $G$-orbit. By Proposition 8.2, $(w_{1},w_{2})$ and $(w^{\prime}_{1},w^{\prime}_{2})$ are in the same $G$-orbit. Replacing $v$ by an appropriate element in its $G$-orbit, we may assume that $w_{1}=w^{\prime}_{1}$ and $w_{2}=w^{\prime}_{2}$. If $w_{1}=0$ or $w_{2}=0$, then we have $G\cdot v=G\cdot v^{\prime}$ by Proposition 8.1. From now on we assume that none of $w_{1}$ or $w_{2}$ is zero. Set
$r^{(1)}_{(1,-\mathrm{i})}:=x_{1}y_{2}+\mathrm{i}x_{2}y_{1},\quad r^{(2)}_{(1,-\mathrm{i})}:=x_{1}y_{1}^{3}-\mathrm{i}x_{2}y_{2}^{3},\quad r_{(1,\mathrm{i})}:=x_{1}y_{2}-\mathrm{i}x_{2}y_{1}.$
Then $r^{(1)}_{(1,-\mathrm{i})}$, $r^{(2)}_{(1,-\mathrm{i})}$, $t^{3}$ are relative invariants of weight $(1,-\mathrm{i})\in\widehat{G}$, whereas $r_{(1,\mathrm{i})}$, $t$ are relative invariants of weight $(1,\mathrm{i})$. Therefore we have the invariants
$f_{1}:=t^{4},\quad f_{2}:=r^{(1)}_{(1,-\mathrm{i})}t,\quad f_{3}:=r^{(2)}_{(1,-\mathrm{i})}t,\quad f_{4}:=r_{(1,\mathrm{i})}t^{3}.$
All of the above invariants have degree at most $5$. By assumption, $f_{j}(v)=f_{j}(v^{\prime})$ for $j=1,2,3,4$. The equality $t^{4}(v)=t^{4}(v^{\prime})$ implies that $u^{\prime}\in\{u,-u,\mathrm{i}u,-\mathrm{i}u\}$. In particular, if $u=0$, then $u^{\prime}=0$, and hence $v=v^{\prime}$. It remains to deal with the case $u\neq 0$. Moreover, $u=u^{\prime}$ if and only if $t(v)=t(v^{\prime})$ if and only if $t^{3}(v)=t^{3}(v^{\prime})$. Suppose for contradiction that $u\neq u^{\prime}$. Then $f_{j}(v)=f_{j}(v^{\prime})$ for $j=2,3,4$ and $(w_{1},w_{2})=(w^{\prime}_{1},w^{\prime}_{2})$ imply that
$0=r^{(1)}_{(1,-\mathrm{i})}(v)=r^{(2)}_{(1,-\mathrm{i})}(v)=r_{(1,\mathrm{i})}(v).$
From $0=r^{(1)}_{(1,-\mathrm{i})}(v)=r_{(1,\mathrm{i})}(v)$ we conclude $0=x_{1}(v)y_{2}(v)=x_{2}(v)y_{1}(v)$. Taking into account that $w_{1}\neq 0$ and $w_{2}\neq 0$, we conclude that $0=x_{1}(v)=y_{1}(v)$ or $0=x_{2}(v)=y_{2}(v)$. Then $r^{(2)}_{(1,-\mathrm{i})}(v)=0$ yields that both $x_{1}(v)y_{1}(v)$ and $x_{2}(v)y_{2}(v)$ are zero. We deduce $w_{1}=0$ or $w_{2}=0$, a contradiction. ∎
Proposition 8.4.
We have $\beta_{\mathrm{sep}}(G,W_{1}\oplus W_{2}\oplus U_{(-1,1)})\leq 6$ and $\beta_{\mathrm{sep}}(G,W_{1}\oplus W_{2}\oplus U_{(1,-1)})\leq 6$.
Proof.
First we deal with $V:=W_{1}\oplus W_{2}\oplus U_{(-1,1)}$. Let $v=(w_{1},w_{2},u)\in V$ and $v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2},u^{\prime})\in V$ such that $f(v)=f(v^{\prime})$ for all $f\in K[V]^{G}$ with $\deg(f)\leq 6$. We need to show that $v$ and $v^{\prime}$ have the same $G$-orbit. In the same way as in the proof of Proposition 8.3, we may assume that $w_{1}=w^{\prime}_{1}$, $w_{2}=w^{\prime}_{2}$ and none of $w_{1}$ or $w_{2}$ is zero. It follows that
$r_{(-1,1)}=x_{1}^{2}+x_{2}^{2},\quad r_{(-1,-1)}s_{(1,-1)}=(x_{1}^{2}-x_{2}^{2})y_{1}y_{2},\quad s_{(-1,1)}=y_{1}^{2}+y_{2}^{2}$
can not all vanish at $v$. Suppose that $f$ is one of the above polynomials, with $f(v)\neq 0$. Since $f$ is a relative invariant of weight $(-1,1)$ of degree at most $4$, we get that $ft\in K[V]^{G}$ has degree at most $5$. Moreover, the equality $(ft)(v)=(ft)(v^{\prime})$ implies $u=u^{\prime}$, hence $v=v^{\prime}$.
The proof for $V=W_{1}\oplus W_{2}\oplus U_{(1,-1)}$ is similar, the only difference is that we use the relative invariants
$s_{(1,-1)}=y_{1}y_{2},\quad s_{(-1,1)}r_{(-1,-1)}=(y_{1}^{2}+y_{2}^{2})(x_{1}^{2}-x_{2}^{2}),\quad s_{(-1,-1)}r_{(-1,1)}=(y_{1}^{2}-y_{2}^{2})(x_{1}^{2}+x_{2}^{2}).$
Proposition 8.5.
We have $\mu(G)\leq 2$, and hence $\kappa(G)\leq 3$.
Proof.
Let $S$ be an intersection independent set of subgroups of $G$ with $|S|=\mu(G)$. Clearly $G\notin S$. If $S$ contains a cyclic subgroup of prime power order, then $|S|\leq 2$ by Lemma 3.8. From now on assume that $S$ contains no cyclic subgroup of prime power order. All such proper subgroups of $G$ contain the subgroup $\langle a^{2},b^{2}\rangle$. If $\langle a^{2},b^{2}\rangle\in S$, then $|S|=1$. Otherwise $S$ consists of two order $8$ subgroups of $G$. In either case, $|S|\leq 2$. The inequality $\mu(G)\leq 2$ is proved, and hence $\kappa(G)\leq 1+\mu(G)\leq 3$ holds by (9). ∎
Theorem 8.6.
Assume that $K$ has an element of multiplicative order $8$. Then we have $\beta_{\mathrm{sep}}(\mathrm{C}_{4}\rtimes\mathrm{C}_{4})=6$.
Proof.
The factor group of $G$ modulo its central normal subgroup $\langle a^{2}b^{2}\rangle$ is isomorphic to the quaternion group $\mathrm{Dic}_{8}$ of order $8$. Therefore $\beta_{\mathrm{sep}}^{K}(G)\geq\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{8})=6$ (in fact we have $\beta_{\mathrm{sep}}(G,W_{2})=6$, as we saw in Section 4).
To prove the reverse inequality, by Lemma 3.1 we may assume that $K$ is large enough so that Lemma 3.6 applies and we can reduce to the study of multiplicity-free representations. By Proposition 8.5 we have $\kappa(G)\leq 3$, and therefore by Lemma 3.7 it is sufficient to show that $\beta_{\mathrm{sep}}(V)\leq 6$ if $V$ is the direct sum of three pairwise non-isomorphic $G$-modules. Assume that $V$ is such a $G$-module. By Proposition 8.2, Proposition 8.1 we know that $\beta_{\mathrm{sep}}(G,V)\leq\beta(G,V)\leq 6$, unless $V=W_{1}\oplus W_{2}\oplus U_{\chi}$ for some non-trivial character $\chi\in\widehat{G}$. The cases when $\chi\in\{(1,\mathrm{i}),\ (-1,1),\ (1,-1)\}$ are covered by Proposition 8.3, Proposition 8.4. The remaining representations are of the form $\rho\circ\alpha$, where $\rho$ is one of the above three representations, and $\alpha$ is an automorphism of $G$. So the proof is complete by Lemma 3.11. ∎
8.2. The group $\mathrm{C}_{5}\rtimes\mathrm{C}_{4}$
In this section
$G=\mathrm{C}_{5}\rtimes\mathrm{C}_{4}=\langle a,b\mid a^{5}=b^{4}=1,\ bab^{-1}=a^{2}\rangle.$
Assume that $K$ has an element $\xi$ of multiplicative order $20$. Then $\omega:=\xi^{4}$ has multiplicative order $5$, and $\mathrm{i}:=\xi^{5}$ has multiplicative order $4$. Consider the following irreducible $4$-dimensional representation of $G$:
$\psi:a\mapsto\begin{bmatrix}\omega&0&0&0\\ 0&\omega^{2}&0&0\\ 0&0&\omega^{4}&0\\ 0&0&0&\omega^{3}\end{bmatrix},\quad b\mapsto\begin{bmatrix}0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ 1&0&0&0\end{bmatrix}.$
The other irreducible representations of $G$ are $1$-dimensional, and can be labelled by the group $\widehat{G}=\{\pm\mathrm{i},\pm 1\}\leq K^{\times}$, where $\chi\in\widehat{G}$ is identified with the representation
$\chi:a\mapsto 1,\ b\mapsto\chi$
(note that $\langle a\rangle$ is the commutator subgroup of $G$). Write $W$ for the vector space $K^{4}$ endowed with the representation $\psi$, and for $\chi\in\widehat{G}$ denote by $U_{\chi}$ the vector space $K$ endowed with the representation $\chi$, and set $U:=\bigoplus_{\chi\in\widehat{G}}U_{\chi}$.
Set
$h_{0}:=x_{1}x_{3}-x_{2}x_{4},\qquad h_{1}:=x_{1}x_{2}^{2}+x_{3}x_{4}^{2},\qquad h_{2}:=x_{2}x_{3}^{2}+x_{4}x_{1}^{2},$
$h_{3}:=x_{1}^{3}x_{2}+x_{3}^{3}x_{4},\qquad h_{4}:=x_{2}^{3}x_{3}+x_{4}^{3}x_{1},$
and consider the following elements in $K[W]^{G}$:
$f_{1}:=x_{1}x_{3}+x_{2}x_{4},\qquad f_{2}:=x_{1}x_{2}x_{3}x_{4},$
$f_{3}:=h_{1}+h_{2},\qquad f_{4}:=h_{1}h_{2},\qquad f_{5}:=(h_{1}-h_{2})h_{0},$
$f_{6}:=h_{3}+h_{4},\qquad f_{7}:=x_{1}^{5}+x_{2}^{5}+x_{3}^{5}+x_{4}^{5},$
$f_{8}:=x_{1}^{6}x_{3}+x_{1}x_{3}^{6}+x_{2}^{6}x_{4}+x_{2}x_{4}^{6}$
$f_{9}:=(h_{3}-h_{4})h_{0}.$
The following result was obtained using the CoCalc platform [43]:
Proposition 8.7.
Suppose additionally that $K$ has characteristic zero. Then the elements $f_{1},\dots,f_{8}$ form a minimal homogeneous generating system of the $K$-algebra $K[W]^{G}$.
It turns out that the elements of degree at most $6$ from the above generating system are sufficient to separate $G$-orbits in $W$:
Proposition 8.8.
If for $v,v^{\prime}\in W$ we have that $f_{i}(v)=f_{i}(v^{\prime})$ for all $i=1,\dots,7$, then $f_{8}(v)=f_{8}(v^{\prime})$. In particular, if $K$ has characteristic zero, then the $G$-invariants $f_{1},\dots,f_{7}$ separate the $G$-orbits in $W$.
Proof.
Let $v,v^{\prime}\in W$ such that $f_{j}(v)=f_{j}(v^{\prime})$ for $j=1,\dots,7$. In order to prove $f_{8}(v)=f_{8}(v^{\prime})$, we will prove the existence of a chain of subsets
$\{f_{1},\dots,f_{7}\}=S^{(0)}\subset S^{(1)}\subset\dots\subset S^{(k)}\subset K[V]$
and elements $v^{\prime}=w^{(0)},\dots,w^{(k)}$ in the $G$-orbit of $v^{\prime}$ such that
  • $f(v)=f(w^{(j)})$ for all $f\in S^{(j)}$ and all $j=0,\dots,k$;
  • $f_{8}\in S^{(k)}.$
Then $f_{8}(v)=f_{8}(w^{(k)})$, and since $f_{8}$ is a $G$-invariant and $G\cdot v^{\prime}=G\cdot w^{(k)}$, we have $f_{8}(w^{(k)})=f_{8}(v^{\prime})$ (and hence $G\cdot v=G\cdot v^{\prime}$ by Proposition 8.7 when $\mathrm{char}(K)=0$). We shall denote by $T^{(j)}$ the subalgebra of $K[V]$ generated by $S^{(j)}$; obviously, $f(v)=f(w^{(j)})$ holds for all $f\in T^{(j)}$ as well, so any element of $T^{(j)}$ can be added to $S^{(j)}$.
From $f_{1}(v)=f_{1}(v^{\prime})$ and $f_{2}(v)=f_{2}(v^{\prime})$ we deduce
$\{(x_{1}x_{3})(v),(x_{2}x_{4})(v)\}=\{(x_{1}x_{3})(v^{\prime}),(x_{2}x_{4})(v^{\prime})\}.$
So either $(x_{1}x_{3})(v)=(x_{1}x_{3})(v^{\prime})$ and $(x_{2}x_{4})(v)=(x_{2}x_{4})(v^{\prime})$, or $(x_{1}x_{3})(v)=(x_{1}x_{3})(b\cdot v^{\prime})$ and $(x_{2}x_{4})(v)=(x_{2}x_{4})(b\cdot v^{\prime})$. Therefore we can take $w^{(1)}=v^{\prime}$ or $w^{(1)}=b\cdot v^{\prime}$, and
$S^{(1)}=S^{(0)}\cup\{x_{1}x_{3},x_{2}x_{4}\}.$
In particular, this implies $h_{0}\in T^{(1)}$.
Case I.: $f_{2}(v)\neq 0$ and $h_{0}(v)\neq 0$. From $f_{3}(v)=f_{3}(w^{(1)})$, $f_{5}(v)=f_{5}(w^{(1)})$, and $h_{0}(v)=h_{0}(w^{(1)})\neq 0$ we infer that $h_{1}(v)=h_{1}(w^{(1)})$ and $h_{2}(v)=h_{2}(w^{(1)})$. The product of the two summands of $h_{1}$ (respectively $h_{2}$) is $(x_{1}x_{3})(x_{2}x_{4})^{2}$ (respectively $(x_{1}x_{3})^{2}(x_{2}x_{4})$), which belongs to the subalgebra $T^{(1)}$ of $K[V]$. It follows that
$\{(x_{1}x_{2}^{2})(v),(x_{3}x_{4}^{2})(v)\}=\{(x_{1}x_{2}^{2})(w^{(1)}),(x_{3}x_{4}^{2})(w^{(1)})\}\qquad\text{ and }$
$\{(x_{2}x_{3}^{2})(v),(x_{4}x_{1}^{2})(v)\}=\{(x_{2}x_{3}^{2})(w^{(1)}),(x_{4}x_{1}^{2})(w^{(1)})\}.$
Similarly, from $f_{6}(v)=f_{6}(w^{(1)})$ and $f_{9}(v)=f_{9}(w^{(1)})$ (note that $f_{9}=f_{1}f_{6}-2f_{4}$, hence $f_{9}(v)=f_{9}(v^{\prime})=f_{9}(w^{(1)})$) we get
$\{(x_{1}^{3}x_{2})(v),(x_{3}^{3}x_{4})(v)\}=\{(x_{1}^{3}x_{2})(w^{(1)}),(x_{3}^{3}x_{4})(w^{(1)})\}$
(and also $\{(x_{2}^{3}x_{3})(v),(x_{4}^{3}x_{1})(v)\}=\{(x_{2}^{3}x_{3})(w^{(1)}),(x_{4}^{3}x_{1})(w^{(1)})\}$, but we do not use it below). Note that the elements of $S^{(1)}$ are $b^{2}$-invariant, whereas $b^{2}$ interchanges $x_{1}x_{2}^{2}$ and $x_{3}x_{4}^{2}$, $x_{2}x_{3}^{2}$ and $x_{4}x_{1}^{2}$, $x_{1}^{3}x_{2}$ and $x_{3}^{3}x_{4}$. It follows that with $w^{(2)}=w^{(1)}$ or $w^{(2)}=b^{2}\cdot w^{(1)}$ one of the following sets can be taken as $S^{(2)}$:
  1. (i)
    $S^{(2)}=S^{(1)}\cup\{x_{1}x_{2}^{2},x_{3}x_{4}^{2},x_{2}x_{3}^{2},x_{4}x_{1}^{2}\}$
  2. (ii)
    $S^{(2)}=S^{(1)}\cup\{x_{1}x_{2}^{2},x_{3}x_{4}^{2},x_{1}^{3}x_{2},x_{3}^{3}x_{4}\}$
  3. (iii)
    $S^{(2)}=S^{(1)}\cup\{x_{2}x_{3}^{2},x_{4}x_{1}^{2},x_{1}^{3}x_{2},x_{3}^{3}x_{4}\}$
In case (i), we have $(x_{1}x_{2}^{2})^{2}(x_{2}x_{3}^{2})=x_{2}^{5}(x_{1}x_{3})^{2}\in T^{(2)}$. Since $0\neq(x_{1}x_{3})(v)=(x_{1}x_{3})(w^{(2)})$, we conclude that $x_{2}(v)^{5}=x_{2}(w^{(2)})^{5}$. For an appropriate $s\in\{0,1,2,3,4\}$, we have $x_{2}(v)=x_{2}(a^{s}\cdot w^{(2)})$. Set $w^{(3)}:=a^{s}\cdot w^{(2)}$. Since the elements of $S^{(2)}$ are $\langle a\rangle$-invariant, we may take
$S^{(3)}:=S^{(2)}\cup\{x_{2}\}.$
Then from $(x_{1}x_{2}^{2})(v)=(x_{1}x_{2}^{2})(w^{(3)})$ and $x_{2}(v)=x_{2}(w^{(3)})\neq 0$ we get $x_{1}(v)=x_{1}(w^{(3)})$. Given that, from $(x_{1}x_{3})(v)=(x_{1}x_{3})(w^{(3)})$ and $x_{1}(v)=x_{1}(w^{(3)})\neq 0$ we infer $x_{3}(v)=x_{3}(w^{(3)})$. Finally, from $(x_{2}x_{4})(v)=(x_{2}x_{4})(w^{(3)})$ and $x_{2}(v)=x_{2}(w^{(3)})\neq 0$ we get $x_{4}(v)=x_{4}(w^{(3)})$. So keeping $w^{(4)}:=w^{(3)}$ we can take
$S^{(4)}:=S^{(3)}\cup\{x_{1},x_{3},x_{4}\}.$
Then $f_{8}\in T^{(4)}$, hence with $w^{(5)}=w^{(4)}$ we can take $S^{(5)}=S^{(4)}\cup\{f_{8}\}$, and we are done. The cases when $S^{(2)}$ is as in (ii) or (iii) can be dealt with similarly.
Case II.: $h_{0}(v)=0$. Then we have
(24) $(x_{1}x_{3})(w^{(1)})=(x_{1}x_{3})(v)=(x_{2}x_{4})(v)=(x_{2}x_{4})(w^{(1)}).$
Observe that
(25) $f_{8}=(x_{1}x_{3})(x_{1}^{5}+x_{3}^{5})+(x_{2}x_{4})(x_{2}^{5}+x_{4}^{5}).$
Therefore by (24) we obtain
$f_{8}(v)=(x_{1}x_{3})(v)f_{7}(v)=(x_{1}x_{3})(w^{(1)})f_{7}(w^{(1)})=f_{8}(w^{(1)}),$
so with $w^{(2)}=w^{(1)}$ we can take $S^{(2)}=S^{(1)}\cup\{f_{8}\}$.
Case III.: $h_{0}(v)\neq 0$ and $f_{2}(v)=0$. By symmetry, we may assume that $x_{1}(v)=0$. From $(x_{1}x_{3})(v)=(x_{1}x_{3})(w^{(1)})$ we deduce that $x_{1}(w^{(1)})=0$ or $x_{3}(w^{(3)})=0$. Since $b^{2}$ interchanges $x_{1}$ and $x_{3}$, and fixes the elements of $S^{(1)}$, we may take $w^{(2)}=w^{(1)}$ or $w^{(2)}=b^{2}\cdot w^{(1)}$ and
$S^{(2)}=S^{(1)}\cup\{x_{1}\}.$
Moreover, we have
(26) $x_{1}(v)=0=x_{1}(w^{(2)})$
and hence
(27) $0\neq h_{0}(v)=-(x_{2}x_{4})(v)=-(x_{2}x_{4})(w^{(2)})=h_{0}(w^{(2)}).$
We have
$h_{1}(v)-h_{2}(v)=\frac{f_{5}(v)}{h_{0}(v)}=\frac{f_{5}(w^{(2)})}{h_{0}(w^{(2)})}=h_{1}(w^{(2)})-h_{2}(w^{(2)}).$
Together with $h_{1}(v)+h_{2}(v)=f_{3}(v)=f_{3}(w^{(2)})=h_{1}(w^{(2)})+h_{2}(w^{(2)})$ and with (26) this implies
$(x_{3}x_{4}^{2})(v)=h_{1}(v)=h_{1}(w^{(2)})=(x_{3}x_{4}^{2})(w^{(2)})\text{ and }$
$(x_{2}x_{3}^{2})(v)=h_{2}(v)=h_{2}(w^{(2)})=(x_{2}x_{3}^{2})(w^{(2)}).$
Therefore with $w^{(3)}=w^{(2)}$ we can take
$S^{(3)}=S^{(2)}\cup\{x_{2}x_{3}^{2},x_{3}x_{4}^{2}\}.$
The equality
$(x_{2}x_{3}^{2})^{2}(x_{3}x_{4}^{2})=(x_{3}^{5})(x_{2}x_{4})^{2}\in T^{(3)}$
with $(x_{2}x_{4})(v)=(x_{2}x_{4})(w^{(3)})\neq 0$ (see (27)) imply that
(28) $x_{3}^{5}(v)=x_{3}^{5}(w^{(3)}).$
So we may take $w^{(4)}=w^{(3)}$ and
$S^{(4)}:=S^{(3)}\cup\{x_{3}^{5}\}.$
Then $f_{7}$, $x_{1}$, $x_{3}^{5}$ all belong to $S^{(4)}$, implying that $x_{2}^{5}+x_{4}^{5}=f_{7}-x_{1}^{5}-x_{3}^{5}\in T^{(4)}$. So we may take $w^{(5)}=w^{(4)}$ and
$S^{(5)}=S^{(4)}\cup\{x_{2}^{5}+x_{4}^{5}\}$
Thus $x_{1}x_{3}$, $x_{2}x_{4}$, $x_{1}^{5}+x_{3}^{5}$, $x_{2}^{5}+x_{4}^{5}$ all belong to $T^{(5)}$, hence equality (25) shows that $f_{8}\in T^{(5)}$. Therefore $f_{8}(v)=f_{8}(w^{(5)})$, and we are done. ∎
Theorem 8.9.
Assume that $K$ has characteristic 00, and contains an element of multiplicative order $20$. Then we have the equality $\beta_{\mathrm{sep}}^{K}(\mathrm{C}_{5}\rtimes\mathrm{C}_{4})=6$.
Proof.
The subgroup $\langle a,b^{2}\rangle$ of $G$ is isomorphic to $\mathrm{D}_{10}$, the dihedral group of order $10$, hence by equation (2) and Theorem 2.1 we have the inequality $\beta_{\mathrm{sep}}^{K}(G)\geq 6$.
In view of Lemma 3.6, to prove the reverse inequality take $v=(w,u),v^{\prime}=(w^{\prime},u^{\prime})\in V=W\oplus U$ such that
(29) $f(v)=f(v^{\prime})\text{ holds for all homogeneous }f\in K[V]^{G}\text{ with }\deg(f)\leq 6.$
We have to show that $G\cdot v=G\cdot v^{\prime}$. By Proposition 8.8 we have $G\cdot w=G\cdot w^{\prime}$, so replacing $v^{\prime}$ by an appropriate element in its $G$-orbit we may assume that $w=w^{\prime}$. Moreover, $G/G^{\prime}=G/\langle a\rangle\cong\langle b\rangle\cong\mathrm{C}_{4}$, and $\mathsf{D}(\mathrm{C}_{4})=4$, so $G\cdot u=G\cdot u^{\prime}$ as well, implying that $u^{\prime}_{\mathrm{i}}\in\{\pm u_{\mathrm{i}},\pm\mathrm{i}u_{\mathrm{i}}\}$, $u^{\prime}_{-1}=\pm u_{-1}$, and $u^{\prime}_{1}=u_{1}$. So it is sufficient to deal with the case when $w\neq 0$ and $u_{\chi}\neq 0$ for some $\chi\in\widehat{G}\setminus\{1\}$.
The eigenvalues of $\psi(g)$ for a non-identity element $g\in\langle a\rangle$ differ from $1$, therefore $\mathrm{Stab}_{G}(w)\cap\langle a\rangle=\{1_{G}\}$, and hence $\mathrm{Stab}_{G}(w)$ is isomorphic to a subgroup of $\mathrm{C}_{4}$.
Case I.: $\mathrm{Stab}_{G}(w)\cong\mathrm{C}_{4}$. Then $\mathrm{Stab}_{G}(w)G^{\prime}=G$, hence from $G\cdot u=G\cdot u^{\prime}$ it follows that there exists an element $h\in\mathrm{Stab}_{G}(w)$ with $h\cdot u=u^{\prime}$. Thus we have $h\cdot(w,u)=(w,u^{\prime})$, and we are done.
Case II.: $\mathrm{Stab}_{G}(w)\cong\mathrm{C}_{2}$. Then $\mathrm{Stab}_{G}(w)$ is conjugate in $G$ to $\langle b^{2}\rangle$, so $\mathrm{Stab}_{G}(w)=g\langle b^{2}\rangle g^{-1}$ for some $g\in G$. Replacing $(w,u)$ and $(w,u^{\prime})$ by $g^{-1}\cdot(w,u)$ and $g^{-1}\cdot(w,u^{\prime})$ we may assume that $\mathrm{Stab}_{G}(w)=\langle b^{2}\rangle$. It follows that $w=[\nu,\eta,\nu,\eta]^{T}\in K^{4}=V$, where $\nu\neq\eta$. Note that if $m$ is a $b^{2}$-invariant monomial in $K[U]$ of degree at most $2$, then either it is $b$-invariant, or $b\cdot m=-m$. So if $m$ is not $b$-invariant, then both $h_{0}m$ and $(h_{1}-h_{2})m$ are $G$-invariants of degree at most $5$. Note that $h_{0}(v)=\nu^{2}-\eta^{2}$ and $(h_{1}-h_{2})(v)=2\nu\eta(\nu-\eta)$, so $\nu\neq\eta$ implies that $h_{0}(v)$ and $(h_{1}-h_{2})(v)$ can not be simultaneously zero. We infer by (29) that $m(u)=m(u^{\prime})$ holds for all $\langle b^{2}\rangle$-invariant monomials in $K[U]$ of degree at most $2$. As $\mathsf{D}(\mathrm{C}_{2})=2$, we deduce that $u$ and $u^{\prime}$ belong to the same orbit under $\langle b^{2}\rangle=\mathrm{Stab}_{G}(w)$, implying in turn that $v=(w,u)$ and $v^{\prime}=(w,u^{\prime})$ belong to the same $G$-orbit.
Case III.: $\mathrm{Stab}_{G}(w)=\{1_{G}\}$. Clearly it is sufficient to show that $u_{\chi}=u^{\prime}_{\chi}$ for all $\chi\in\widehat{G}\setminus\{1\}=\{-1,\pm\mathrm{i}\}$. First we deal with $\chi=-1$. Using the CoCalc platform [43] we obtained that a minimal homogeneous generating system of $K[W\oplus U_{-1}]^{G}$ consists of the generators $f_{1},\dots,f_{8}$ of $K[V]^{G}$ from Proposition 8.7, together with $t_{-1}^{2}$, $h_{0}t_{-1}$, $(h_{1}-h_{2})t_{-1}$, $(h_{3}-h_{4})t_{-1}$, $(x_{1}^{5}-x_{2}^{5}+x_{3}^{5}-x_{4}^{5})t_{-1}$. So all the generators involving $t_{-1}$ have degree at most $6$. This implies that $G\cdot(w,u_{-1})=G\cdot(w,u^{\prime}_{-1})$. Taking into account that the stabilizer of $w$ is trivial, this means that $u_{-1}=u^{\prime}_{-1}$.
Next we deal with $\chi=-\mathrm{i}$ and show that that $u_{-\mathrm{i}}=u^{\prime}_{-\mathrm{i}}$ (the argument for $u_{\mathrm{i}}=u^{\prime}_{\mathrm{i}}$ is obtained by obvious modification). Since we know already $G\cdot u_{-\mathrm{i}}=G\cdot u^{\prime}_{-\mathrm{i}}$, i.e. $t_{-\mathrm{i}}^{4}(u)=t_{-\mathrm{i}}^{4}(u^{\prime})$, our claim is equivalent to
(30) $t_{-\mathrm{i}}(u)=t_{-\mathrm{i}}(u^{\prime})\text{ or }t_{-\mathrm{i}}^{3}(u)=t_{-\mathrm{i}}^{3}(u^{\prime}).$
The latter holds if there exists a relative invariant $f\in K[W]^{G,\mathrm{i}}$ (respectively, $f\in K[W]^{G,-\mathrm{i}}$) with $\deg(f)\leq 5$ (respectively, $\deg(f)\leq 3$) with $f(w)\neq 0$, because then $ft_{-\mathrm{i}}$ (respectively, $ft_{-\mathrm{i}}^{3}$) is a $G$-invariant of degree a most $6$, and therefore $f(w)t_{-\mathrm{i}}(u)=f(w)t_{-\mathrm{i}}(u^{\prime})$ (respectively, $f(w)t_{\mathrm{i}}^{3}(u)=f(w)t_{\mathrm{i}}^{3}(u^{\prime})$), so (30) holds. Suppose for contradiction that $w$ belongs to the common zero locus of $K[W]^{G,\mathrm{i}}_{\leq 5}$ and $K[W]^{G,-\mathrm{i}}_{\leq 3}$. Set
$k_{1}:=x_{1}x_{2}^{2}-x_{3}x_{4}^{2},\qquad k_{2}:=x_{2}x_{3}^{2}-x_{4}x_{1}^{2}$
$k_{3}:=x_{1}^{3}x_{2}-x_{3}^{3}x_{4},\qquad k_{4}:=x_{2}^{3}x_{3}-x_{4}^{3}x_{1}$
$k_{5}:=x_{1}^{5}-x_{3}^{5},\qquad k_{6}:=x_{2}^{5}-x_{4}^{5}.$
Then
$k_{1}-\mathrm{i}k_{2},\quad k_{3}-\mathrm{i}k_{4},\quad k_{5}-\mathrm{i}k_{6}\in K[V]^{G,\mathrm{i}}_{\leq 5}$
$k_{1}+\mathrm{i}k_{2}\in K[V]^{G,-\mathrm{i}}_{\leq 3}.$
So the above four relative invariants all vanish at $w$. In particular, it follows that $k_{1}(w)=0$ and $k_{2}(w)=0$, i.e.
(31) $(x_{1}x_{2}^{2})(w)=(x_{3}x_{4}^{2})(w)\text{ and }(x_{2}x_{3}^{2})(w)=(x_{4}x_{1}^{2})(w).$
Case III.a: $x_{j}(w)=0$ for some $j\in\{1,2,3,4\}$; by symmetry we may assume that $x_{1}(w)=0$. By (31) we conclude $x_{3}(w)=0$ or $x_{4}(w)=0=x_{2}(w)$. In the latter case by $(k_{5}-\mathrm{i}k_{6})(w)=0$ we deduce $x_{3}(w)=0$, leading to the contradiction that $w=0$. So $x_{1}(w)=0=x_{3}(w)$. Then $(k_{5}-\mathrm{i}k_{6})(w)=0$ imply $x_{2}^{5}(w)=x_{4}^{5}(w)$, so (as $w\neq 0$) we have that $x_{4}(w)=\omega^{j}x_{2}(w)$ for some $j\in\{0,1,2,3,4\}$. Then one can easily check that the stabilizer of $w$ is a conjugate of the subgroup $\langle b^{2}\rangle$ of $G$, a contradiction.
Case III.b: $(x_{1}x_{2}x_{3}x_{4})(w)\neq 0$. From (31) we deduce
$x_{3}(w)=x_{1}(w)x_{2}(w)^{2}x_{4}(w)^{-2}\text{ and then }x_{4}(w)=x_{2}(w)(x_{1}(w)x_{2}(w)^{2}x_{4}(w)^{-2})^{2}x_{1}(w)^{-2}.$
The latter equality implies
$x_{2}(w)^{5}=x_{4}(w)^{5},$
which together with $(k_{5}-\mathrm{i}k_{6})(w)=0$ yields
$x_{1}(w)^{5}=x_{3}(w)^{5}.$
So there exist unique $j,k\in\{0,1,2,3,4\}$ with
$x_{3}(w)=\omega^{j}x_{1}(w)\text{ and }x_{4}(w)=\omega^{k}x_{2}(w).$
From (31) it follows that
$(j,k)\in\{(0,0),(1,2),(2,4),(3,1),(4,3).$
If $(j,k)=(0,0)$, then $b^{2}\in\mathrm{Stab}_{G}(w)$, a contradiction. If $(j,k)=(1,2)$, then $a^{4}b^{2}\in\mathrm{Stab}_{G}(w)$, a contradiction. Similarly we get to a contradiction for all other possible $(j,k)$. ∎
8.3. The group $\mathrm{C}_{7}\rtimes\mathrm{C}_{3}$
In this section
$G=\mathrm{C}_{7}\rtimes\mathrm{C}_{3}=\langle a,b\mid a^{7}=b^{3}=1_{G},\quad bab^{-1}=a^{2}\rangle$
is the only non-abelian group of order $21$.
It is proved in [9, Theorem 4.1] that for any prime $p$ congruent to $1$ modulo $3$, the separating Noether number of the non-abelian semidirect product $\mathrm{C}_{p}\rtimes\mathrm{C}_{3}$ equals $p+1$, under the assumption that the base field $K$ contains an element of multiplicative order $3p$ ([9] works over an algebraically closed base field, but the representation providing the exact lower bound for the separating Noether number is defined over any field containing an element of multiplicative order $3p$, and in order to prove that $p+1$ is also an upper bound, we may pass to the algebraic closure using Lemma 3.1).
As the special case $p=7$, we get the following:
Corollary 8.10.
Assume that $K$ contains an element of multiplicative order $21$. Then we have $\beta_{\mathrm{sep}}^{K}(\mathrm{C}_{7}\rtimes\mathrm{C}_{3})=8$.
8.4. The group $\mathrm{M}_{27}$
In this section
$G=\mathrm{M}_{27}=\langle a,b\mid a^{9}=b^{3}=1,\ bab^{-1}=a^{4}\rangle\cong\mathrm{C}_{9}\rtimes\mathrm{C}_{3}$
is the non-abelian group of order $27$ with an index $3$ cyclic subgroup. Assume that $K$ contains an element $\omega$ of multiplicative order $9$, and consider the following irreducible $3$-dimensional representations of $G$:
$\psi_{1}:a\mapsto\begin{bmatrix}\omega&0&0\\ 0&\omega^{4}&0\\ 0&0&\omega^{7}\end{bmatrix},\quad b\mapsto\begin{bmatrix}0&1&0\\ 0&0&1\\ 1&0&0\end{bmatrix}.$
$\psi_{2}:a\mapsto\begin{bmatrix}\omega^{2}&0&0\\ 0&\omega^{8}&0\\ 0&0&\omega^{5}\end{bmatrix},\quad b\mapsto\begin{bmatrix}0&1&0\\ 0&0&1\\ 1&0&0\end{bmatrix}.$
The commutator subgroup of $G$ is $G^{\prime}=\langle a^{3}\rangle$, and $G/G^{\prime}\cong\mathrm{C}_{3}\times\mathrm{C}_{3}$. So the remaining irreducible representations of $G$ are $1$-dimensional and can be labeled by $\widehat{G}=\{\varepsilon,\varepsilon^{2},1\}\times\{\varepsilon,\varepsilon^{2},1\}\leq K^{\times}\times K^{\times}$, where $\varepsilon:=\omega^{3}$ has multiplicative order $3$. Identify $\chi=(\chi_{1},\chi_{2})\in\widehat{G}$ with the representation
$\chi:a\mapsto\chi_{1},\qquad b\mapsto\chi_{2}.$
Write $W_{j}$ for the vector space $K^{3}$ endowed with the representation $\psi_{j}$, and write $U_{\chi}$ for the $1$-dimensional vector space $K$ endowed by the representation $\chi$, and set $U:=\bigoplus_{\chi\in\widehat{G}}U_{\chi}$.
The following result was obtained using the CoCalc platform [43]:
Proposition 8.11.
Assume in addition that $\mathrm{char}(K)=0$. Then we have $\beta(G,W_{1}\oplus W_{2})=9$.
Lemma 8.12.
If $\mathrm{Stab}_{G}(v)$ is non-trivial for some $v\in W_{1}\oplus W_{2}$, then either $v=0$ (and then $\mathrm{Stab}_{G}(v)=G$), or $v\neq 0$ and then $\mathrm{Stab}_{G}(v)\in\{\langle b\rangle,\langle a^{3}b\rangle,\langle a^{6}b\rangle\}$ (in particular, then $|\mathrm{Stab}_{G}(v)|=3$). Moreover, for a non-zero $w_{1}\in W_{1}$ and $w_{2}\in W_{2}$ we have that
  • $\mathrm{Stab}_{G}(w_{1})=\langle b\rangle\iff w_{1}\in K[1,1,1]^{T}$;
  • $\mathrm{Stab}_{G}(w_{2})=\langle b\rangle\iff w_{2}\in K[1,1,1]^{T}$;
  • $\mathrm{Stab}_{G}(w_{1})=\langle a^{3}b\rangle\iff w_{1}\in K[\varepsilon^{2},\varepsilon,1]^{T}$;
  • $\mathrm{Stab}_{G}(w_{2})=\langle a^{3}b\rangle\iff w_{2}\in K[\varepsilon,\varepsilon^{2},1]^{T}$;
  • $\mathrm{Stab}_{G}(w_{1})=\langle a^{6}b\rangle\iff w_{1}\in K[\varepsilon,\varepsilon^{2},1]^{T}$;
  • $\mathrm{Stab}_{G}(w_{2})=\langle a^{6}b\rangle\iff w_{2}\in K[\varepsilon^{2},\varepsilon,1]^{T}$.
Proof.
One can check by basic linear algebra that $\psi_{j}(g)$ has the eigenvalue $1$ for some $g\in G$ and $j\in\{1,2\}$ if and only if $g$ belongs to one of the order $3$ subgroups $\langle b\rangle$, $\langle a^{3}b\rangle$, $\langle a^{6}b\rangle$ of $G$. Computing the corresponding eigenvectors we obtain the result. ∎
Recall that given a set $S\subset K[V]$ of polynomials, $\mathcal{V}(S)$ stands for the common zero locus in $V$ of the elements of $S$.
Lemma 8.13.
  • (i)
    For $\chi\in\{(\varepsilon,1),(\varepsilon^{2},1)\}$ we have
    $\mathcal{V}(f\in K[W_{1}\oplus W_{2}]^{G,\chi}\mid\deg(f)\leq 6)=\{0\}.$
  • (ii)
    For $\chi\in\{(\varepsilon,\varepsilon),(\varepsilon,\varepsilon^{2}),(\varepsilon^{2},\varepsilon),(\varepsilon^{2},\varepsilon^{2}),(1,\varepsilon),(1,\varepsilon^{2})\}$ we have
    $\mathcal{V}(f\in K[W_{1}\oplus W_{2}]^{G,\chi}\mid\deg(f)\leq 9)=\{v\in W_{1}\oplus W_{2}\mid\mathrm{Stab}_{G}(v)\neq\{1_{G}\}\}.$
Proof.
(i) For $\chi=(\varepsilon,1)$, consider the following relative invariants in $K[W_{1}\oplus W_{2}]^{G,(\varepsilon,1)}$:
$f_{(\varepsilon,1)}^{(1)}:=x_{1}x_{2}x_{3},$ $\qquad h_{(\varepsilon,1)}^{(1)}:=(y_{1}y_{2}y_{3})^{2},$
$f_{(\varepsilon,1)}^{(2)}:=x_{1}^{3}+x_{2}^{3}+x_{3}^{3},$ $\qquad h_{(\varepsilon,1)}^{(2)}:=y_{1}^{6}+y_{2}^{6}+y_{3}^{6},$
$f_{(\varepsilon,1)}^{(3)}:=x_{1}^{5}x_{3}+x_{2}^{5}x_{1}+x_{3}^{5}x_{2},$ $\qquad h_{(\varepsilon,1)}^{(3)}:=y_{1}^{2}y_{2}+y_{2}^{2}y_{3}+y_{3}^{2}y_{1}.$
It is easy to see that $0=f_{(\varepsilon,1)}^{(1)}(v)=f_{(\varepsilon,1)}^{(2)}(v)=f_{(\varepsilon,1)}^{(3)}(v)$ implies $0=x_{1}(v)=x_{2}(v)=x_{3}(v)$, and similarly $0=h_{(\varepsilon,1)}^{(1)}(v)=h_{(\varepsilon,1)}^{(2)}(v)=h_{(\varepsilon,1)}^{(3)}(v)$ implies $0=y_{1}(v)=y_{2}(v)=y_{3}(v)$. For $\chi=(\varepsilon^{2},1)$ we just need to interchange the roles of the variable sets $\{x_{1},x_{2},x_{3}\}$ and $\{y_{1},y_{2},y_{3}\}$ in the relative invariants constructed above.
(ii) Take $\chi\in\{(\varepsilon,\varepsilon),(\varepsilon,\varepsilon^{2}),(\varepsilon^{2},\varepsilon),(\varepsilon^{2},\varepsilon^{2}),(1,\varepsilon),(1,\varepsilon^{2})\}$. None of $b,a^{3}b,a^{6}b$ belongs to $\ker(\chi)$, hence by Lemma 8.12, $\mathrm{Stab}_{G}(v)\neq\{1_{G}\}$ if and only if $\mathrm{Stab}_{G}(v)\nsubseteq\ker(\chi)$. Therefore the inclusion $\supseteq$” holds by Lemma 3.10. We turn to the proof of the reverse inclusion $\subseteq$”. So assume that all elements of $K[W_{1}\oplus W_{2}]^{G,\chi}_{\leq 9}$ vanish at $(w_{1},w_{2})$. We have to show that $\mathrm{Stab}_{G}(w_{1},w_{2})\neq\{1_{G}\}$.
Case I.: $\chi=(\varepsilon,\varepsilon)$. Consider the following relative invariants in $K[W_{1}\oplus W_{2}]^{G,(\varepsilon,\varepsilon)}$:
$f_{(\varepsilon,\varepsilon)}^{(1)}$ $:=x_{1}^{3}+\varepsilon^{2}x_{2}^{3}+\varepsilon x_{3}^{3}\qquad$ $h_{(\varepsilon,\varepsilon)}^{(1)}:=y_{1}^{6}+\varepsilon^{2}y_{2}^{6}+\varepsilon y_{3}^{6}$
$f_{(\varepsilon,\varepsilon)}^{(2)}$ $:=x_{1}^{5}x_{3}+\varepsilon^{2}x_{2}^{5}x_{1}+\varepsilon x_{3}^{5}x_{2}\qquad$ $h_{(\varepsilon,\varepsilon)}^{(2)}:=y_{1}^{2}y_{2}+\varepsilon^{2}y_{2}^{2}y_{3}+\varepsilon y_{3}^{2}y_{1}$
$f_{(\varepsilon,\varepsilon)}^{(3)}$ $:=x_{1}^{7}x_{3}^{2}+\varepsilon^{2}x_{2}^{7}x_{1}^{2}+\varepsilon x_{3}^{7}x_{2}^{2}\qquad$ $h_{(\varepsilon,\varepsilon)}^{(3)}:=y_{1}^{5}y_{3}^{4}+\varepsilon^{2}y_{2}^{5}y_{1}^{4}+\varepsilon y_{3}^{5}y_{2}^{4}$
$k_{(\varepsilon,\varepsilon)}^{(1)}$ $:=x_{1}y_{1}+\varepsilon^{2}x_{2}y_{2}+\varepsilon x_{3}y_{3}\qquad$ $h_{(\varepsilon,\varepsilon)}^{(4)}:=y_{1}y_{2}y_{3}(y_{1}^{3}+\varepsilon^{2}y_{2}^{3}+\varepsilon y_{3}^{3})$
$k_{(\varepsilon,\varepsilon)}^{(2)}$ $:=x_{1}^{2}y_{1}^{5}+\varepsilon^{2}x_{2}^{2}y_{2}^{5}+\varepsilon x_{3}^{2}y_{3}^{5}\qquad$
The above relative invariants have degree at most $9$, so each of them vanishes at $(w_{1},w_{2})\in W_{1}\oplus W_{2}$. In particular, $0=f_{(\varepsilon,\varepsilon)}^{(1)}(w_{1})=f_{(\varepsilon,\varepsilon)}^{(2)}(w_{1})=f_{(\varepsilon,\varepsilon)}^{(3)}(w_{1})$. Then
$0=\det\begin{bmatrix}x_{1}^{3}&x_{2}^{3}&x_{3}^{3}\\ x_{1}^{5}x_{3}&x_{2}^{5}x_{1}&x_{3}^{5}x_{2}\\ x_{1}^{7}x_{3}^{2}&x_{2}^{7}x_{1}^{2}&x_{3}^{7}x_{2}^{2}\end{bmatrix}(w_{1})=(x_{1}x_{2}x_{3})^{4}(x_{1}x_{2}-x_{3}^{2})(x_{1}x_{3}-x_{2}^{2})(x_{2}x_{3}-x_{1}^{2}))(w_{1}).$
If $x_{j}(w_{1})=0$ for some $j\in\{1,2,3\}$, then $0=f_{(\varepsilon,\varepsilon)}^{(1)}(w_{1})=f_{(\varepsilon,\varepsilon)}^{(2)}(w_{1})=f_{(\varepsilon,\varepsilon)}^{(3)}(w_{1})$ implies $0=x_{1}(w_{1})=x_{2}(w_{1})=x_{3}(w_{1})$, and so $\mathrm{Stab}_{G}(w_{1},w_{2})=\mathrm{Stab}_{G}(w_{2})$.
Suppose next that $(x_{1}x_{2}x_{3})(w_{1})\neq 0$. Then one of $x_{1}x_{2}-x_{3}^{2}$, $x_{1}x_{3}-x_{2}^{2}$, $x_{2}x_{3}-x_{1}^{2}$ vanishes at $w_{1}$. Assume that
(32) $(x_{1}x_{2}-x_{3}^{2})(w_{1})=0$
(the other cases follow by cyclic symmetry). From (32) and $f_{(\varepsilon,\varepsilon)}^{(1)}(w_{1})=0$ we deduce
(33) $\varepsilon(x_{1}x_{2}x_{3})(w_{1})=-x_{1}^{3}(w_{1})-\varepsilon^{2}x_{2}^{3}(w_{1}).$
From (32), (33) and $f_{(\varepsilon,\varepsilon)}^{(2)}(w_{1})=0$ we deduce
00 0 $=(x_{1}^{5}x_{3}+\varepsilon^{2}x_{2}^{5}x_{1}+\varepsilon(x_{1}x_{2})^{2}x_{3}x_{2})(w_{1})$
$=(x_{1}^{5}x_{3}+\varepsilon^{2}x_{2}^{5}x_{1}-(x_{1}^{3}+\varepsilon^{2}x_{2}^{3})x_{1}x_{2}^{2})(w_{1})$
$=x_{1}^{4}(w_{1})(x_{1}x_{3}-x_{2}^{2})(w_{1}),$
so we have
(34) $(x_{1}x_{3})(w_{1})=(x_{2}^{2})(w_{1})=0.$
From $0=f_{(\varepsilon,\varepsilon)}^{(3)}(w_{1})$, (32) and (34) we deduce
00 0 $=(x_{1}^{5}(x_{1}x_{3})^{2}+\varepsilon^{2}x_{2}^{7}x_{1}^{2}+\varepsilon(x_{1}x_{2})^{3}x_{2}^{2}x_{3})(w_{1})$
$=(x_{1}^{5}x_{2}^{4}+\varepsilon^{2}x_{2}^{7}x_{1}^{2}+\varepsilon x_{2}^{2}x_{1}^{2}x_{2}^{5})(w_{1})$
$=(x_{1}^{2}x_{2}^{4})(w_{1})(x_{1}^{3}-x_{2}^{3})(w_{1}),$
so we have
(35) $x_{1}^{3}(w_{1})=x_{2}^{3}(w_{1}).$
So $x_{1}(w_{1})=x_{2}(w_{1})$ or $x_{1}(w_{1})=\varepsilon x_{2}(w_{1})$ or $x_{1}(w_{1})=\varepsilon^{2}x_{2}(w_{1})$. Taking into account (32) and (34) we conclude that $w_{1}$ is a non-zero scalar multiple of $[1,1,1]^{T}$,$[\varepsilon^{2},\varepsilon,1]^{T}$, or $[\varepsilon,\varepsilon^{2},1]^{T}\}$. By Lemma 8.12 it follows that $\mathrm{Stab}_{G}(w_{1})\in\{\langle b\rangle,\langle a^{3}b\rangle,\langle a^{6}b\rangle\}$. In particular, we are done if $w_{2}=0$, since then $\mathrm{Stab}_{G}(w_{1},w_{2})=\mathrm{Stab}_{G}(w_{1})$.
Assume next that $w_{2}\neq 0$. Let us take into account that $0=h_{(\varepsilon,\varepsilon)}^{(1)}(w_{2})=h_{(\varepsilon,\varepsilon)}^{(2)}(w_{2})=h_{(\varepsilon,\varepsilon)}^{(3)}(w_{2})=h_{(\varepsilon,\varepsilon)}^{(4)}(w_{2})$. Since $w_{2}\neq 0$, the above equalities imply that $y_{j}(w_{2})\neq 0$ for $j=1,2,3$. Therefore $0=h_{(\varepsilon,\varepsilon)}^{(4)}(w_{2})$ implies
(36) $(y_{1}^{3}+\varepsilon^{2}y_{2}^{3}+\varepsilon y_{3}^{3})(w_{2})=0.$
It follows that
$0=\det\begin{bmatrix}1&1&1\\ y_{1}^{3}&y_{2}^{3}&y_{3}^{3}\\ y_{1}^{6}&y_{2}^{6}&y_{3}^{6}\end{bmatrix}(w_{2})=(y_{2}^{3}-y_{1}^{3})(w_{2})(y_{3}^{3}-y_{1}^{3})(w_{2})(y_{3}^{3}-y_{2}^{3})(w_{2}).$
Assume that
(37) $y_{1}^{3}(w_{2})=y_{2}^{3}(w_{2})$
(the cases when $y_{1}^{3}(w_{2})=y_{3}^{3}(w_{2})$ or $y_{2}^{3}(w_{2})=y_{3}^{3}(w_{2})$ follow by cyclic symmetry). From (36), (37) and $0=h_{(\varepsilon,\varepsilon)}^{(2)}(w_{2})$ we deduce
$0=\det\begin{bmatrix}1&1&1\\ y_{1}^{2}y_{2}&y_{2}^{2}y_{3}&y_{3}^{2}y_{1}\\ y_{1}^{3}&y_{2}^{3}&y_{3}^{3}\end{bmatrix}(w_{2})=\det\begin{bmatrix}1&0&1\\ y_{1}^{2}y_{2}&y_{2}(y_{2}y_{3}-y_{1}^{2})&y_{3}^{2}y_{1}\\ y_{1}^{3}&0&y_{3}^{3}\end{bmatrix}(w_{2}),$
hence
(38) $y_{1}^{2}(w_{2})=(y_{2}y_{3})(w_{2})$
or
(39) $y_{1}^{3}(w_{2})=y_{3}^{3}(w_{2}).$
Now (37), (38) clearly imply that $w_{2}$ is a non-zero scalar multiple of $[1,1,1]^{T}$, $[\varepsilon,\varepsilon^{2},1]^{T}$, or $[\varepsilon^{2},\varepsilon,1]^{T}\}$, implying in turn by Lemma 8.12 that $\mathrm{Stab}_{G}(w_{2})\in\{\langle b\rangle,\langle a^{3}b\rangle,\langle a^{6}b\rangle\}$.
If (39) holds, then together with (37) it implies that $y_{2}(w_{2})=\nu_{2}y_{1}(w_{2})$ and $y_{3}(w_{2})=\nu_{3}y_{1}(w_{2})$, where $\{\nu_{2},\nu_{3}\}\in\{1,\varepsilon,\varepsilon^{2}\}$. Recall that $y_{1}(w_{2})\neq 0$. From $h_{(\varepsilon,\varepsilon)}^{(2)}(w_{2})=0$ we infer $\nu_{2}+\varepsilon^{2}\nu_{2}^{2}\nu_{3}+\varepsilon\nu_{3}^{2}=0$, and hence $\nu_{2}\nu_{3}=1$ or $\nu_{2}\nu_{3}=\varepsilon^{2}$. On the other hand, $h_{(\varepsilon,\varepsilon)}^{(3)}(w_{2})=0$ yields $\nu_{3}+\varepsilon^{2}\nu_{2}^{2}+\varepsilon\nu_{3}^{2}\nu_{2}=0$, and hence $\nu_{2}\nu_{3}=1$ or $\nu_{2}\nu_{3}=\varepsilon$. Thus necessarily we have $\nu_{2}\nu_{3}=1$. So $\nu_{2}=\nu_{3}=1$ (i.e. $w_{2}\in K[1,1,1]^{T}$) or $(\nu_{2},\nu_{3})=(\varepsilon,\varepsilon^{2})$ (i.e. $w_{2}\in K[\varepsilon,\varepsilon^{2},1]^{T}$) or $(\nu_{2},\nu_{3})=(\varepsilon^{2},\varepsilon)$ (i.e. $w_{2}\in K[\varepsilon^{2},\varepsilon,1]^{T}$). It follows by Lemma 8.12 that $\mathrm{Stab}_{G}(w_{2})\in\{\langle b\rangle,\langle a^{3}b\rangle,\langle a^{6}b\rangle\}$.
Thus we showed that both $w_{1}$ and $w_{2}$ are scalar multiples of $[1,1,1]^{T}$, $[\varepsilon^{2},\varepsilon,1]^{T}$, or $[\varepsilon,\varepsilon^{2},1]^{T}$. If $w_{1}=0$ or $w_{2}=0$, then we have the desired $\mathrm{Stab}_{G}(w_{1},w_{2})\neq\{1_{G}\}$ by Lemma 8.12. If both $w_{1}$ and $w_{2}$ are non-zero scalar multiples of $[1,1,1]^{T}$, $[\varepsilon^{2},\varepsilon,1]^{T}$, or $[\varepsilon,\varepsilon^{2},1]^{T}$, then again using Lemma 8.12 one can verify by direct computation that for such a pair $(w_{1},w_{2})$, we have
$k_{(\varepsilon,\varepsilon)}^{(1)}(w_{1},w_{2})=0=k_{(\varepsilon,\varepsilon)}^{(2)}(w_{1},w_{2})\iff\mathrm{Stab}_{G}(w_{1})=\mathrm{Stab}_{G}(w_{2}).$
Therefore $\mathrm{Stab}_{G}(w_{1},w_{2})=\mathrm{Stab}_{G}(w_{1})\neq\{1_{G}\}$.
Case II.: $\chi=(1,\varepsilon)$. Consider the following relative invariants in $K[W_{1}\oplus W_{2}]^{G,(1,\varepsilon)}$:
$f_{(1,\varepsilon)}^{(1)}$ $:=(x_{1}x_{2}x_{3})^{2}(x_{1}^{3}+\varepsilon^{2}x_{2}^{3}+\varepsilon x_{3}^{3})\qquad$ $h_{(1,\varepsilon)}^{(1)}:=y_{1}^{2}y_{3}+\varepsilon^{2}y_{2}^{2}y_{1}+\varepsilon y_{3}^{2}y_{2}$
$f_{(1,\varepsilon)}^{(2)}$ $:=x_{1}x_{2}x_{3}(x_{1}^{6}+\varepsilon^{2}x_{2}^{6}+\varepsilon x_{3}^{6})\qquad$ $h_{(1,\varepsilon)}^{(2)}:=y_{1}^{4}y_{3}^{2}+\varepsilon^{2}y_{2}^{4}y_{1}^{2}+\varepsilon y_{3}^{4}y_{2}^{2}$
$f_{(1,\varepsilon)}^{(3)}$ $:=x_{1}x_{2}^{2}+\varepsilon^{2}x_{2}x_{3}^{2}+\varepsilon x_{3}x_{1}^{2}\qquad$ $h_{(1,\varepsilon)}^{(3)}:=y_{1}^{9}+\varepsilon^{2}y_{2}^{9}+\varepsilon y_{3}^{9}$
$f_{(1,\varepsilon)}^{(4)}$ $:=x_{1}^{9}+\varepsilon^{2}x_{2}^{9}+\varepsilon x_{3}^{9}\qquad$ $k_{(1,\varepsilon)}^{(1)}:=x_{1}y_{2}+\varepsilon^{2}x_{2}y_{3}+\varepsilon x_{3}y_{1}$
$k_{(1,\varepsilon)}^{(2)}:=x_{1}^{2}y_{2}^{2}+\varepsilon^{2}x_{2}^{2}y_{3}^{2}+\varepsilon x_{3}^{2}y_{1}^{2}.$
The above relative invariants have degree at most $9$, so each of them vanishes at $(w_{1},w_{2})\in W_{1}\oplus W_{2}$. If $x_{j}(w_{1})=0$ for some $j\in\{1,2,3\}$, then $0=f_{(1,\varepsilon)}^{(3)}(w_{1})=f_{(1,\varepsilon)}^{(4)}(w_{1})$ implies $w_{1}=0$, and thus $\mathrm{Stab}_{G}(w_{1},w_{2})=\mathrm{Stab}_{G}(w_{2})$.
Otherwise $(x_{1}x_{2}x_{3})(w_{1})\neq 0$, and $0=f_{(1,\varepsilon)}^{(1)}(w_{1})=f_{(1,\varepsilon)}^{(2)}(w_{1})$ implies
$0=\det\begin{bmatrix}1&1&1\\ x_{1}^{3}(w_{1})&x_{2}^{3}(w_{1})&x_{3}^{3}(w_{1})\\ x_{1}^{6}(w_{1})&x_{2}^{6}(w_{1})&x_{3}^{6}(w_{1})\end{bmatrix}=(x_{2}^{3}-x_{1}^{3})(x_{3}^{3}-x_{1}^{3})(x_{3}^{3}-x_{2}^{3})(w_{1}).$
By symmetry it is sufficient to deal with the case when
$x_{1}^{3}(w_{1})=x_{2}^{3}(w_{1}).$
By $0=f_{(1,\varepsilon)}^{(1)}(w_{1})$, $0=f_{(1,\varepsilon)}^{(3)}$ and $(x_{1}x_{2}x_{3})^{2}(w_{1})\neq 0$ we have
$0=\det\begin{bmatrix}1&1&1\\ x_{1}^{3}&x_{2}^{3}&x_{3}^{3}\\ x_{1}x_{2}^{2}&x_{2}x_{3}^{2}&x_{3}x_{1}^{2}\end{bmatrix}(w_{1}).$
Taking into account $x_{1}^{3}(w_{1})=x_{2}^{3}(w_{1})$ we end up with
(40) $0=x_{2}(x_{1}x_{2}-x_{3}^{2})(w_{1})(x_{3}^{3}-x_{1}^{3})(w_{1}).$
If $(x_{1}x_{2})(w_{1})=x_{3}^{2}(w_{1})$, then we have
$0=f_{(1,\varepsilon)}^{(3)}(w_{1})=x_{1}x_{2}^{2}+\varepsilon^{2}x_{2}(x_{1}x_{2})+\varepsilon x_{3}x_{1}^{2})(w_{1})=\varepsilon x_{1}(w_{1})(x_{1}x_{3}-x_{2}^{2})(w_{1}).$
Thus $(x_{1}x_{3})(w_{1})=x_{2}^{2}(w_{1})$, and this together with $(x_{1}x_{2})(w_{1})=x_{3}^{2}(w_{1})$ implies that $w_{1}$ is a non-zero scalar multiple of $[1,1,1]^{T}$, $[\varepsilon,\varepsilon^{2},1]^{T}$, or $[\varepsilon^{2},\varepsilon,1]^{T}$.
If $x_{3}^{3}(w_{1})=x_{1}^{3}(w_{1})$ (the other alternative from (40), then (recall that $x_{2}^{3}(w_{1})=x_{1}^{3}(w_{1})$) there exist some cubic roots $\nu_{2},\nu_{3}$ of $1$ such that $w_{1}$ is a non-zero scalar multiple of $[1,\nu_{2},\nu_{3}]$. Then $0=f_{(1,\varepsilon)}^{(3)}(w_{1})$ reduces to
$\nu_{2}^{2}+\varepsilon^{2}\nu_{2}\nu_{3}^{2}+\varepsilon\nu_{3}=0.$
The sum of three cubic roots of $1$ is zero only if the three summands are $1,\varepsilon,\varepsilon^{2}$ in some order. We conclude that $w_{1}$ is a non-zero scalar multiple of $[1,1,1]^{T}$, $[\varepsilon,\varepsilon^{2},1]^{T}$, or $[\varepsilon^{2},\varepsilon,1]^{T}$.
In particular, if $w_{2}=0$, then $\mathrm{Stab}_{G}(w_{1},w_{2})=\mathrm{Stab}_{G}(w_{1})\neq\{1_{G}\}$ by Lemma 8.12, and we are done.
Assume next that $w_{2}\neq 0$. Then by $0=h_{(1,\varepsilon)}^{(1)}(w_{2})=h_{(1,\varepsilon)}^{(3)}(w_{2})$ we deduce that none of $y_{1}(w_{2})$, $y_{2}(w_{2})$, $y_{3}(w_{2})$ is zero. From $0=h_{(1,\varepsilon)}^{(1)}(w_{2})=h_{(1,\varepsilon)}^{(2)}(w_{2})$ we get
$0=\det\begin{bmatrix}1&1&1\\ y_{1}y_{2}^{2}&y_{2}y_{3}^{2}&y_{3}y_{1}^{2}\\ y_{1}^{2}y_{2}^{4}&y_{2}^{2}y_{3}^{4}&y_{3}^{2}y_{1}^{4}\end{bmatrix}(w_{2})=(y_{1}y_{2}y_{3})(y_{1}y_{3}-y_{2}^{2})(y_{3}^{2}-y_{1}y_{2})(y_{1}^{2}-y_{2}y_{3})(w_{2}).$
By symmetry we may assume that $y_{2}^{2}(w_{2})=(y_{1}y_{3})(w_{2})$. Then we have
$0=h_{(1,\varepsilon)}^{(1)}(w_{2})=(y_{1}^{2}y_{3}+\varepsilon^{2}(y_{1}y_{3})y_{1}+\varepsilon y_{3}^{2}y_{2})(w_{2})=\varepsilon y_{3}(w_{2})(y_{2}y_{3}-y_{1}^{2})(w_{2}).$
Thus we have
$\frac{y_{2}^{2}(w_{2})}{y_{1}(w_{2})}=y_{3}(w_{2})=\frac{y_{1}^{2}(w_{2})}{y_{2}(w_{2})},$
and so $y_{1}^{3}(w_{2})=y_{2}^{3}(w_{2})$. Obviously, this together with $y_{2}^{2}(w_{2})=y_{1}(w_{2})y_{3}(w_{2})$ means that $w_{2}$ is a non-zero scalar multiple of $[1,1,1]^{T}$, $[\varepsilon,\varepsilon^{2},1]^{T}$, or $[\varepsilon^{2},\varepsilon,1]^{T}$.
So both $w_{1}$ and $w_{2}$ belong to $K[1,1,1]^{T}\cup K[\varepsilon,\varepsilon^{2},1]^{T}\cup K[\varepsilon^{2},\varepsilon,1]^{T}$. If $w_{1}=0$ or $w_{2}=0$, then $\mathrm{Stab}_{G}(w_{1},w_{2})\neq\{1_{G}\}$ by Lemma 8.12. If both $w_{1}$ and $w_{2}$ are non-zero, then again using Lemma 8.12 one can verify by direct computation that
$k_{(1,\varepsilon)}^{(1)}(w_{1},w_{2})=0=k_{(1,\varepsilon)}^{(2)}(w_{1},w_{2})\iff\mathrm{Stab}_{G}(w_{1})=\mathrm{Stab}_{G}(w_{2}).$
Therefore $\mathrm{Stab}_{G}(w_{1},w_{2})=\mathrm{Stab}_{G}(w_{1})\neq\{1_{G}\}$.
The map $a\mapsto a^{2}$, $b\mapsto b$ extends to an automorphism $\alpha$ of $G$, and for $\chi=(\varepsilon,\varepsilon)$ we have $\chi\circ\alpha=(\varepsilon^{2},\varepsilon)$, whereas for $\chi=(\varepsilon,1)$ we have $\chi\circ\alpha=(\varepsilon^{2},1)$. So by Lemma 3.11, statement (ii) holds also for the weight $\chi=(\varepsilon^{2},\varepsilon)$ and $\chi=(\varepsilon^{2},1)$. The only information on $\varepsilon$ used in the constructions of relative invariants above was that it has multiplicative order $3$. Therefore we can replace it by the other element of multiplicative order $3$, namely by $\varepsilon^{2}$, and we get that (ii) holds also for the weights $\chi\in\{\varepsilon,\varepsilon^{2}),(\varepsilon^{2},\varepsilon^{2}),(1,\varepsilon^{2})\}$. ∎
Theorem 8.14.
Assume that $K$ has characteristic zero, and it contains an element of multiplicative order $9$. Then we have the equality $\beta_{\mathrm{sep}}^{K}(\mathrm{M}_{27})=10$.
Proof.
In view of Lemma 3.6, take $v=(w_{1},w_{2},u),v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2},u^{\prime})\in W_{1}\oplus W_{2}\oplus U$ with
(41) $f(v)=f(v^{\prime})\text{ for all homogeneous }f\in K[W_{1}\oplus W_{2}\oplus U]^{G}\text{ where }\deg(f)\leq 10.$
We need to show that $G\cdot v=G\cdot v^{\prime}$. Replacing $v^{\prime}$ by an appropriate element in its $G$-orbit, we may assume by Proposition 8.11 that $(w_{1},w_{2})=(w^{\prime}_{1},w^{\prime}_{2})$. Moreover, $G\cdot u=G\cdot u^{\prime}$, since $\mathsf{D}(G/G^{\prime})=\mathsf{D}(\mathrm{C}_{3}\times\mathrm{C_{3}})=5$. Therefore it is sufficient to deal with the case when $(w_{1},w_{2})\in W_{1}\oplus W_{2}$ is non-zero.
Case I.: $\mathrm{Stab}_{G}(w_{1},w_{2})\neq\{1_{G}\}$. Take any $\langle b\rangle$-invariant monomial $m\in K[U]^{\langle b\rangle}$ with $\deg(m)\leq 3$. It belongs to $K[W_{1}\oplus W_{2}\oplus U]^{G,\chi}$ for some $\chi\in\{\varepsilon,1),(\varepsilon^{2},1),(1,1)\}$. By Lemma 8.13 (i) there exists an $f\in K[W_{1}\oplus W_{2}\oplus U]^{G,\chi^{-1}}$ of degree at most $6$ with $f(w_{1},w_{2})\neq 0$. Note that $fm\in K[W_{1}\oplus W_{2}\oplus U]^{G}$ has degree at most $9$. It follows by (41) that $(fm)(v)=(fm)(v^{\prime})$, implying in turn that $m(u)=m(u^{\prime})$. This holds for all monomials $m\in K[U]^{\langle b\rangle}$ with $\deg(m)\leq 3$. Since $\mathsf{D}(\langle b\rangle)=3$, we conclude that $u$ and $u^{\prime}$ belong to the same $\langle b\rangle$-orbit. Note that $\mathrm{Stab}_{G}(w_{1},w_{2})G^{\prime}=\langle b\rangle G^{\prime}$ by Lemma 8.12. So the $\langle b\rangle$-orbits of $u$ and $u^{\prime}$ coincide with their $\mathrm{Stab}_{G}(w_{1},w_{2})$-orbits. Thus $\mathrm{Stab}_{G}(w_{1},w_{2})\cdot u=\mathrm{Stab}_{G}(w_{1},w_{2})\cdot u^{\prime}$, and therefore $G\cdot(w_{1},w_{2},u)=G\cdot(w^{\prime}_{1},w^{\prime}_{2},u^{\prime})$.
Case II.: $\mathrm{Stab}_{G}(w_{1},w_{2})=\{1_{G}\}$. We claim that $u=u^{\prime}$; that is, we claim that $u_{\chi}=u^{\prime}_{\chi}$ for all $\chi\in\widehat{G}$. This is obvious for $\chi=(1,1)\in\widehat{G}$, since $u$ and $u^{\prime}$ have the same $G$-orbit. For $\chi\in\widehat{G}\setminus\{(1,1)\}$, by Lemma 8.13 there exists an $f\in K[W_{1}\oplus W_{2}\oplus U]^{G,\chi^{-1}}$ with $\deg(f)\leq 9$ and $f(w_{1},w_{2})\neq 0$. Then $ft_{\chi}$ is a $G$-invariant of degree at most $10$. Thus by (41) we have $(ft_{\chi})(w_{1},w_{2},u)=(ft_{\chi})(w_{1},w_{2},u^{\prime})$, implying in turn that $t_{\chi}(u)=t_{\chi}(u^{\prime})$, i.e. $u_{\chi}=u^{\prime}_{\chi}$. This finishes the proof of the inequality $\beta_{\mathrm{sep}}^{K}(G)\leq 10$.
To see the reverse inequality, consider the $G$-module $W_{1}\oplus U_{(1,\varepsilon^{2})}$. Consider the vectors $v:=([1,0,0]^{T},\varepsilon)$ and $v^{\prime}:=([1,0,0]^{T},\varepsilon^{2})$. We have $(f_{(1,\varepsilon)}^{(4)}t_{(1,\varepsilon^{2})})(v)=\varepsilon$, whereas $(f_{(1,\varepsilon)}^{(4)}t_{(1,\varepsilon^{2})})(v^{\prime})=\varepsilon^{2}$, so the invariant $(f_{(1,\varepsilon)}^{(4)}t_{(1,\varepsilon^{2})}$) separates $v$ and $v^{\prime}$. We claim that no $G$-invariant of degree at most $9$ separates $v$ and $v^{\prime}$. The elements of $K[W_{1}]$ and $K[U_{(1,\varepsilon^{2})}]^{G}=K[t_{(1,\varepsilon)}^{3}]$ agree on $v$ and $v^{\prime}$. Suppose for contradiction that there exists a multihomogeneous invariant $f=ht_{(1,\varepsilon^{2})}$ (respectively $f=ht_{(1,\varepsilon^{2})}^{2}$) of degree at most $9$ with $f(v)\neq f(v^{\prime})$, where $h\in K[W_{1}]^{G,(1,\varepsilon)}$ (respectively $h\in K[W_{1}]^{G,(1,\varepsilon^{2})}$). Then $h$ has a monomial of the form $x^{d}$ with non-zero coefficients (since $0=x_{2}(v)=x_{3}(v)=x_{2}(v^{\prime})=x_{3}(v^{\prime})$). Moreover, $x_{1}^{d}$ must be an $\langle a\rangle$-invariant monomial. However, the smallest $d$ for which $x_{1}^{d}$ is $\langle a\rangle$-invariant is $9$. This is a contradiction, because the degree of $h$ is strictly less than $9$. The inequality $\beta_{\mathrm{sep}}(G,W_{1}\oplus U_{(1,\varepsilon)})\geq 10$ is proved. ∎